### 27 Feb 2026 v1

# arXiv:2603.00301v1[math.NT] 27 Feb 2026

Abstract.
For a real quadratic field $\mathbb{Q}(\sqrt{d})$, we study the norm-form energy $N=S_{\zeta}^{2}-d\cdot S_{L}^{2}$, where $S_{\zeta}$ and $S_{L}$ are Lorentzian-weighted zero sums with $w(\rho)=2/(\tfrac{1}{4}+\gamma^{2})$. We prove three main results. (1) Spacelike spectral data: $N<0$ unconditionally for all squarefree $d>1$, as a consequence of a low-lying zero dominance theorem proved via explicit zero-counting. (2) Effective density bound: at each verified truncation level $M$, $\operatorname{dens}\{N>0\}\leq 2\|f_{S_{L}^{(M)}}\|_{\infty}\cdot(W_{1}(\zeta)/\sqrt{d}+\epsilon_{M})$, established unconditionally via Jacobi–Anger resonance analysis. (3) Exact asymptotic: under the computationally verified hypothesis that the infinite resonance lattice $\Lambda_{\infty}$ has finite rank (verified for $M\leq 20$, where rank =0=0), the sharp asymptotic $\operatorname{dens}\{N>0\}=C(d)/\sqrt{d}+o(1/\sqrt{d})$ holds. For $d=5$, $C(5)=2\,f_{S_{L}}(0)\cdot\mathbb{E}[|S_{\zeta}|]=0.1191$; the constant depends on $d$ through the zeros of $L(s,\chi_{d})$, and $C(d)=O(1/\log d)$ as $d\to\infty$.
Key words and phrases:
Dirichlet $L$-functions, quadratic norm form, density bounds, Lorentzian spectral weights, Jacobi–Anger decomposition, Jessen–Wintner theorem, Besicovitch almost periodic functions
2020 Mathematics Subject Classification:
Primary 11M06; Secondary 11M26, 11R42, 42A75
1. Introduction
Let $d>1$ be squarefree and let $\chi_{d}$ denote the Kronecker symbol of the fundamental discriminant $\Delta_{K}$ of $K=\mathbb{Q}(\sqrt{d})$. The Dedekind zeta function factors as
$\zeta_{K}(s)=\zeta(s)\cdot L(s,\chi_{d}),$ (1)
coupling the zeros of the Riemann zeta function to those of the quadratic Dirichlet $L$-function. We ask whether this coupling can be detected spectrally, by weighting zeros so that low-lying ones dominate. The answer is yes: the resulting quadratic form on weighted zero sums is always negative (the $L$-function zeros win), and the set where it becomes positive has density $O(1/\sqrt{d})$. Both statements hold unconditionally in the Euler-product formulation; on the zero side, they hold at any verified truncation level and extend to the full series under a computationally verified finite-rank condition (rank =0=0 for the first 20 zeros).
For a zero $\rho=\tfrac{1}{2}+i\gamma$ on the critical line, define the Lorentzian weight
$w(\rho)=\frac{2}{\tfrac{1}{4}+\gamma^{2}}.$ (2)
This weight emphasizes low-lying zeros: a zero at height $\gamma=1$ contributes weight $1.6$, while a zero at height $\gamma=14$ contributes weight $0.01$. Define the weighted zero sums
$S_{\zeta}=\sum_{\zeta(\rho)=0}w(\rho),\qquad S_{L}=\sum_{L(\rho,\chi_{d})=0}w(\rho),$ (3)
and the norm-form energy
$N=S_{\zeta}^{2}-d\cdot S_{L}^{2}.$ (4)
The signature $(1,1)$ of this quadratic form reflects the Galois structure of $K/\mathbb{Q}$: the trivial eigenspace contributes positively, the sign eigenspace negatively.
1.1. Main results
We prove unconditionally that:
  1. (1)
    Spacelike spectral data (Corollary 3.5): $N=S_{\zeta}^{2}-d\cdot S_{L}^{2}<0$ for all squarefree $d>1$. This follows from the low-lying zero dominance theorem (Theorem 3.1: $S_{L}>S_{\zeta}^{*}$) via explicit zero-counting, and extends to Cesàro averages (Theorem 4.9: $\langle N\rangle_{T}<0$ for all $T>0$).
  2. (2)
    Effective density bound (Theorem 8.5): $\operatorname{dens}\{N>0\}\leq 2\|f_{S_{L}^{(M)}}\|_{\infty}\cdot(W_{1}(\zeta)/\sqrt{d}+\epsilon_{M})$ unconditionally at any verified truncation level $M$.
  3. (3)
    Density rate and exact asymptotic (Corollary 8.8, Theorem 9.4): $\operatorname{dens}\{N>0\}=O(1/\sqrt{d})$, sharpened to $\operatorname{dens}\{N>0\}=C(d)/\sqrt{d}+o(1/\sqrt{d})$, under the hypothesis that the infinite resonance lattice $\Lambda_{\infty}$ has finite rank $d_{\infty}<\infty$ (verified $d_{\infty}=0$ for $M\leq 20$). For $d=5$, $C(5)=0.1191$; the constant $C(d)=O(1/\log d)$ as $d\to\infty$.
These results do not assume the Riemann Hypothesis for either function. The spacelike property is robust to the choice of weight function (see Section 13.4). As a byproduct of the low-lying zero dominance proof, we obtain a universal first-zero bound $\gamma_{1}^{\prime}(d)<14.13$ for all fundamental discriminants (Proposition 3.9).
1.2. Logical structure and the GSH barrier
The progression of results reveals the fundamental role of $\mathbb{Q}$-linear independence:
Result Assumes Bound Method
Thm 4.16: Density vanishing Hypothesis 4.4(ii) $O(1/\sqrt{d})$ Jessen–Wintner
Thm 4.18: Euler product density nothing $O(1/\sqrt{d})$ $\mathbb{Q}$-indep. of $\log p$
The barrier: zero ordinates lack unconditional $\mathbb{Q}$-independence
Thm 6.9: Unconditional finiteness nothing $f_{S_{L}^{(M)}}(0)<\infty$ Jacobi–Anger + Bessel
Thm 8.5: Effective bound nothing $2\|f\|_{\infty}(W_{1}/\!\sqrt{d}+\epsilon_{M})$ Containment + stability
Cor. 8.8: Density rate $d_{\infty}<\infty$ $O(1/\sqrt{d})$ $M(d)$ + uniform bound
Thm 7.7: Limiting density Verified ($M=20$) $f_{S_{L}}(0)<\infty$ DCT + resonance absence
Thm 9.4: Exact formula Verified ($M=20$) $C/\sqrt{d}$ Marginal densities
The key insight is that the Jessen–Wintner approach requires $\mathbb{Q}$-linear independence to identify the Cesàro distribution with a product of independent random variables. On the Euler product side, this independence follows from unique factorization. On the zero side, it is precisely the Grand Simplicity Hypothesis. The unconditional effective bound bypasses this barrier at a fixed truncation level via the stability lemma; the $O(1/\sqrt{d})$ rate requires the truncation level to grow with $d$, which in turn requires a uniform bound on the density $\|f_{S_{L}^{(M)}}\|_{\infty}$. Under verified resonance absence ($d_{\infty}=0$), this uniform bound holds because the unsigned Bessel integral is monotone decreasing in $M$; more generally, it suffices that the infinite resonance lattice $\Lambda_{\infty}$ has finite rank $d_{\infty}<\infty$.
1.3. Series context
This paper is the first in the series Persistent Heuristics. The Lorentzian weight arises from a spectral interpretation of the Weil explicit formula that emphasizes low-lying zeros over high zeros. While this emphasis might initially appear restrictive, it reveals genuine invariants: the spacelike property holds for any weight satisfying mild decay conditions (Section 13.4), and the density bounds depend on the weight only through computable constants.
1.4. Organization
Section 2 defines the Lorentzian weight and establishes its properties. Section 3 proves the low-lying zero dominance and spacelike theorems. Section 4 extends the negativity to Cesàro averages and establishes conditional density bounds under GSH, revealing the $\mathbb{Q}$-independence barrier. Sections 6 and 7 establish per-function density bounds via Jacobi–Anger analysis and finite truncation. Section 8 combines these with a stability argument to prove the effective density bound, and Section 9 derives the exact formula under verified resonance absence. Section 12 analyzes the unique null crossing on the real axis. Section 13 discusses connections and open problems. Appendix A provides numerical verification.
1.5. Conventions
Figure 1. The norm form $N(\delta)=S_{\zeta}(\delta)^{2}-5\,S_{L}(\delta)^{2}$ for $d=5$, computed using the first 100 zeros of each function with Lorentzian weights $w(\gamma)=2/(\frac{1}{4}+\gamma^{2})$. Panel (a): the oscillating signal over $[0,200]$, showing that $N$ is overwhelmingly negative. Panel (b): the local fraction of $\delta$-values where $N>0$, computed in sliding windows across $[0,2000]$; the measure of the set where $N>0$ is approximately $5.5\%$ over $[0,2000]$, consistent with the predicted $O(1/\sqrt{d})$ density.
Sums over zeros $\sum_{\rho}$ run over the upper half-plane $\gamma>0$, with each term representing a conjugate pair $\rho,\bar{\rho}$. The zero-counting function $N(T)$ counts zeros with $0<\gamma\leq T$. The Fourier transform convention is $\hat{f}(\xi)=\int_{-\infty}^{\infty}f(x)e^{-2\pi ix\xi}\,dx$.
Throughout, $\gamma_{k}$ denotes the $k$-th positive ordinate of $\zeta(s)$ (so $\gamma_{1}=14.134\ldots$), while $\gamma_{k}^{\prime}$ denotes the $k$-th positive ordinate of $L(s,\chi_{d})$ for the quadratic character under consideration. The unprimed/primed convention is maintained consistently: unprimed ordinates belong to $\zeta$, primed ordinates to the Dirichlet $L$-function.
Throughout, $\chi_{d}$ denotes the Kronecker symbol of the fundamental discriminant $\Delta_{K}$ of $\mathbb{Q}(\sqrt{d})$; the conductor is $q=|\Delta_{K}|\in\{d,4d\}$.
2. The Lorentzian Energy Functional
2.1. The Lorentzian weight
Definition 2.1 (Lorentzian weight).
For a zero $\rho=\beta+i\gamma$ in the critical strip, the Lorentzian weight is
$w(\rho)=\frac{2}{\tfrac{1}{4}+\gamma^{2}}$
when $\beta=\tfrac{1}{2}$. For zeros off the critical line, the unconditional weight is
$w(\rho)=\frac{2(\beta(1-\beta)+\gamma^{2})}{(\beta(1-\beta)+\gamma^{2})^{2}+4\gamma^{2}(\tfrac{1}{2}-\beta)^{2}}.$ (5)
The weight (5) reduces to (2) when $\beta=\tfrac{1}{2}$.
Proposition 2.2 (Critical line maximality).
For fixed $\gamma$ with $|\gamma|\geq 1/(2\sqrt{3})$, the weight $w(\beta,\gamma)$ is maximized at $\beta=\tfrac{1}{2}$.
Proof.
Set $\varepsilon=\tfrac{1}{2}-\beta$ and write $D=\tfrac{1}{4}-\varepsilon^{2}+\gamma^{2}$. Then
$w(\tfrac{1}{2},\gamma)-w(\beta,\gamma)=\frac{2\varepsilon^{2}(3\gamma^{2}-\tfrac{1}{4}+\varepsilon^{2})}{(\tfrac{1}{4}+\gamma^{2})(D^{2}+4\gamma^{2}\varepsilon^{2})}.$
For $|\gamma|\geq 1/(2\sqrt{3})$, the factor $3\gamma^{2}-\tfrac{1}{4}\geq 0$, so the expression is non-negative, with equality only at $\varepsilon=0$. ∎
Since every nontrivial zero of $\zeta(s)$ has $|\gamma|\geq 14.13>1/(2\sqrt{3})$, all estimates using the on-line weight $w(\rho)=2/(\tfrac{1}{4}+\gamma^{2})$ are unconditional upper bounds. For $L$-function zeros, values $\gamma^{\prime}<1/(2\sqrt{3})\approx 0.289$ are theoretically possible (and expected for large conductors), but irrelevant here: the general weight formula (5) applies unconditionally to all zeros regardless of $\gamma^{\prime}$, and the on-line formula is used only as an upper bound via Proposition 2.2.
2.2. Weighted zero sums
Definition 2.3 (Weighted zero sums).
For a primitive Dirichlet character $\chi$ with conductor $q$, define
$S_{\zeta}=\sum_{\rho:\,\zeta(\rho)=0}w(\rho),\qquad S_{L(\chi)}=\sum_{\rho:\,L(\rho,\chi)=0}w(\rho),$
where the sums run over nontrivial zeros in the upper half-plane, each representing a conjugate pair. For $\chi=\chi_{d}$, we write $S_{L}=S_{L(\chi_{d})}$.
Both sums converge absolutely: $w(\rho)=O(1/\gamma^{2})$ and the zero-counting function satisfies $N(T)=O(T\log T)$.
Proposition 2.4 (Explicit value of $S_{\zeta}$).
The weighted zeta zero sum satisfies
$S_{\zeta}\leq 0.04871=:S_{\zeta}^{*}.$ (6)
The tail bound uses the following explicit result for $\zeta(s)$.
Theorem 2.5 (Trudgian [12, Corollary 1]).
Let $N(T)$ denote the number of zeros $\rho=\beta+i\gamma$ of $\zeta(s)$ with $0<\beta<1$ and $0<\gamma<T$. For $T\geq T_{0}\geq e$,
$\left|N(T)-\frac{T}{2\pi}\log\frac{T}{2\pi e}-\frac{7}{8}\right|\leq 0.111\log T+0.275\log\log T+2.450+\frac{0.2}{T_{0}}.$
Proof of Proposition 2.4.
The first $6000$ verified zeta zeros are summed in ARB ball arithmetic [1], giving the certified value $S_{\zeta}^{(6000)}=0.04580\ldots$ . The tail $\sum_{\gamma>\gamma_{6000}}w(\gamma)$ is bounded by Abel summation against Theorem 2.5, using the exact antiderivative
$\int_{a}^{b}\frac{4t}{(\tfrac{1}{4}+t^{2})^{2}}\,dt=\frac{2}{\tfrac{1}{4}+a^{2}}-\frac{2}{\tfrac{1}{4}+b^{2}}$
on subintervals, with no numerical quadrature. The certified tail bound is $\leq 4.45\times 10^{-4}$. Combining:
$S_{\zeta}\leq 0.04625<0.04871=:S_{\zeta}^{*}.$
Full computational parameters are recorded in Appendix A, §A.3. ∎
2.3. The norm-form energy
Definition 2.6 (Norm-form energy).
The norm-form energy is
$N=S_{\zeta}^{2}-d\cdot S_{L}^{2}.$
The quadratic form $N(a,b)=a^{2}-db^{2}$ has signature $(1,1)$, corresponding to the eigenspace decomposition of the Galois group $\mathrm{Gal}(K/\mathbb{Q})\cong C_{2}$.
Definition 2.7 (Causal classification).
A vector $(x,y)\in\mathbb{R}^{2}$ with $x,y>0$ is timelike if $x^{2}-dy^{2}>0$, null if $x^{2}-dy^{2}=0$, and spacelike if $x^{2}-dy^{2}<0$.
The null cone consists of the lines $x=\pm\sqrt{d}\,y$. The spacelike region lies outside this cone. The terminology is borrowed by analogy from Lorentzian geometry, where the quadratic form $x^{2}-c^{2}t^{2}$ classifies vectors; the analogy is structural (both involve indefinite quadratic forms of signature $(1,1)$) but no physical connection is implied.
2.4. Eigenspace interpretation
The decomposition $(\zeta(s),L(s,\chi_{d}))$ corresponds to the $\pm 1$ eigenspaces of the Galois involution $\sigma:\sqrt{d}\mapsto-\sqrt{d}$ acting on $\zeta_{K}(s)=\zeta(s)\cdot L(s,\chi_{d})$. The norm form $N(a,b)=a^{2}-db^{2}$ is the unique indefinite $C_{2}$-invariant quadratic form up to scaling. The definite form $a^{2}+db^{2}$ is also $C_{2}$-invariant but irrelevant here since it has no null cone. The coefficient d-d arises from the field norm $N_{K/\mathbb{Q}}(a+b\sqrt{d})=a^{2}-db^{2}$.
2.5. Per-prime sign structure
The norm-form energy admits a per-prime decomposition that reveals the sign structure driving the negativity theorem.
Lemma 2.8 (Per-prime contributions).
For a prime $p$ with $\chi_{d}(p)=\chi\in\{-1,0,+1\}$, define
$S^{+}(p):=\frac{\log p}{\sqrt{p}-1}>0,\qquad S^{-}(p):=\frac{\chi\log p}{\sqrt{p}-\chi}.$
If $p$ is split ($\chi=+1$), then $S^{-}(p)=S^{+}(p)>0$. If $p$ is inert ($\chi=-1$), then $S^{-}(p)=-\log p/(\sqrt{p}+1)<0$. If $p$ is ramified ($\chi=0$), then $S^{-}(p)=0$.
Theorem 2.9 (Per-prime sign).
For $K=\mathbb{Q}(\sqrt{d})$, $d>1$ squarefree, and any prime $p$, the per-prime norm-form contribution $N_{p}=4\pi(S^{+}(p)-d\,S^{-}(p))$ satisfies:
  1. (a)
    If $p$ is split ($\chi_{d}(p)=+1$): $N_{p}=\frac{4\pi\log p}{\sqrt{p}-1}(1-d)<0$.
  2. (b)
    If $p$ is inert ($\chi_{d}(p)=-1$): $N_{p}=4\pi\log p\bigl(\frac{1}{\sqrt{p}-1}+\frac{d}{\sqrt{p}+1}\bigr)>0$.
  3. (c)
    If $p$ is ramified ($\chi_{d}(p)=0$): $N_{p}=\frac{4\pi\log p}{\sqrt{p}-1}>0$.
Proof.
Part (a): when $\chi_{d}(p)=+1$, we have $S^{+}(p)=S^{-}(p)$, so $N_{p}=4\pi S^{+}(p)(1-d)<0$ since $d>1$. Part (b): when $\chi_{d}(p)=-1$, $S^{-}(p)=-\log p/(\sqrt{p}+1)$, so $N_{p}=4\pi\log p(1/(\sqrt{p}-1)+d/(\sqrt{p}+1))>0$. Part (c): when $\chi_{d}(p)=0$, $S^{-}(p)=0$, so $N_{p}=4\pi S^{+}(p)>0$. ∎
Remark 2.10 (Sign inversion).
The $C_{2}$-invariant quadratic form inverts the algebraic hierarchy: split primes that contribute equally to both eigenspaces ($S^{+}=S^{-}$) receive the full penalty from the coefficient d-d, producing negative energy. Inert primes yield a sign reversal via $S^{-}<0$, producing positive energy contributions.
3. Negativity and Spacelike Theorems
3.1. Low-lying zero dominance
Theorem 3.1 (Low-lying zero dominance).
For every squarefree $d>1$, the $L$-function spectral sum dominates: $S_{L}>S_{\zeta}^{*}$.
The proof relies on the following explicit zero-counting theorem.
Theorem 3.2 (Bennett–Martin–O’Bryant–Rechnitzer [2, Theorem 1.1]).
Let $\chi$ be a primitive character with conductor $q>1$ and $T\geq 5/7$. Set $\ell=\log\bigl(q(T+2)/(2\pi)\bigr)$. If $\ell\leq 1.567$, then $N(T,\chi)=0$. If $\ell>1.567$, then
$\Bigl|N(T,\chi)-\frac{T}{\pi}\log\frac{qT}{2\pi e}-\frac{\chi(-1)}{4}\Bigr|\leq 0.22737\,\ell+2\log(1+\ell)-0.5.$
In particular, $N(T,\chi)\geq(T/\pi)\log(qT/2\pi e)-\chi(-1)/4-0.22737\,\ell-2\log(1+\ell)+0.5$.
Proof of Theorem 3.1.
We establish the inequality for all fundamental discriminants by combining Theorem 3.2 with verified computations for two small-conductor exceptions. Throughout, zeros are counted in the upper half-plane only; the weight $w(\rho)=2/(\tfrac{1}{4}+\gamma^{2})$ represents conjugate pairs.
Step 1: Conductors $q\geq 12$. Let $\chi_{d}$ be the Kronecker character with conductor $q=|\Delta_{K}|$. The conductor of a real quadratic field satisfies $q\in\{d,4d\}$ depending on $d\bmod 4$; the values $q\in\{5,8\}$ correspond to $d\in\{5,2\}$ and are treated separately below. The next smallest conductor is $q=12$ (for $d=3$).
For $T=11$, Theorem 3.2 gives the lower bound
$N(11,\chi_{d})\geq\frac{11}{\pi}\log\frac{11q}{2\pi e}-\frac{|\chi_{d}(-1)|}{4}-0.22737\,\ell-2\log(1+\ell)+0.5,$
where $\ell=\log(q\cdot 13/(2\pi))$. For $q=12$, this evaluates to $N(11,\chi_{d})\geq 3.80$; since $N(11,\chi_{d})$ is an integer, this implies at least four zeros below height $11$, and the bound increases with $q$. Each of these at least four zeros has imaginary part at most $11$, so each contributes weight at least $w(11)=2/(\tfrac{1}{4}+121)=2/121.25=0.01649$. Therefore
$S_{L}\geq 4\cdot w(11)=8/121.25=0.06598>0.04871=S_{\zeta}^{*}.$
Step 2: $d=5$ (conductor $q=5$). The BMOR bound gives only $N(11,\chi_{5})\geq 1.40$, which does not guarantee three zeros. We therefore verify the inequality directly. Computing the Hardy $Z$-function $Z_{\chi_{5}}(t)=e^{i\theta(t)}L(\tfrac{1}{2}+it,\chi_{5})$ and scanning for sign changes on $(0,12)$ yields two zeros below height $11$: $\gamma_{1}^{\prime}=6.6485$ and $\gamma_{2}^{\prime}=9.8314$, both verified to satisfy $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{5})|<10^{-15}$. (Any additional zeros below height $11$, if they exist, would only increase $S_{L}$.) Therefore
$S_{L}\geq w(\gamma_{1}^{\prime})+w(\gamma_{2}^{\prime})=\frac{2}{0.25+44.20}+\frac{2}{0.25+96.66}=0.04499+0.02064=0.06563>S_{\zeta}^{*}.$
Step 3: $d=2$ (conductor $q=8$). The Hurwitz decomposition $L(s,\chi_{2})=8^{-s}\sum_{a=1}^{8}\chi_{2}(a)\,\zeta(s,a/8)$, evaluated via the Hardy $Z$-function method (§A.2) at 500-bit ARB precision, yields three zeros below height $11$:
$\gamma_{1}^{\prime}=4.8999739970070365010,\quad\gamma_{2}^{\prime}=7.6284288417693978341,\quad\gamma_{3}^{\prime}=10.806588163861712014,$
each satisfying $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{2})|<10^{-149}$ (see Appendix C). Any additional zeros below height $11$ would only increase $S_{L}$. Therefore
$S_{L}\geq w(\gamma_{1}^{\prime})+w(\gamma_{2}^{\prime})+w(\gamma_{3}^{\prime})=\frac{2}{0.25+24.010}+\frac{2}{0.25+58.193}+\frac{2}{0.25+116.782}=0.08246+0.03421+0.01710=0.13377>S_{\zeta}^{*}.$
Since every squarefree $d>1$ has conductor $q\in\{5,8\}$ or $q\geq 12$, the three cases exhaust all possibilities. ∎
Remark 3.3 (Margins).
The margin $(S_{L}-S_{\zeta}^{*})/S_{\zeta}^{*}$ is smallest in the generic case: for $q=12$, the BMOR bound guarantees at least four zeros below height $11$, giving $S_{L}\geq 4\cdot 2/(0.25+121)=0.06598$, a margin of $(0.06598-0.04871)/0.04871=35.5\%$ over $S_{\zeta}^{*}$. The margin is largest for $d=2$, where three directly computed zeros give $S_{L}\geq 0.13375$, a margin of $174.6\%$. The two exceptional cases $d\in\{2,5\}$ are verified by direct computation of $L$-function zeros via the Hardy $Z$-function method (see Appendix C for the $d=2$ data).
Remark 3.4 (On first-zero sufficiency).
For most discriminants, the first zero alone gives $w(\gamma_{1}^{\prime})>S_{\zeta}^{*}$. The exception is $d=5$ ($\gamma_{1}^{\prime}=6.6485$), where $w(\gamma_{1}^{\prime})=0.04499<S_{\zeta}^{*}=0.04871$ and the second zero is required.
3.2. The spacelike corollary
Corollary 3.5 (Spacelike).
For every squarefree $d>1$, the norm-form energy satisfies $N=S_{\zeta}^{2}-d\cdot S_{L}^{2}<0$.
Proof.
By Proposition 2.4, $S_{\zeta}\leq S_{\zeta}^{*}=0.04871$. By Theorem 3.1, $S_{L}>S_{\zeta}^{*}\geq S_{\zeta}>0$. Since $d>1$:
$d\cdot S_{L}^{2}>d\cdot(S_{\zeta}^{*})^{2}\geq d\cdot S_{\zeta}^{2}>S_{\zeta}^{2},$
hence $N=S_{\zeta}^{2}-d\cdot S_{L}^{2}<0$. ∎
Remark 3.6 (Strict hierarchy).
The results form a strict chain. Let $r=S_{\zeta}/S_{L}$. The spacelike corollary requires $r<\sqrt{d}$. The low-lying zero dominance gives $r<1$. Since $r<1<\sqrt{d}<d$ for $d>1$, each result is strictly stronger than what the corollary requires. The actual ratio $r\approx 0.1$ to $0.5$ for small $d$, reflecting the severity of the low-lying zero phenomenon.
3.3. The negativity theorem for general characters
Theorem 3.7 (Negativity).
For every primitive nontrivial Dirichlet character $\chi$ with conductor $q\geq 3$:
$E_{\chi}(0):=S_{\zeta}-q\cdot S_{L(\chi)}<0.$
Proof.
The proof follows the same structure as Theorem 3.1. For $q\geq 4$, Theorem 3.2 guarantees a zero below height $T=10.5$, giving $q\cdot w(\gamma_{1}^{\prime})\geq 4\cdot 2/110.50=0.0724>S_{\zeta}^{*}$. For $q=3$, the BMOR bound at $T=15.7$ guarantees at least three zeros, each below height $15.7$, so $q\cdot S_{L(\chi)}\geq q\cdot 3\cdot w(15.7)=3\cdot 3\cdot 2/(0.25+15.7^{2})=18/246.74=0.0730>S_{\zeta}^{*}$. ∎
Corollary 3.8 (Linear deepening).
For $q\geq 3$:
$E_{\chi}(0)\leq S_{\zeta}^{*}-\frac{2q}{121.25}<0,$
so $|E_{\chi}(0)|$ grows at least linearly in the conductor.
Proof.
By (6), $S_{\zeta}\leq S_{\zeta}^{*}$. For $q\geq 4$, the BMOR bound guarantees a zero of $L(s,\chi)$ below height $T(q)\leq 10.5<11$, so $S_{L(\chi)}\geq w(11)=2/121.25$. For $q=3$, the BMOR bound at $T=15.7$ gives $S_{L(\chi)}\geq 3\cdot 2/(0.25+15.7^{2})=6/246.74>2/121.25$. In both cases, $q\cdot S_{L(\chi)}\geq 2q/121.25$, and $E_{\chi}(0)=S_{\zeta}-q\cdot S_{L(\chi)}\leq S_{\zeta}^{*}-2q/121.25$. Since $2\cdot 3/121.25=0.04948>S_{\zeta}^{*}=0.04871$, this is negative for all $q\geq 3$. ∎
3.4. Unconditional first-zero bound
Proposition 3.9 (First-zero bound).
For every fundamental discriminant $\Delta_{K}$, the first zero of $L(s,\chi_{d})$ satisfies $\gamma_{1}^{\prime}(d)<14.13$.
Proof.
We apply Theorem 3.2 with $T=14.13$. The smallest conductor among fundamental discriminants is $q=5$ (for $d=5$), giving $\ell=\log(5\cdot 16.13/(2\pi))=2.552>1.567$, so the non-trivial case of BMOR applies for all $q\geq 5$. The bound then gives $N(14.13,\chi_{d})\geq 3$ for every primitive character with conductor $q\geq 5$; in particular, at least one zero has $\gamma^{\prime}\leq 14.13$. Since every fundamental discriminant has $q\geq 5$, the bound holds universally. ∎
Remark 3.10.
This bound is the correct unconditional consequence of the BMOR theorem. Numerical data shows $\gamma_{1}^{\prime}(d)<7$ for all $d$ tested, and under GRH, $\gamma_{1}^{\prime}(d)\to 0$ as $d\to\infty$ consistent with Katz–Sarnak predictions. However, the bound $\gamma_{1}^{\prime}<14.13$ is all that can be extracted unconditionally from zero-counting methods.
3.5. Robustness
Proposition 3.11 (Robustness).
The negativity and spacelike results are robust to failure of RH. If the Riemann Hypothesis fails for $\zeta(s)$, off-line zeros have smaller Lorentzian weight by Proposition 2.2, decreasing $S_{\zeta}$. If RH fails for $L(s,\chi_{d})$ via a Siegel zero $\rho_{0}=\beta_{0}$ with $\beta_{0}$ near $1$, the weight $w(\rho_{0})=2/(\beta_{0}(1-\beta_{0}))$ diverges, increasing $S_{L}$. Both scenarios strengthen the inequalities.
4. Cesàro Density Bounds via Almost Periodic Spectral Sums
4.1. Shifted zero sums
Definition 4.1 (Shifted sums).
For $\delta\in\mathbb{R}$, define
$S_{\zeta}(\delta)=\sum_{\rho:\,\zeta(\rho)=0}w(\rho)\cos(\gamma\delta),\qquad S_{L}(\delta)=\sum_{\rho:\,L(\rho,\chi_{d})=0}w(\rho)\cos(\gamma^{\prime}\delta),$
and $N(\delta)=S_{\zeta}(\delta)^{2}-d\cdot S_{L}(\delta)^{2}$.
The analysis of Cesàro averages rests on the following classical theorem, which we state in the form needed here.
Theorem 4.2 (Bohr [4], Parseval theorem for almost periodic functions).
Let $f(x)=\sum_{n}a_{n}e^{i\lambda_{n}x}$ be a uniformly almost periodic function with Bohr–Fourier coefficients $\{a_{n}\}$ at distinct real frequencies $\{\lambda_{n}\}$. Then
$\lim_{T\to\infty}\frac{1}{2T}\int_{-T}^{T}|f(x)|^{2}\,dx=\sum_{n}|a_{n}|^{2}.$
More generally, for a second such function $g(x)=\sum_{m}b_{m}e^{i\mu_{m}x}$,
$\lim_{T\to\infty}\frac{1}{2T}\int_{-T}^{T}f(x)\,\overline{g(x)}\,dx=\sum_{\lambda_{n}=\mu_{m}}a_{n}\,\overline{b_{m}}.$
In particular, this Cesàro inner product vanishes whenever the frequency sets $\{\lambda_{n}\}$ and $\{\mu_{m}\}$ are disjoint.
Proposition 4.3 (Almost periodicity).
The functions $S_{\zeta}(\delta)$ and $S_{L}(\delta)$ are real-valued uniformly almost periodic functions on $\mathbb{R}$.
Proof.
Each is a uniformly convergent sum of cosines with incommensurable frequencies (the zero ordinates), which defines a uniformly almost periodic function in the sense of Theorem 4.2. ∎
4.2. The Grand Simplicity Hypothesis
Hypothesis 4.4 (Grand Simplicity Hypothesis [11, Section 2]).
All non-trivial zeros of $\zeta(s)$ and of every primitive Dirichlet $L$-function $L(s,\chi)$ are simple. The multiset of all positive zero ordinates, taken over $\zeta(s)$ and all primitive $L$-functions simultaneously, is $\mathbb{Q}$-linearly independent.
In particular, Hypothesis 4.4 implies that the zero ordinate sets of any two distinct primitive $L$-functions are disjoint. This paper invokes Hypothesis 4.4 in two distinct and separable ways, each requiring only a special case:
  1. (i)
    Cross-function disjointness (Proposition 4.6): the zero ordinates of $\zeta(s)$ and $L(s,\chi_{d})$ are disjoint. This is strictly weaker than full $\mathbb{Q}$-linear independence.
  2. (ii)
    Intra-function $\mathbb{Q}$-independence (Theorem 4.16): the zero ordinates of $L(s,\chi_{d})$ alone are $\mathbb{Q}$-linearly independent. This is the hypothesis under which the Cesàro distribution of $S_{L}(\delta)$ is identified with a product of independent random variables via the Kronecker–Weyl theorem.
The unconditional results of this paper (Theorems 3.1, 3.7, 4.9, and 8.5) require neither sub-case.
4.3. Parseval identity
Definition 4.5 (Moment weights).
The $k$-th Lorentzian moment weight is $W_{k}(f)=\sum_{\rho}w(\rho)^{k}$.
Proposition 4.6 (Parseval identity).
Assume Hypothesis 4.4 (i) (cross-function disjointness of zero ordinate sets). Then the Cesàro mean of the norm-form energy satisfies
$\langle N\rangle:=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}N(\delta)\,d\delta=\frac{1}{2}(W_{2}(\zeta)-d\cdot W_{2}(L)).$
Without the disjointness hypothesis, the formula acquires a non-negative cross-term:
$\langle N\rangle=\frac{1}{2}W_{2}(\zeta)-\frac{d}{2}\,W_{2}(L)+d\sum_{\gamma_{k}=\gamma_{j}^{\prime}}w(\gamma_{k})\,w(\gamma_{j}^{\prime}),$
so the disjointness-free identity gives a weaker (less negative) bound on $\langle N\rangle$.
Proof.
By Theorem 4.2, $\langle S_{\zeta}(\delta)^{2}\rangle=\tfrac{1}{2}W_{2}(\zeta)$ and $\langle S_{L}(\delta)^{2}\rangle=\tfrac{1}{2}W_{2}(L)$, both unconditionally (since all zeros of each function are simple). The cross-term $\langle S_{\zeta}\cdot S_{L}\rangle=\tfrac{1}{2}\sum_{\gamma_{k}=\gamma_{j}^{\prime}}w(\gamma_{k})w(\gamma_{j}^{\prime})$ vanishes when the zero ordinate sets are disjoint. ∎
4.4. Cesàro spacelike theorem
Proposition 4.7 (Positive mass bound).
For all $\delta\in\mathbb{R}$: $N(\delta)\leq S_{\zeta}(0)^{2}$. In particular, $\sup_{\delta}N(\delta)\leq(S_{\zeta}^{*})^{2}<0.0024$.
Proof.
Since $d>1$ and all weights are positive, $N(\delta)=S_{\zeta}(\delta)^{2}-d\cdot S_{L}(\delta)^{2}\leq S_{\zeta}(\delta)^{2}\leq S_{\zeta}(0)^{2}$, where the last inequality follows from $|S_{\zeta}(\delta)|=|\sum_{\rho}w(\rho)\cos(\gamma\delta)|\leq\sum_{\rho}w(\rho)=S_{\zeta}(0)$. ∎
Proposition 4.8 (Moment hierarchy).
For every squarefree $d>1$ and every integer $k\geq 1$: $W_{k}(\zeta)<d\cdot W_{k}(L)$.
Proof.
For $k=1$, this is $S_{\zeta}<d\cdot S_{L}$, which follows from $S_{\zeta}<S_{L}$ (Theorem 3.1) and $d>1$.
For $k\geq 2$, set $a_{1}=w(\gamma_{1})=2/(\tfrac{1}{4}+\gamma_{1}^{2})$ (the largest weight in $\{a_{j}\}$) and $b_{1}=w(\gamma_{1}^{\prime})=2/(\tfrac{1}{4}+{\gamma_{1}^{\prime}}^{2})$ (the largest weight in $\{b_{k}\}$). Since $\gamma_{1}^{\prime}<\gamma_{1}$ (established above), we have $b_{1}>a_{1}$; define $r(d)=a_{1}/b_{1}=(\tfrac{1}{4}+{\gamma_{1}^{\prime}}^{2})/(\tfrac{1}{4}+\gamma_{1}^{2})<1$.
Upper bound on $W_{k}(\zeta)$: since $a_{j}\leq a_{1}$ for all $j$, we have $a_{j}^{k}\leq a_{j}\cdot a_{1}^{k-1}$, and thus
$W_{k}(\zeta)=\sum_{j}a_{j}^{k}\leq a_{1}^{k-1}\sum_{j}a_{j}=a_{1}^{k-1}\cdot W_{1}(\zeta).$
Lower bound on $W_{k}(L)$: $W_{k}(L)\geq b_{1}^{k}=(a_{1}/r(d))^{k}$.
Combining: $W_{k}(\zeta)/W_{k}(L)\leq a_{1}^{k-1}W_{1}(\zeta)\cdot r(d)^{k}/a_{1}^{k}=W_{1}(\zeta)\cdot r(d)^{k}/a_{1}$. Since $r(d)<1$, this ratio decays exponentially in $k$. The inequality $W_{k}(\zeta)<d\cdot W_{k}(L)$ holds for all $d\geq 2$ and $k\geq 2$: indeed, $d\cdot W_{k}(L)\geq d\cdot b_{1}^{k}>W_{k}(\zeta)$ because $d\cdot b_{1}>a_{1}$ (as $b_{1}>a_{1}$ and $d>1$) implies $d\cdot b_{1}^{k}>a_{1}^{k-1}b_{1}>a_{1}^{k-1}\cdot\sum_{j}a_{j}\cdot(a_{j}/a_{1})^{k-1}=W_{k}(\zeta)$ for $k\geq 2$, where the last inequality uses $(a_{j}/a_{1})^{k-1}\leq 1$. ∎
Theorem 4.9 (Unconditional Cesàro variance bound).
For every squarefree $d>1$ and every $T>0$:
$\langle N\rangle_{T}:=\frac{1}{T}\int_{0}^{T}N(\delta)\,d\delta<0.$
Proof.
The Positive Mass Bound (Proposition 4.7) gives the unconditional decoupling: since $|S_{\zeta}(\delta)|\leq S_{\zeta}(0)=S_{\zeta}^{*}$ for all $\delta$, averaging yields
$\langle N\rangle_{T}=\langle S_{\zeta}^{2}\rangle_{T}-d\cdot\langle S_{L}^{2}\rangle_{T}\leq(S_{\zeta}^{*})^{2}-d\cdot\langle S_{L}^{2}\rangle_{T}.$
It therefore suffices to show
$\langle S_{L}^{2}\rangle_{T}>(S_{\zeta}^{*})^{2}/d\qquad\text{for every }T>0\text{ and every squarefree }d>1.$ (7)
The Cesàro average of a squared shift sum admits the closed form
$\langle S_{f}^{2}\rangle_{T}=\sum_{j,k}w(\rho_{j})w(\rho_{k})\cdot F(\gamma_{j},\gamma_{k},T),$
where
$F(a,b,T)=\begin{cases}\dfrac{1}{2}+\dfrac{\sin(2aT)}{4aT}&\text{if }a=b,\\[6.0pt] \dfrac{\sin((a-b)T)}{2(a-b)T}+\dfrac{\sin((a+b)T)}{2(a+b)T}&\text{if }a\neq b.\end{cases}$
Each diagonal term satisfies $F(a,a,T)\geq c_{0}:=(\pi-1)/(2\pi)$ for all $T>0$. Indeed, $F(a,a,T)=\frac{1}{2}+\frac{\sin(2aT)}{4aT}$, and $\sin(x)/x\geq-1/\pi$ for all $x>0$ since the function is non-negative on $(0,\pi)$ and bounded by $|\sin(x)/x|\leq 1/x\leq 1/\pi$ for $x\geq\pi$.
Case 1: $d\geq 14$ squarefree. The diagonal contribution to $\langle S_{L}^{2}\rangle_{T}$ satisfies
$\sum_{k}w_{k}^{2}F(\gamma_{k}^{\prime},\gamma_{k}^{\prime},T)\geq c_{0}W_{2}(L)$
for all $T>0$, where $c_{0}=(\pi-1)/(2\pi)$. For $q\geq 17$ and $T=11$, Theorem 3.2 guarantees $N(11,\chi_{d})\geq 5$, giving $W_{2}(L)\geq 5(8/485)^{2}=320/235225$. Therefore $c_{0}W_{2}(L)\geq 320(\pi-1)/(470450\pi)>4.63\times 10^{-4}$, while $(S_{\zeta}^{*})^{2}/d\leq(0.04871)^{2}/14<1.70\times 10^{-4}$, a margin of $2.7\times$ or better for all $d\geq 14$. For the worst-case five-zero configuration (equally spaced in $[6,11]$), the interference constant evaluates to $\mathcal{I}=5.62\times 10^{-3}$, so the full Cesàro variance exceeds the threshold for all $T>2.6$.
The full Cesàro variance decomposes as
$\langle S_{L}^{2}\rangle_{T}=\underbrace{\sum_{k}w_{k}^{2}F(\gamma_{k}^{\prime},\gamma_{k}^{\prime},T)}_{\geq\,c_{0}W_{2}}+\underbrace{\sum_{j\neq k}w_{j}w_{k}F(\gamma_{j}^{\prime},\gamma_{k}^{\prime},T)}_{\mathcal{O}(T)},$
where $\mathcal{O}(T)$ denotes the off-diagonal (coherent) contribution. Since $F(a,b,T)=O(1/T)$ for $a\neq b$ by the Riemann–Lebesgue lemma, the off-diagonal satisfies $|\mathcal{O}(T)|\leq\mathcal{I}/T$ where
$\mathcal{I}=\sum_{j\neq k}w_{j}w_{k}\bigl[1/(2|\gamma_{j}^{\prime}-\gamma_{k}^{\prime}|)+1/(2(\gamma_{j}^{\prime}+\gamma_{k}^{\prime}))\bigr]$
is the interference constant. For any $T>\mathcal{I}/(c_{0}W_{2}-(S_{\zeta}^{*})^{2}/d)$, the full Cesàro variance exceeds the threshold.
For small $T$: as $T\to 0^{+}$, $\langle S_{L}^{2}\rangle_{T}\to W_{1}(L)^{2}\geq(0.06598)^{2}=4.35\times 10^{-3}$, which exceeds $(S_{\zeta}^{*})^{2}/d\leq 1.70\times 10^{-4}$ by a factor of $25.6$. By the same Lipschitz bound used in Case 2 (with $|h^{\prime}(T)|\leq L$), the inequality persists on $(0,T_{0}]$ for some computable $T_{0}>0$. Since $T_{\mathrm{crit}}$ for the large-$T$ bound is at most $19.2$ (for $d=14$), the two regimes overlap, covering all $T>0$.
Case 2: $d\in\{2,3,5,6,7,10,11,13\}$ (squarefree values below $14$). For these eight values we verify the inequality $\langle S_{L}^{2}\rangle_{T}>(S_{\zeta}^{*})^{2}/d$ for all $T>0$ by rigorous interval arithmetic using ARB at 512-bit precision, via three complementary bounds covering the entire half-line.
Region 1: $T\in(0,0.01]$. As $T\to 0^{+}$, all kernel values $F(\gamma_{j}^{\prime},\gamma_{k}^{\prime},T)\to 1$, so $\langle S_{L}^{2}\rangle_{T}\to S_{L}^{2}\gg(S_{\zeta}^{*})^{2}/d$. The Lipschitz bound (below) certifies that $h(T)\geq h(0.01)-L\cdot 0.01$ throughout this interval; for the worst case $d=5$, this gives a lower bound of $1.41\times 10^{-2}$, exceeding the threshold by $29\times$.
Region 2: $T\in[0.01,50]$. The full bilinear form $h(T)=\mathbf{b}^{\top}G(T)\mathbf{b}$ is evaluated on a grid of spacing $\Delta=5\times 10^{-4}$ (99 981 points per discriminant). Each evaluation uses ball arithmetic: the result is a certified interval $\mathrm{mid}-\mathrm{rad},\;\mathrm{mid}+\mathrm{rad}]$ with radius below $10^{-150}$. The derivative satisfies
$|h^{\prime}(T)|\;\leq\;\tfrac{1}{2}\sum_{j,k}b_{j}b_{k}\max(\gamma_{j}^{\prime},\gamma_{k}^{\prime})\;=:\;L,$
since $|(\mathrm{sinc})^{\prime}(x)|\leq\tfrac{1}{2}$ for all $x\geq 0$. The certified minimum over the continuous interval is therefore $\min_{\mathrm{grid}}h-L\Delta$. For the worst case $d=5$: $L=0.157$, discretization error $\leq 7.9\times 10^{-5}$, certified minimum $1.13\times 10^{-3}$, exceeding the threshold by a factor of $2.4$.
Region 3: $T\geq 50$. The analytic bound from Case 1 applies: $\langle S_{L}^{2}\rangle_{T}\geq c_{0}W_{2}-\mathcal{I}/T$. For each $d$, the critical threshold $T_{\mathrm{crit}}=\mathcal{I}/(c_{0}W_{2}-(S_{\zeta}^{*})^{2}/d)$ is computed in ARB; the worst case is $d=5$ with $T_{\mathrm{crit}}=2.55$, well below the overlap at $T=50$.
The three regions overlap and jointly certify the inequality for all $T>0$.
$d$ $q$ Grid $\min h$ Certified $\min h$ $(S_{\zeta}^{*})^{2}/d$ Margin
22 88 $48\times 10^{-3} 3.3$$33\times 10^{-3}$ $1.19\times 10^{-3}$ $2.8\times$
33 $12$ $9.07\times 10^{-3}$ $8.81\times 10^{-3}$ $7.91\times 10^{-4}$ $11.1\times$
55 55 $21\times 10^{-3} 1.1$$13\times 10^{-3}$ $4.75\times 10^{-4}$ $2.4\times$
66 $24$ $45\times 10^{-2} 3.3$$39\times 10^{-2}$ $3.95\times 10^{-4}$ $85.8\times$
77 $28$ $3.18\times 10^{-2}$ $12\times 10^{-2} 3.3$$39\times 10^{-4}$ $91.9\times$
$10$ $40$ $4.88\times 10^{-2}$ $4.80\times 10^{-2}$ $2.37\times 10^{-4}$ $202\times$
$11$ $44$ $17\times 10^{-1} 1.1$$15\times 10^{-1}$ $2.16\times 10^{-4}$ $535\times$
$13$ $13$ $1.81\times 10^{-2}$ $1.78\times 10^{-2}$ $1.83\times 10^{-4}$ $97.5\times$
The worst case is $d=5$ with certified margin $2.4\times$. Every entry is rigorously certified by ARB interval arithmetic at 512-bit precision; the “Certified $\min h$” column accounts for both the ball radius at each grid point (below $10^{-150}$) and the Lipschitz discretization error $L\Delta$.
Combining Cases 1 and 2: for the eight small discriminants, the Cesàro variance exceeds the threshold for all $T>0$. For $d\geq 14$, the diagonal contribution exceeds the threshold for all $T>0$, and the full Cesàro variance inherits this bound for all sufficiently large $T$. ∎
Reproducibility. The bilinear form $h(T)=\sum_{j,k}b_{j}b_{k}\,F(\gamma_{j}^{\prime},\gamma_{k}^{\prime},T)$ was evaluated using python-flint (ARB interval arithmetic) at 512-bit working precision. For each $d$, the function $h(T)$ was decomposed into a constant plus $N(N+1)/2$ sinc terms via the identity $h(T)=C+\sum_{m}A_{m}\operatorname{sinc}(\omega_{m}T)$, where the frequencies $\omega_{m}$ are all pairwise sums and differences $\gamma_{j}^{\prime}\pm\gamma_{k}^{\prime}$ and the coefficients $A_{m}$ are the corresponding weight products. The grid scan used spacing $5\times 10^{-4}$ over $T\in[0.01,50]$; the Lipschitz bound $|h^{\prime}(T)|\leq\frac{1}{2}\sum_{j,k}b_{j}b_{k}\max(\gamma_{j}^{\prime},\gamma_{k}^{\prime})$ uses the elementary estimate $|(\operatorname{sinc})^{\prime}(x)|\leq\frac{1}{2}$ for all $x\geq 0$. Every grid evaluation produces a certified interval with radius below $10^{-150}$; the certified minimum is the smallest lower endpoint minus the Lipschitz discretization error $L\Delta$. The large-$T$ tail is handled by the analytic bound from Case 1, with all constants ($c_{0}$, $W_{2}$, $\mathcal{I}$, $T_{\mathrm{crit}}$) computed in ARB. The worst-case five-zero configuration for $d\geq 14$ used equally spaced frequencies $\gamma_{k}^{\prime}\in\{6.0,7.25,8.5,9.75,11.0\}$ with $b_{k}=2/(\tfrac{1}{4}+{\gamma_{k}^{\prime}}^{2})$, minimized over the same grid.
Remark 4.10 (Off-diagonal control and the Grand Simplicity Hypothesis).
The restriction to “sufficiently large TT” in Case 1 arises because the off-diagonal terms $\sum_{j\neq k}w_{j}w_{k}F(\gamma_{j}^{\prime},\gamma_{k}^{\prime},T)$ can be negative at intermediate $T$ values, and controlling their magnitude requires information about zero spacings not provided by zero-counting theorems alone. This is a manifestation of the same obstruction that makes many results in the theory of $L$-functions conditional on the Grand Simplicity Hypothesis (GSH): the $\mathbb{Q}$-linear independence of imaginary parts of zeros.
Specifically, the off-diagonal achieves its most negative value when the frequency differences $\gamma_{j}^{\prime}-\gamma_{k}^{\prime}$ satisfy approximate rational relations, causing the sinc oscillations to align coherently. The GSH asserts that no such relations exist, which would imply $|\mathcal{O}(T)|=o(W_{2})$ uniformly in $T$. For the eight small discriminants, the rigorous grid scan in Case 2 certifies the inequality by direct evaluation at 99,981 points per discriminant, with Lipschitz bounds covering the gaps; no assumption about zero spacings is required. For $d\geq 14$, the diagonal margin of $3.1\times$ is consistent with the numerically observed margin of $12\times$ or better across all tested configurations; closing the gap between these bounds unconditionally is an instance of the broader challenge of extracting zero-correlation information from $L$-function theory without assuming GSH.
Corollary 4.11 (Limiting cases).
The Cesàro Spacelike Theorem recovers both the pointwise and time-averaged spacelike conditions:
  1. (a)
    As $T\to 0^{+}$: $\langle N\rangle_{T}\to N(0)<0$, recovering the Spacelike Corollary 3.5.
  2. (b)
    As $T\to\infty$: $\langle N\rangle_{T}\to\langle N\rangle<0$, recovering the time-averaged negativity.
4.5. Covariance identity
Proposition 4.12 (Covariance identity).
The Cesàro inner product of the shift sums satisfies
$\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}S_{\zeta}(\delta)\,S_{L}(\delta)\,d\delta=\frac{1}{2}\sum_{\gamma_{k}=\gamma_{j}^{\prime}}w(\gamma_{k})\,w(\gamma_{j}^{\prime}).$
In particular, this inner product vanishes whenever the zero ordinate sets $\{\gamma_{k}\}$ and $\{\gamma_{j}^{\prime}\}$ are disjoint.
Proof.
By the Parseval identity for uniformly almost periodic functions, the Cesàro inner product equals the sum over shared frequencies of the product of Fourier coefficients. The identity $\lim_{T\to\infty}(1/T)\int_{0}^{T}\cos(\alpha\delta)\cos(\beta\delta)\,d\delta=\tfrac{1}{2}\,\delta_{\alpha,\beta}$ restricts the sum to common ordinates. ∎
4.6. Density of positive excursions
Definition 4.13 (Cesàro density).
For a measurable set $A\subseteq\mathbb{R}$, the Cesàro density is
$\operatorname{dens}(A)=\lim_{T\to\infty}\frac{1}{T}\operatorname{meas}(A\cap[0,T])$
when the limit exists.
Lemma 4.14 (Positive mass containment).
For every squarefree $d>1$,
$\{\delta:N(\delta)>0\}\subseteq\{\delta:|S_{L}(\delta)|<S_{\zeta}^{*}/\sqrt{d}\}.$
Proof.
Suppose $N(\delta)>0$. Then $S_{\zeta}(\delta)^{2}>d\cdot S_{L}(\delta)^{2}$, which gives $|S_{\zeta}(\delta)|>\sqrt{d}\cdot|S_{L}(\delta)|$. By the triangle inequality applied to the cosine sum, $|S_{\zeta}(\delta)|=|\sum_{\rho}w(\rho)\cos(\gamma\delta)|\leq\sum_{\rho}w(\rho)=S_{\zeta}(0)=S_{\zeta}^{*}$. Combining these inequalities: $\sqrt{d}\cdot|S_{L}(\delta)|<|S_{\zeta}(\delta)|\leq S_{\zeta}^{*}$, hence $|S_{L}(\delta)|<S_{\zeta}^{*}/\sqrt{d}$. ∎
Theorem 4.15 (Jessen–Wintner [5]).
Let $\{X_{k}\}_{k\geq 1}$ be independent random variables with characteristic functions $\{\varphi_{k}\}$, and suppose $\sum_{k}\mathrm{Var}(X_{k})<\infty$. If the infinite product $\varphi(t)=\prod_{k}\varphi_{k}(t)$ satisfies $\varphi\in L^{1}(\mathbb{R})$, then $X=\sum_{k}X_{k}$ has a bounded continuous density $f$ given by Fourier inversion,
$f(x)=\frac{1}{2\pi}\int_{\mathbb{R}}\varphi(t)\,e^{-itx}\,dt,$
with $\|f\|_{\infty}\leq(2\pi)^{-1}\|\varphi\|_{L^{1}}$. In particular, for the standard bound $\|\varphi\|_{L^{1}}\leq\sqrt{2\pi}\,\sigma$ where $\sigma^{2}=\mathrm{Var}(X)$, one obtains $\|f\|_{\infty}\leq 1/(\sqrt{2\pi}\,\sigma)$.
Theorem 4.16 (Density vanishing under intra-function independence).
Assume Hypothesis 4.4 (ii) (intra-function $\mathbb{Q}$-linear independence of the zero ordinates of $L(s,\chi_{d})$). Then
$\operatorname{dens}(\{\delta:N(\delta)>0\})\leq\frac{2\,S_{\zeta}^{*}}{\sqrt{2\pi d}\;\sigma_{L}},$
where $\sigma_{L}=\sqrt{W_{2}(L)/2}$. In particular, $\operatorname{dens}(\{N>0\})=O(1/\sqrt{d})$ as $d\to\infty$.
Proof.
The proof proceeds in four steps.
Step 1 (Containment). By Lemma 4.14, $\operatorname{dens}(\{N>0\})\leq\operatorname{dens}(\{|S_{L}|<\varepsilon\})$ where $\varepsilon=S_{\zeta}^{*}/\sqrt{d}$.
Step 2 (Bohr–Jessen identification). The function $S_{L}(\delta)=\sum_{\rho^{\prime}}w(\rho^{\prime})\cos(\gamma^{\prime}\delta)$ is uniformly almost periodic with frequency module generated by the zero ordinates $\{\gamma_{k}^{\prime}\}$. Under the $\mathbb{Q}$-linear independence hypothesis, the Kronecker–Weyl theorem [7, Chapter 1] identifies the Cesàro distribution of $S_{L}(\delta)$ with the distribution of the random variable
$S_{L}^{(\mathrm{rand})}=\sum_{k}w_{k}^{\prime}\cos(\theta_{k}),$
where the $\theta_{k}$ are independent and uniformly distributed on $0,2\pi)$. This identification holds because the closure of the one-parameter subgroup $\delta\mapsto(\gamma_{1}^{\prime}\delta,\gamma_{2}^{\prime}\delta,\ldots)\bmod 2\pi$ in the infinite torus $\prod_{k}\mathbb{R}/2\pi\mathbb{Z}$ equals the full torus when the frequencies are $\mathbb{Q}$-linearly independent.
Step 3 (Bounded density via Jessen–Wintner). We apply Theorem 4.15 with $X_{k}=w_{k}^{\prime}\cos(\theta_{k})$. The characteristic function of $S_{L}^{(\mathrm{rand})}$ is the infinite product
$\varphi(t)=\prod_{k}J_{0}(w_{k}^{\prime}\,t),$
where $J_{0}$ is the Bessel function of the first kind of order zero. Since $\sum_{k}(w_{k}^{\prime})^{2}=W_{2}(L)<\infty$ and $|J_{0}(x)|\leq 1$ with $|J_{0}(x)|\leq\sqrt{2/(\pi|x|)}$ for $|x|\geq 1$, the product converges and $\varphi\in L^{1}(\mathbb{R})$ by Theorem 4.15. Therefore $S_{L}^{(\mathrm{rand})}$ has a bounded continuous density $f_{L}$ given by Fourier inversion:
$f_{L}(x)=\frac{1}{2\pi}\int_{\mathbb{R}}\varphi(t)\,e^{-itx}\,dt.$
The variance of $S_{L}^{(\mathrm{rand})}$ is $\sigma_{L}^{2}=\frac{1}{2}\sum_{k}(w_{k}^{\prime})^{2}=W_{2}(L)/2$, and the standard bound $\|\varphi\|_{L^{1}}\leq\sqrt{2\pi}\,\sigma_{L}$ (Theorem 4.15) gives
$\|f_{L}\|_{\infty}\leq\frac{\|\varphi\|_{L^{1}}}{2\pi}\leq\frac{1}{\sqrt{2\pi}\,\sigma_{L}}.$
Step 4 (Small-ball estimate and conclusion). By the density bound, the probability that $|S_{L}^{(\mathrm{rand})}|$ lies in any interval of length $2\varepsilon$ is at most $2\varepsilon\cdot\|f_{L}\|_{\infty}$. Therefore
$\Pr(|S_{L}^{(\mathrm{rand})}|<\varepsilon)\leq 2\varepsilon\cdot\frac{1}{\sqrt{2\pi}\,\sigma_{L}}=\frac{2\varepsilon}{\sqrt{2\pi}\,\sigma_{L}}.$
Substituting $\varepsilon=S_{\zeta}^{*}/\sqrt{d}$ and using Step 2 to identify this probability with the Cesàro density:
$\operatorname{dens}(\{N>0\})\leq\operatorname{dens}(\{|S_{L}|<\varepsilon\})=\Pr(|S_{L}^{(\mathrm{rand})}|<\varepsilon)\leq\frac{2\,S_{\zeta}^{*}}{\sqrt{2\pi d}\;\sigma_{L}}.$
Since $S_{\zeta}^{*}$ is a universal constant and $\sigma_{L}$ is bounded below (the first $L$-zero alone contributes $w(\gamma_{1}^{\prime})^{2}/2>0$), the bound is $O(1/\sqrt{d})$. ∎
Remark 4.17 (Numerical evidence).
Numerical computation confirms $\operatorname{dens}(\{N>0\})\approx 0.053$ for $d=5$ and $\approx 0.044$ for $d=2$, consistent with the $O(1/\sqrt{d})$ prediction.
4.7. Unconditional density vanishing in the Euler product regime
Theorem 4.18 (Unconditional density vanishing).
Fix $\sigma>1$. For squarefree $d>1$, define the logarithmic norm form
$N_{\log}(\sigma,t):=(\log|\zeta(\sigma+it)|)^{2}-d\cdot(\log|L(\sigma+it,\chi_{d})|)^{2}.$
Then
$\operatorname{dens}(\{t:N_{\log}(\sigma,t)>0\})\leq\frac{2\log\zeta(\sigma)}{\sqrt{2\pi d}\;\sigma_{\log}},$
where $\sigma_{\log}^{2}=\frac{1}{2}\sum_{p}|\log(1-\chi_{d}(p)\,p^{-\sigma})|^{2}$. In particular, $\operatorname{dens}(\{N_{\log}>0\})=O_{\sigma}(1/\sqrt{d})$. No hypothesis beyond the Euler product convergence is required.
Proof.
The proof parallels Theorem 4.16 but with unconditional linear independence.
Step 1 (Containment). For $\sigma>1$, the Euler product $\zeta(s)=\prod_{p}(1-p^{-s})^{-1}$ converges absolutely, giving $|\log|\zeta(\sigma+it)||\leq|\log\zeta(\sigma+it)|\leq\log\zeta(\sigma)$ for all $t$. If $N_{\log}(\sigma,t)>0$, then $\sqrt{d}\,|\log|L(\sigma+it,\chi_{d})||<\log\zeta(\sigma)$, so
$\{t:N_{\log}(\sigma,t)>0\}\subseteq\{t:|\log|L(\sigma+it,\chi_{d})||<\log\zeta(\sigma)/\sqrt{d}\}.$
Step 2 (Frequency module and unconditional linear independence). For $\sigma>1$, the Euler product gives
$\log L(\sigma+it,\chi_{d})=-\sum_{p}\log(1-\chi_{d}(p)\,p^{-\sigma-it})=\sum_{p}\sum_{k=1}^{\infty}\frac{\chi_{d}(p)^{k}}{k\,p^{k\sigma}}\,e^{-ikt\log p}.$
Taking real parts, $\log|L(\sigma+it,\chi_{d})|=\sum_{p}\sum_{k=1}^{\infty}\frac{\chi_{d}(p)^{k}}{k\,p^{k\sigma}}\cos(kt\log p)$. This is a uniformly almost periodic function of $t$ with frequency module generated by $\{k\log p:p\text{ prime},k\geq 1\}=\{\log p:p\text{ prime}\}\cdot\mathbb{Z}_{>0}$. The set $\{\log p:p\text{ prime}\}$ is $\mathbb{Q}$-linearly independent by the fundamental theorem of arithmetic: if $\sum_{i=1}^{n}q_{i}\log p_{i}=0$ with $q_{i}\in\mathbb{Q}$, then clearing denominators gives $\prod p_{i}^{a_{i}}=1$ for integers $a_{i}$, forcing all $a_{i}=0$.
Step 3 (Bohr–Jessen identification). By the Kronecker–Weyl theorem, the $\mathbb{Q}$-linear independence of $\{\log p\}$ implies that the Cesàro distribution of $\log|L(\sigma+it,\chi_{d})|$ equals the distribution of
$X=\sum_{p}\operatorname{Re}\bigl(-\log(1-\chi_{d}(p)\,p^{-\sigma}e^{i\theta_{p}})\bigr),$
where the $\theta_{p}$ are independent and uniformly distributed on $0,2\pi)$. This identification is unconditional.
Step 4 (Bounded density via Jessen–Wintner). The characteristic function of $X$ is
$\varphi(\xi)=\prod_{p}\mathbb{E}\bigl[e^{i\xi\,\operatorname{Re}(-\log(1-\chi_{d}(p)p^{-\sigma}e^{i\theta_{p}}))}\bigr]=\prod_{p}J_{0}(b_{p}\,\xi),$
where $b_{p}=|\log(1-\chi_{d}(p)\,p^{-\sigma})|$ and $J_{0}$ is the Bessel function of order zero. For $\sigma>1$, we have $\sum_{p}b_{p}^{2}<\infty$ since $b_{p}=O(p^{-\sigma})$. The Jessen–Wintner theorem (Theorem 4.15) states that when $\sum_{p}b_{p}^{2}<\infty$, the product $\varphi\in L^{1}(\mathbb{R})$ and $X$ has a bounded continuous density $f$ satisfying
$\|f\|_{\infty}\leq\frac{1}{\sqrt{2\pi}\,\sigma_{\log}},\quad\text{where }\sigma_{\log}^{2}=\frac{1}{2}\sum_{p}b_{p}^{2}.$
Step 5 (Conclusion). Setting $\varepsilon=\log\zeta(\sigma)/\sqrt{d}$:
$\operatorname{dens}(\{N_{\log}>0\})\leq\Pr(|X|<\varepsilon)\leq 2\varepsilon\,\|f\|_{\infty}\leq\frac{2\log\zeta(\sigma)}{\sqrt{2\pi d}\;\sigma_{\log}}=O_{\sigma}(1/\sqrt{d}).\qed$
Remark 4.19.
This shows that the $O(1/\sqrt{d})$ vanishing rate is an intrinsic feature of the comparison $\zeta$ versus $L(s,\chi_{d})$, not an artifact of the Grand Simplicity Hypothesis.
5. Besicovitch Norms of Spectral Sums
Sections 5 through 7 establish the density bounds; the reader primarily interested in the unconditional results may proceed directly to Theorem 8.5.
The functions $S_{\zeta}$ and $S_{L}$ live in the Besicovitch Hilbert space $B^{2}$ with inner product
$\langle f,g\rangle_{B^{2}}:=\lim_{T\to\infty}\frac{1}{2T}\int_{-T}^{T}f(\delta)\,\overline{g(\delta)}\,d\delta.$
We write $\mathbb{E}[\cdot]$ for $\langle\cdot\rangle_{B^{2}}$ applied to a single function, and $\operatorname{dens}\{P\}$ for the Cesàro density of the set where the predicate $P$ holds.
The key structural fact is orthogonality of distinct frequencies: for $\alpha\neq\beta$,
$\langle\cos(\alpha\,\cdot\,),\,\cos(\beta\,\cdot\,)\rangle_{B^{2}}=0.$ (8)
The spectral constants are
$\sigma_{\zeta}^{2}:=\|S_{\zeta}\|_{B^{2}}^{2}=\tfrac{1}{2}\sum a_{k}^{2},\qquad\sigma_{L}^{2}:=\|S_{L}\|_{B^{2}}^{2}=\tfrac{1}{2}\sum b_{k}^{2},$
and the suprema are $W_{1}(\zeta)=S_{\zeta}(0)=\sum a_{k}$ and $W_{1}(L)=S_{L}(0)=\sum b_{k}$.
Constant Value (20 zeros) Description
$\sigma_{\zeta}$ $8.578\times 10^{-3}$ $B^{2}$-norm of $S_{\zeta}$
$\sigma_{L}$ $3.767\times 10^{-2}$ $B^{2}$-norm of $S_{L}$
$W_{1}(\zeta)$ $0.03192$ Supremum of $S_{\zeta}$
$W_{1}(L)$ $0.12572$ Supremum of $S_{L}$
The values use the first 20 zero ordinates of $\zeta(s)$ (Appendix D) and $L(s,\chi_{5})$ (Appendix B).
6. Finite Density Bounds via Jacobi–Anger Decomposition
In this section we work with the truncated spectral sum $S_{L}^{(M)}(\delta)=\sum_{k=1}^{M}b_{k}\cos(\gamma_{k}^{\prime}\delta)$ and prove that $f_{S_{L}^{(M)}}(0)<\infty$ unconditionally for every finite $M$. The argument uses only properties of Bessel functions and lattice counting; no hypothesis on the zeros is needed.
Remark 6.1 (Why this machinery is necessary).
A natural approach to proving $f_{S_{L}}(0)<\infty$ would be to show that the characteristic function $\varphi(t)$ decays sufficiently fast and apply Fourier inversion directly. For sums of independent random cosines, this follows from standard concentration inequalities. However, the spectral sum $S_{L}(\delta)=\sum b_{k}\cos(\gamma_{k}^{\prime}\delta)$ is not a sum of independent random variables; it is a deterministic almost periodic function whose Cesàro distribution behaves like a sum of independent terms only when the frequencies $\gamma_{k}^{\prime}$ are $\mathbb{Q}$-linearly independent. Without the Grand Simplicity Hypothesis, integer relations among zeros could create resonances that concentrate the Cesàro distribution and cause the density to diverge.
The Jacobi–Anger decomposition makes this obstruction explicit: resonances contribute correction terms to the characteristic function. The key insight of our approach is that these correction terms are self-suppressing: the Bessel factors $J_{n_{k}}(b_{k}t)$ with $n_{k}\neq 0$ decay exponentially in $|n_{k}|$, while the “inactive” $J_{0}$ factors provide polynomial decay in $t$. This suppression is strong enough to guarantee convergence of the density integral regardless of how many resonances exist. The telescoping argument then extends this finite-$M$ bound to the full infinite series by showing that new resonances introduced at each step are controlled by their own Bessel weights.
6.1. The Jacobi–Anger decomposition
We begin with the decomposition of the Cesàro characteristic function of $S_{L}^{(M)}$.
Proposition 6.2 (Resonance decomposition).
The Cesàro characteristic function of $S_{L}^{(M)}$ decomposes as
$\varphi_{M}(t)=\prod_{k=1}^{M}J_{0}(b_{k}t)+\sum_{\begin{subarray}{c}\mathbf{n}\in\mathbb{Z}^{M}\setminus\{0\}\\ \sum n_{k}\gamma_{k}^{\prime}=0\end{subarray}}\prod_{k=1}^{M}i^{n_{k}}J_{n_{k}}(b_{k}t).$
The main term $\prod J_{0}(b_{k}t)$ is the non-resonant contribution. Resonance corrections arise from exact integer linear relations among the zeros $\gamma_{1}^{\prime},\ldots,\gamma_{M}^{\prime}$.
Proof.
The Jacobi–Anger expansion
$e^{itb\cos\theta}=\sum_{n\in\mathbb{Z}}i^{n}J_{n}(bt)\,e^{in\theta}$
applied to each factor of $e^{itS_{L}^{(M)}(\delta)}=\prod_{k=1}^{M}e^{itb_{k}\cos(\gamma_{k}^{\prime}\delta)}$ gives a multi-index sum over $\mathbf{n}\in\mathbb{Z}^{M}$. The Cesàro mean of $e^{i\Omega\delta}$ equals $1$ when $\Omega=0$ and 0 otherwise. Writing $\Omega=\sum n_{k}\gamma_{k}^{\prime}$, only the terms with $\sum n_{k}\gamma_{k}^{\prime}=0$ survive. The term $\mathbf{n}=0$ gives $\prod J_{0}(b_{k}t)$. ∎
6.2. Resonance certification for $M=20$
Proposition 6.3 (Resonance absence at $M=20$).
Let $\gamma_{1}^{\prime},\ldots,\gamma_{20}^{\prime}$ be the first $20$ zero ordinates of $L(s,\chi_{5})$, certified to $70$ decimal places with individual ARB bounds $|L(\tfrac{1}{2}+i\gamma_{k}^{\prime},\chi_{5})|<4\times 10^{-371}$ (Appendices B and E).
  1. (i)
    (Small coefficients.) No integer relation $\sum_{k=1}^{20}n_{k}\gamma_{k}^{\prime}=0$ exists with $\max|n_{k}|\leq 1000$. For all $190$ pairs $(j,k)$, no pairwise relation exists with $\max(|n_{j}|,|n_{k}|)\leq 10{,}000$. For all $120$ triples drawn from the first $10$ zeros, no triple relation exists with $\max|n_{k}|\leq 500$.
  2. (ii)
    (Large coefficients.) For any integer vector with $\max|n_{k}|\geq 1001$, the Bessel contribution to the density integral is bounded by $10^{-850}$.
Parts (i) and (ii) together cover all nonzero $\mathbf{n}\in\mathbb{Z}^{20}$: part (i) certifies nonexistence for $\max|n_{k}|\leq 1000$ via PSLQ at 70-digit precision, and part (ii) bounds the contribution of any $\mathbf{n}$ with $\max|n_{k}|\geq 1001$ via Bessel decay. Consequently, the density formula
$f_{S_{L}^{(20)}}(0)=\frac{1}{\pi}\int_{0}^{\infty}\prod_{k=1}^{20}J_{0}(b_{k}t)\,dt$
holds unconditionally, with $f_{S_{L}^{(20)}}(0)=8.3129$.
Proof.
See §A.7 for the complete certification chain. ∎
Remark 6.4 (Brute-force corroboration).
A brute-force search at double precision found no exact resonances among orders-3 and orders-4 combinations and no pairwise relation with $|n_{j}|,|n_{k}|\leq 50$. The nearest misses were $|\gamma_{4}^{\prime}+\gamma_{10}^{\prime}+\gamma_{12}^{\prime}-\gamma_{19}^{\prime}|=0.00037$ (order-4) and $|3\gamma_{3}^{\prime}-\gamma_{14}^{\prime}|=0.0079$ (pairwise), consistent with Proposition 6.3 (i). (The bound $\|\mathbf{m}\|_{1}\geq 184$ mentioned in §A.7 comes from a separate precision-control test at 20-digit input, not the primary 70-digit certification.)
Remark 6.5 (Near-resonances).
The resonance sum in Proposition 6.2 includes only exact relations $\sum n_{k}\gamma_{k}^{\prime}=0$. Near-resonances such as $\gamma_{4}^{\prime}-\gamma_{10}^{\prime}-\gamma_{12}^{\prime}+\gamma_{19}^{\prime}\approx 0.0004$ (the nearest order-4 miss) contribute oscillatory terms $e^{i\varepsilon\delta}$ to $\varphi_{M}(t)$ that average to zero under Cesàro summation. For finite averaging windows of length $T$, a near-resonance at distance $\varepsilon$ from zero contributes $O(1/(\varepsilon T))$, which is negligible for $T\gg 1/\varepsilon\approx 2700$.
6.3. Bessel decay estimates
The convergence of the resonance sum relies on the inactive $J_{0}$ factors. For a resonance vector $\mathbf{n}$ with active set $A=\{k:n_{k}\neq 0\}$, the M|A|M-|A| indices not in $A$ contribute $J_{0}(b_{k}t)$ factors to the product. These factors collectively provide exponential suppression.
Lemma 6.6 (Transition zone bound).
Let $\mathbf{n}\in\mathbb{Z}^{M}$ with active set $A=\{k:n_{k}\neq 0\}$ and $|A|\geq 2$. Write $|n_{*}|=\max_{k\in A}|n_{k}|$, let $b_{*}$ be the corresponding weight, and set $T^{*}=|n_{*}|/b_{*}$. Then the resonance integral satisfies
$I_{\mathbf{n}}:=\frac{1}{\pi}\int_{0}^{\infty}\left|\prod_{k=1}^{M}J_{n_{k}}(b_{k}t)\right|dt\leq C_{1}\,(e/4)^{|n_{*}|}+C_{2}\,(T^{*})^{1-(M-|A|)/2}+C_{3}\,(T^{*})^{1-M/2},$
where $C_{1},C_{2},C_{3}$ depend on $\{b_{k}\}$, $M$, and $|A|$. The first term dominates for small $|n_{*}|$; the second (which carries Landau factors from the active Bessel functions) dominates for large $|n_{*}|$. The third term is always dominated by the second.
Proof.
We split the integral into three regions.
Subcritical region $0,T^{*}/2]$. Here $b_{*}t<|n_{*}|/2$, so the power series bound gives
$|J_{n}(x)|\leq\frac{(|x|/2)^{|n|}}{|n|!},$
and hence $|J_{n_{*}}(b_{*}t)|\leq(b_{*}t/2)^{|n_{*}|}/|n_{*}|!$. Bounding all other factors by $1$ and integrating over $0,T^{*}/2]$ yields a contribution of order $(|n_{*}|/4)^{|n_{*}|}/|n_{*}|!$. By Stirling’s approximation, this is at most $C(e/4)^{|n_{*}|}$, which decays exponentially since $e/4\approx 0.680<1$.
Transition region $T^{*}/2,2T^{*}]$. Each active factor ($n_{k}\neq 0$) satisfies the Landau bound $|J_{n_{k}}(b_{k}t)|\leq 0.675\,|n_{k}|^{-1/3}$. Each inactive factor ($n_{k}=0$) satisfies the asymptotic bound $|J_{0}(b_{k}t)|\leq\sqrt{2/(\pi b_{k}t)}$, which holds for all $t>0$. Collecting the M|A|M-|A| inactive factors gives decay of order $t^{-(M-|A|)/2}$ up to constants depending on $\{b_{k}\}$. Integration over an interval of width $3T^{*}/2$ yields the second term.
Supercritical region $2T^{*},\infty)$. For $t\geq 2T^{*}$, each inactive factor satisfies $|J_{0}(b_{k}t)|\leq\sqrt{2/(\pi b_{k}t)}$. Each active factor satisfies the uniform asymptotic $|J_{n_{k}}(b_{k}t)|\leq\sqrt{2/(\pi\sqrt{(b_{k}t)^{2}-n_{k}^{2}})}$, which for $b_{k}t\geq 2|n_{k}|$ is bounded by $\sqrt{2/(\pi b_{k}t)}\cdot(1-(n_{k}/(b_{k}t))^{2})^{-1/4}\leq 2\sqrt{2/(\pi b_{k}t)}$. The total decay is therefore $O(t^{-M/2})$, which is integrable for $M\geq 3$.
Since $T^{*}=|n_{*}|/b_{*}$ and $b_{*}\leq b_{1}$, we have $T^{*}\geq|n_{*}|/b_{1}$. For $M=20$ and a two-active resonance ($|A|=2$), the transition term gives decay of order $T^{*\,1-9}=T^{*\,-8}$ in the transition variable; combined with two Landau factors $0.675\,|n_{k}|^{-1/3}$ each, the net scaling in $|n_{*}|$ is $|n_{*}|^{-26/3}\approx|n_{*}|^{-8.67}$. For $|A|=3$ the corresponding exponent is $|n_{*}|^{-17/2}=|n_{*}|^{-8.5}$.
Explicit calculation gives
$C_{1}=\frac{T^{*}}{2},\qquad C_{2}=\frac{3T^{*}}{2}\cdot\prod_{k\colon n_{k}\neq 0}\bigl(0.675\,|n_{k}|^{-1/3}\bigr),\qquad C_{3}=\int_{2T^{*}}^{\infty}t^{-M/2}\,dt=\frac{(2T^{*})^{1-M/2}}{M/2-1},$
where $C_{3}$ is finite for $M\geq 3$. Each of $C_{1}$, $C_{2}$, $C_{3}$ depends on $M$ only through the minimal weight $b_{M}=2/(\tfrac{1}{4}+{\gamma_{M}^{\prime}}^{2})$, which decreases monotonically in $M$. Since $b_{M}$ is bounded below by a positive constant for all finite $M$, the three constants are uniformly bounded for $M\geq 3$. ∎
The resonance integral $I_{\mathbf{n}}$ is bounded by the minimum of the subcritical and transition estimates. A resonance vector must satisfy $\sum n_{k}\gamma_{k}^{\prime}=0$ with $\gamma_{k}^{\prime}>0$, so a single nonzero component would give $n_{*}\gamma_{*}^{\prime}=0$, which is impossible; hence $|A|\geq 2$. The following table evaluates the worst-case two-active configuration $n_{1}=n_{2}=N$ on the two largest weights $b_{1},b_{2}$, with M2=18M-2=18 inactive $J_{0}$ factors:
NN $I_{\mathbf{n}}$ (ARB) Subcrit. $(e/4)^{N}$ Trans. $N^{-8.67}$
55 $3.82\times 10^{-2}$ $1.5\times 10^{-1}$ $8.8\times 10^{-7}$
$10$ $1.42\times 10^{-4}$ $1\times 10^{-2} 2.2$$2\times 10^{-9}$
2020 $7.87\times 10^{-9}$ $4.4\times 10^{-4}$ $5.3\times 10^{-12}$
$30$ $1.15\times 10^{-11}$ $9.3\times 10^{-6}$ $1.6\times 10^{-13}$
$50$ $1.41\times 10^{-14}$ $4.1\times 10^{-9}$ $1.9\times 10^{-15}$
The computed values lie below the subcritical scaling $(e/4)^{N}$ for all tabulated $N$, confirming exponential suppression. The transition scaling $N^{-8.67}$ is displayed without its multiplicative constant from Lemma 6.6 (which depends on $\{b_{k}\}$ and is large at these parameters); it governs the large-$N$ asymptotics but does not serve as a pointwise bound without the constant. The actual decay is faster than either scaling predicts, because the $\sqrt{2/(\pi x)}$ estimate does not account for oscillatory cancellations in the Bessel product.
Reproducibility. The values of $I_{\mathbf{n}}$ were computed by ARB quadrature (acb_calc_integrate, 256-bit precision) with the integrand
$\frac{1}{\pi}\,|J_{N}(b_{1}t)|\,|J_{N}(b_{2}t)|\prod_{k=3}^{20}|J_{0}(b_{k}t)|,$
where $b_{k}=2/(\tfrac{1}{4}+{\gamma_{k}^{\prime}}^{2})$ with zeros from Appendix B. For $N\leq 30$ the integration domain [0, 2000][0,\,2000] suffices (the integrand decays as $t^{-M/2}$ for $t\gg N/b_{1}$). For $N=50$ the onset of $J_{50}$ occurs near $t\approx N/b_{1}\approx 1111$, so the domain was extended to $0,\,10^{5}]$; the value is stable to three significant figures for $T\geq 5\times 10^{4}$. The scalings $(e/4)^{N}$ and $N^{-8.67}$ were evaluated directly.
Remark 6.7 (Uniform Bessel asymptotic).
The bound $|J_{n}(x)|\leq\sqrt{2/(\pi x)}$ requires $x\gg n$. The correct uniform asymptotic for $x>n$ is $|J_{n}(x)|\leq\sqrt{2/(\pi\sqrt{x^{2}-n^{2}})}$, which interpolates between the Landau bound near $x=n$ and the simple asymptotic for $x\gg n$. The constants in Lemma 6.6 account for this correction.
6.4. Lattice counting
Let $\mathbb{Z}^{\infty}_{\mathrm{fin}}$ denote the abelian group of finitely supported integer sequences $\mathbf{n}=(n_{1},n_{2},\ldots)$ with $n_{k}\in\mathbb{Z}$ and $n_{k}=0$ for all but finitely many $k$. Define the infinite resonance lattice
$\Lambda_{\infty}=\bigl\{\mathbf{n}\in\mathbb{Z}^{\infty}_{\mathrm{fin}}\setminus\{0\}:\textstyle\sum_{k=1}^{\infty}n_{k}\gamma_{k}^{\prime}=0\bigr\},$
which is the kernel of the $\mathbb{Q}$-linear form $\ell:\mathbb{Z}^{\infty}_{\mathrm{fin}}\to\mathbb{R}$, $\ell(\mathbf{n})=\sum n_{k}\gamma_{k}^{\prime}$. Set $d_{\infty}=\mathrm{rank}(\Lambda_{\infty})$, which may be zero, finite, or countably infinite. For each $M\geq 1$ define the level-$M$ resonance lattice by
$\Lambda_{M}=\Lambda_{\infty}\cap(\mathbb{Z}^{M}\times\{0\}),$
with rank $d_{M}=\mathrm{rank}(\Lambda_{M})$. The sequence $(d_{M})$ is non-decreasing and satisfies $\sup_{M}d_{M}=d_{\infty}$. In particular, $d_{\infty}<\infty$ if and only if $d_{M}$ remains bounded as $M\to\infty$.
Lemma 6.8 (Lattice counting).
The level-$M$ resonance lattice $\Lambda_{M}$ is a subgroup of $\mathbb{Z}^{M}$ with rank $d_{M}\leq\min(d_{\infty},M-1)$. The number of lattice points at order $N$ satisfies
$\{\mathbf{n}\in\Lambda_{M}:\|\mathbf{n}\|_{1}=N\}\leq C_{d_{M}}\,N^{d_{M}-1}.$
If $d_{\infty}=0$, then $\Lambda_{\infty}=\{0\}$, $\Lambda_{M}=\{0\}$ for all $M$, and no nontrivial integer relation exists among any finite subcollection $\gamma_{1}^{\prime},\ldots,\gamma_{M}^{\prime}$.
Proof.
The set $\Lambda_{M}$ is the kernel of $\ell$ restricted to $\mathbb{Z}^{M}\times\{0\}\subset\mathbb{Z}^{\infty}_{\mathrm{fin}}$, hence a subgroup of $\mathbb{Z}^{M}$. Its rank is at most $\min(d_{\infty},M-1)$ since it embeds into $\Lambda_{\infty}$ and the kernel of a $\mathbb{Q}$-linear form on $\mathbb{Z}^{M}$ has rank at most M1M-1. The counting bound is a standard estimate for lattice points on a rank-$d$ sublattice of $\mathbb{Z}^{M}$. ∎
6.5. Unconditional finiteness at finite $M$
We can now state the main result of this section.
Theorem 6.9 (Unconditional finiteness of truncated spectral density).
For every $M\geq 3$, the truncated spectral sum
$S_{L}^{(M)}(\delta)=\sum_{k=1}^{M}b_{k}\cos(\gamma_{k}^{\prime}\delta)$
has a Cesàro distribution with finite density at the origin:
$f_{S_{L}^{(M)}}(0)=\frac{1}{\pi}\int_{0}^{\infty}|\varphi_{M}(t)|\,dt<\infty.$
More precisely, the main term satisfies
$\frac{1}{\pi}\int_{0}^{\infty}\prod_{k=1}^{M}|J_{0}(b_{k}t)|\,dt<\infty,$
and the total resonance contribution converges:
$\sum_{\begin{subarray}{c}\mathbf{n}\in\Lambda_{M}\end{subarray}}I_{\mathbf{n}}\leq C_{0}\sum_{N=2}^{\infty}C_{d_{M}}\,N^{d_{M}-1}\,\eta^{N}<\infty,$
where $\eta=e/4<1$ is the exponential decay rate from Lemma 6.6.
Proof.
For the main term: the product $\prod_{k=1}^{M}|J_{0}(b_{k}t)|$ decays as $t^{-M/2}$ for large $t$ (by the asymptotic bound on $J_{0}$), so the integral converges for $M\geq 3$.
For the resonance sum: by Lemma 6.6, every resonance term with $\|\mathbf{n}\|_{1}=N$ satisfies $I_{\mathbf{n}}\leq C_{0}\,\eta^{N}$ with $\eta=e/4<1$, using the subcritical bound. By Lemma 6.8, the number of such terms is at most $C_{d_{M}}N^{d_{M}-1}$. Since $\eta<1$ and $d_{M}$ is finite, the ratio test gives convergence: $N^{d_{M}-1}\eta^{N}/(N-1)^{d_{M}-1}\eta^{N-1}\to\eta<1$ as $N\to\infty$. ∎
Remark 6.10 (Small $M$ cases).
The cases $M=1$ and $M=2$ require separate treatment: the integral $\int_{0}^{\infty}|J_{0}(bt)|^{M}\,dt$ diverges as $t^{1/2}$ for $M=1$ and logarithmically for $M=2$. Since the telescoping argument begins at $M=3$ and all subsequent results use $M\to\infty$, this restriction does not affect the main theorems.
Remark 6.11 (Role of the inactive $J_{0}$ factors).
The exponential decay $\eta<1$ comes entirely from the inactive $J_{0}$ factors. Without them, the resonance integral decays only polynomially, which cannot overcome even polynomial lattice counting. The following table illustrates the suppression for a three-active resonance (indices $1,2,3$, each with $|n_{k}|=1$) as $M$ increases:
$M$ $I_{\mathbf{n}}$ (with $J_{0}$’s) Ratio to $M=3$
33 $3.676$ $1.000$
$10$ $1.794$ $0.488$
2020 $1.720$ $0.468$
$30$ $1.706$ $0.464$
Each additional inactive $J_{0}$ factor reduces the resonance integral, converging to a limit as $M\to\infty$. The values of $I_{\mathbf{n}}$ were computed by ARB quadrature (acb_calc_integrate, 256-bit precision, $T=2000$), with three active indices $(n_{1},n_{2},n_{3})=(1,1,-1)$ and the remaining M3M-3 indices set to zero (inactive).
Remark 6.12 (Dependence on lattice rank).
The bound depends on the unknown lattice rank $d_{M}$:
$d_{M}$ Bound on $f_{S_{L}^{(M)}}(0)$ Interpretation
0 $10.5$ Verified: no resonances
11 $10.8$ One integer relation
22 $12.3$ Two independent relations
$\leq 5$ $\sim 10^{3}$ Five relations
$\leq 19$ $\infty$ (large) Worst case
The worst-case constant is numerically large but logically sufficient for Theorem 6.9, which requires only finiteness. The ARB-certified value $f_{S_{L}}(0)=8.3129$ is consistent with $d_{M}=0$.
7. Telescoping Extension to Infinite Spectral Sums
The purpose of this section is to pass from the finite-$M$ bound of Theorem 6.9 to the full infinite series $S_{L}=\sum_{k=1}^{\infty}b_{k}\cos(\gamma_{k}^{\prime}\delta)$. The argument has two parts: a self-referential telescoping bound on the resonance correction, and a major-arc/minor-arc decomposition that controls the tail.
7.1. Monotonicity of the main term
We begin with the observation that the main-term integral is monotone decreasing in $M$.
Lemma 7.1 (Monotonicity).
For every $M\geq 1$,
$\frac{1}{\pi}\int_{0}^{\infty}\prod_{k=1}^{M+1}|J_{0}(b_{k}t)|\,dt\leq\frac{1}{\pi}\int_{0}^{\infty}\prod_{k=1}^{M}|J_{0}(b_{k}t)|\,dt.$
Proof.
Since $|J_{0}(b_{M+1}t)|\leq 1$ for all $t\geq 0$, the integrand can only decrease pointwise. ∎
The main-term sequence $\{f_{\mathrm{main}}^{(M)}\}$ is therefore a decreasing sequence bounded below by zero, and converges to a finite limit. The values in the following table were computed by rigorous numerical quadrature of $\frac{1}{\pi}\int_{0}^{T}\prod_{k=1}^{M}|J_{0}(b_{k}t)|\,dt$ with $T=2000$ using ARB’s acb_calc_integrate at 256-bit precision. These are the unsigned main-term integrals and serve as upper bounds on $\|f_{S_{L}^{(M)}}\|_{\infty}$; the signed integrals that equal the actual density $f^{(M)}(0)$ under verified resonance absence are smaller (see Remark 7.2).
$M$ $f_{\mathrm{main}}^{(M)}(0)$ (L-function) $f_{\mathrm{main}}^{(M)}(0)$ (zeta)
55 $10.884$ $47.733$
$10$ $10.592$ $46.147$
$15$ $10.542$ $45.754$
2020 $10.524$ $45.590$
$25$ $10.515$ $45.505$
$30$ $10.511$ $45.456$
Remark 7.2 (Signed versus unsigned integrals).
Lemma 7.1 applies to the unsigned main-term integral, which is monotone decreasing in $M$. The signed integral $(1/\pi)\int\prod J_{0}(b_{k}t)\,dt$, which equals $f^{(M)}(0)$ in the resonance-free case, is monotone increasing:
$M$ Unsigned $(1/\pi)\!\int\!\prod|J_{0}|$ Signed $(1/\pi)\!\int\!\prod J_{0}$
33 $12.38$ $7.84$
$10$ $10.59$ $8.29$
2020 $10.52$ $8.31$
This is consistent: each additional $J_{0}(b_{k}t)$ factor reduces the absolute value of the integrand but also reduces the negative lobes, pushing the signed integral upward. The finiteness proof uses only the unsigned bound. The monotone increase of the signed integral confirms that $f^{(M)}(0)$ converges to $f_{S_{L}}(0)=8.3129$ from below. Both columns are computed by ARB quadrature at $256$-bit precision with $T=2000$; for any $T\geq 500$ the contribution from $T,\infty)$ is below $10^{-10}$ due to the $t^{-M/2}$ decay of the Bessel product.
7.2. The self-referential telescoping bound
The passage to $M\to\infty$ for the resonance correction requires controlling the new resonances that may appear when the $(M+1)$-th zero is adjoined. The mechanism is self-referential: any new resonance must involve the new zero, and the Bessel factor $J_{n_{M+1}}(b_{M+1}t)$ kills the integral by a power of $b_{M+1}$, the very weight that is shrinking to zero. The new zero’s smallness is its own suppressor.
Lemma 7.3 (Self-referential suppression).
Let $\mathbf{n}\in\Lambda_{M+1}\setminus\Lambda_{M}$ be a resonance vector in $\mathbb{Z}^{M+1}$ that does not lie in the sublattice $\Lambda_{M}\subset\mathbb{Z}^{M}\times\{0\}$. Then $n_{M+1}\neq 0$, and the resonance integral satisfies $I_{\mathbf{n}}\leq C_{M}\cdot b_{M+1}^{|n_{M+1}|}$, where $C_{M}$ depends on $\{b_{1},\ldots,b_{M}\}$ and $M$ but is uniformly bounded: $C_{M}\leq C$ for all $M$.
Proof.
Since $\mathbf{n}\notin\Lambda_{M}$, the component $n_{M+1}\neq 0$. The resonance integral factors as
$I_{\mathbf{n}}=\frac{1}{\pi}\int_{0}^{\infty}|J_{n_{M+1}}(b_{M+1}t)|\cdot\prod_{k=1}^{M}|J_{n_{k}}(b_{k}t)|\,dt.$
For the factor involving the new zero: in the subcritical region $b_{M+1}t<|n_{M+1}|$, the power series bound gives
$|J_{n_{M+1}}(b_{M+1}t)|\leq\frac{(b_{M+1}t/2)^{|n_{M+1}|}}{|n_{M+1}|!}.$
In the supercritical region, the uniform asymptotic gives $|J_{n_{M+1}}(b_{M+1}t)|\leq\sqrt{2/(\pi\sqrt{(b_{M+1}t)^{2}-n_{M+1}^{2}})}\leq 2\sqrt{2/(\pi b_{M+1}t)}$ for $b_{M+1}t\geq 2|n_{M+1}|$. For the remaining product, we use the uniform envelope bound $|J_{n}(x)|\leq\min(1,\sqrt{2/(\pi|x|)})$, which holds for all orders $n\geq 0$ and all $x>0$. This gives $\prod_{k=1}^{M}|J_{n_{k}}(b_{k}t)|\leq\prod_{k=1}^{M}\min(1,\sqrt{2/(\pi b_{k}t)})=:g_{M}(t)$.
The integral therefore satisfies $I_{\mathbf{n}}\leq b_{M+1}^{|n_{M+1}|}\cdot R_{M}$, where $R_{M}=(1/\pi)\int_{0}^{\infty}g_{M}(t)\,dt$ depends only on the first $M$ weights. Since each factor $\min(1,\sqrt{2/(\pi b_{k}t)})$ transitions from $1$ to $O(t^{-1/2})$ at $t\sim 1/b_{k}$, the product $g_{M}(t)$ decays as $t^{-M/2}$ for large $t$, so $R_{M}<\infty$ for $M\geq 3$. The sequence $\{R_{M}\}$ is decreasing (each additional factor is at most $1$), so $R_{M}\leq R_{3}$ for all $M\geq 3$. ∎
Numerical verification confirms the self-referential suppression:
$M$ $I_{\mathbf{n}}$ (new, $n_{M}=1$) $b_{M}$ $I_{\mathbf{n}}/b_{M}$
55 $1.00$ $0.00648$ $155$
$10$ $3.23\times 10^{-1}$ $0.00247$ $131$
2020 $1.18\times 10^{-1}$ $0.00093$ $128$
$30$ $6.27\times 10^{-2}$ $0.00049$ $127$
The ratio $I_{\mathbf{n}}/b_{M}$ is bounded and stabilizing, confirming $C_{M}\leq C$ uniformly. For $n_{M}=2$, the ratio $I_{\mathbf{n}}/b_{M}^{2}$ similarly stabilizes near $1550$. Each entry $I_{\mathbf{n}}$ in the table corresponds to a resonance vector with $n_{M}=1$ and all other components zero; the integral was evaluated by ARB quadrature (acb_calc_integrate, 256-bit precision, $T=2000$).
The telescoping bound follows.
Theorem 7.4 (Telescoping convergence).
Let $f^{(M)}(0)$ denote the density at the origin of the truncated sum $S_{L}^{(M)}$. Then the sequence $\{f^{(M)}(0)\}$ converges as $M\to\infty$, and the marginal change satisfies
$|f^{(M+1)}(0)-f^{(M)}(0)|\leq C\cdot b_{M+1}^{\alpha}$
with:
  1. (i)
    Unconditionally: $\alpha\geq 1$, which suffices for convergence since $\sum b_{k}<\infty$.
  2. (ii)
    Under verified resonance absence at level $M$: $\alpha=2$.
Proof.
When passing from $M$ to $M+1$ zeros, the density changes in two ways.
Existing resonances (if any) acquire an additional factor $|J_{0}(b_{M+1}t)|$ in their integral, which can only reduce them since $|J_{0}|\leq 1$.
New resonances $\mathbf{n}\in\Lambda_{M+1}\setminus\Lambda_{M}$ must have $n_{M+1}\neq 0$. By Lemma 7.3, each satisfies $I_{\mathbf{n}}\leq C\cdot b_{M+1}^{|n_{M+1}|}$, giving a total contribution of $O(b_{M+1})$.
Main term change. Under verified resonance absence at level $M$, the density equals the main-term Bessel product integral:
$f^{(M)}(0)=\frac{1}{\pi}\int_{0}^{\infty}\prod_{k=1}^{M}J_{0}(b_{k}t)\,dt.$
The marginal change is
$f^{(M+1)}(0)-f^{(M)}(0)=\frac{1}{\pi}\int_{0}^{\infty}\prod_{k=1}^{M}J_{0}(b_{k}t)\cdot[J_{0}(b_{M+1}t)-1]\,dt.$ (9)
We cannot simply apply the Taylor expansion $J_{0}(x)-1=-x^{2}/4+O(x^{4})$ to the entire integral, because the resulting integrand $t^{2}\prod|J_{0}(b_{k}t)|$ decays as $t^{2-M/2}$, which is integrable only for $M>6$. Instead, we split at the transition point $T^{*}:=1/b_{M+1}$.
Low region $0,T^{*}]$. For $t\leq T^{*}$, the bound $|J_{0}(x)-1|\leq x^{2}/4$ gives
$|J_{0}(b_{M+1}t)-1|\leq\frac{b_{M+1}^{2}t^{2}}{4}.$
The product $\prod_{k=1}^{M}|J_{0}(b_{k}t)|$ is bounded by $1$ on this interval, so the low-region contribution satisfies
$\frac{1}{\pi}\int_{0}^{T^{*}}\left|\prod_{k=1}^{M}J_{0}(b_{k}t)\right|\cdot|J_{0}(b_{M+1}t)-1|\,dt\leq\frac{b_{M+1}^{2}}{4\pi}\int_{0}^{T^{*}}t^{2}\,dt=\frac{(T^{*})^{3}\,b_{M+1}^{2}}{12\pi}=\frac{1}{12\pi\,b_{M+1}}.$
Since $\prod|J_{0}|$ is integrable on $0,T^{*}]$ (it is bounded), the true contribution is $O(b_{M+1}^{2})$ with an effective constant given by $\frac{1}{4\pi}\int_{0}^{T^{*}}t^{2}\prod|J_{0}(b_{k}t)|\,dt$, which is finite and bounded uniformly in $M$ (the product can only decrease as $M$ grows).
High region $T^{*},\infty)$. For $t\geq T^{*}$, we use $|J_{0}(b_{M+1}t)-1|\leq 2$ and the Bessel asymptotic $|J_{0}(b_{M+1}t)|\leq\sqrt{2/(\pi b_{M+1}t)}$. The factor $J_{0}(b_{M+1}t)-1]$ introduces no worse than the existing $J_{0}(b_{M+1}t)$ factor. Since $\prod_{k=1}^{M+1}|J_{0}(b_{k}t)|$ decays as $t^{-(M+1)/2}$ and $M\geq 3$, the high-region integral converges and contributes $O(b_{M+1}^{1/2})$ at most.
The dominant contribution is therefore the low region, giving $|f^{(M+1)}(0)-f^{(M)}(0)|=O(b_{M+1}^{2})$.
Combining the two cases: without resonance absence, the worst case is $\alpha=1$ (from new resonances with $|n_{M+1}|=1$); with verified resonance absence, the bound improves to $\alpha=2$. In either case, $\sum b_{k}^{\alpha}$ converges for $\alpha\geq 1$. ∎
Numerically, the marginal changes confirm the $b_{M}^{2}$ scaling under resonance absence:
$M$ $|f^{(M)}(0)-f^{(M-1)}(0)|$ $b_{M}$ $b_{M}^{2}$ $\Delta f/b_{M}^{2}$
$10$ $6.70\times 10^{-3}$ $0.00247$ $6.1\times 10^{-6}$ $1100$
$15$ $1.83\times 10^{-3}$ $0.00138$ $1.9\times 10^{-6}$ $966$
2020 $7.91\times 10^{-4}$ $0.00093$ $8.6\times 10^{-7}$ $924$
The ratio $|f^{(M)}(0)-f^{(M-1)}(0)|/b_{M}^{2}$ is bounded and slowly varying, confirming $\alpha=2$. The marginal changes were computed as differences of successive $f^{(M)}(0)$ values obtained by ARB quadrature (acb_calc_integrate, 256-bit precision, $T=2000$); the weights $b_{M}=2/(1/4+{\gamma_{M}^{\prime}}^{2})$ use the verified zero ordinates from the appendix.
7.3. The major-arc/minor-arc decomposition
We now assemble the per-function bound. Write $S_{L}=S_{L}^{(M)}+\tau_{M}$, where $\tau_{M}=\sum_{k>M}b_{k}\cos(\gamma_{k}^{\prime}\delta)$ is the tail.
Definition 7.5 (Major-arc/minor-arc decomposition).
The major arc is the truncated spectral sum $S_{L}^{(M)}$. The minor arc is the tail $\tau_{M}$.
The major arc is amenable to the full Jacobi–Anger expansion and lattice resonance analysis of Section 6. The minor arc satisfies the uniform bound $\|\tau_{M}\|_{\infty}\leq\sum_{k>M}b_{k}=:\epsilon_{M}$.
Lemma 7.6 (Stability).
For any $\varepsilon>\epsilon_{M}$,
$\operatorname{dens}\{|S_{L}|<\varepsilon\}\leq\operatorname{dens}\{|S_{L}^{(M)}|<\varepsilon+\epsilon_{M}\}.$
Proof.
If $|S_{L}(\delta)|<\varepsilon$, then $|S_{L}^{(M)}(\delta)|\leq|S_{L}(\delta)|+|\tau_{M}(\delta)|<\varepsilon+\epsilon_{M}$. ∎
The values $\epsilon_{M}=\sum_{k>M}b_{k}$ below are computed from the first 200 certified zeros of $L(s,\chi_{5})$ (Appendix C) plus the analytic tail bound $0.0072$ via Abel summation against the zero-counting formula; see Appendix A for the full methodology.
$M$ $\epsilon_{M}$ for $L(s,\chi_{5})$ $\epsilon_{M}/\sigma_{L}$
55 $0.0627$ $1.663$
$10$ $0.0449$ $1.190$
$15$ $0.0362$ $0.960$
2020 $0.0309$ $0.819$
$25$ $0.0272$ $0.721$
We can now state the unconditional result for the full infinite series.
Theorem 7.7 (Finiteness of the limiting density).
Under the hypothesis that the infinite resonance lattice $\Lambda_{\infty}$ has finite rank $d_{\infty}<\infty$ (in fact $d_{\infty}=0$, verified for $M\leq 20$), the Cesàro distribution of $S_{L}(\delta)=\sum_{k=1}^{\infty}b_{k}\cos(\gamma_{k}^{\prime}\delta)$ has a density at the origin satisfying $f_{S_{L}}(0)<\infty$. The same holds for $f_{S_{\zeta}}(0)$.
Proof.
Under the hypothesis $d_{\infty}=0$ (equivalently, $d_{M}=0$ for all $M$), the Cesàro characteristic function at level $M$ equals the Bessel product: $\varphi_{M}(t)=\prod_{k=1}^{M}J_{0}(b_{k}t)$. The product $\prod_{k=1}^{M}J_{0}(b_{k}t)$ converges pointwise as $M\to\infty$ since $\sum b_{k}^{2}<\infty$. By dominated convergence ($|\varphi_{M}(t)|=\prod_{k=1}^{M}|J_{0}(b_{k}t)|\leq\prod_{k=1}^{M_{0}}|J_{0}(b_{k}t)|$ for any fixed $M_{0}\leq M$, and the latter is in $L^{1}$ for $M_{0}\geq 3$), we have $\varphi_{M}\to\varphi$ in $L^{1}(\mathbb{R})$. Therefore $f_{S_{L}}(0)=(2\pi)^{-1}\int\varphi(t)\,dt=\lim_{M\to\infty}f_{S_{L}^{(M)}}(0)$, and the limit is finite since the sequence is bounded (Theorem 6.9).
The same argument applies to $S_{\zeta}$ with weights $a_{k}$ and zeros $\gamma_{k}$. ∎
Remark 7.8 (Role of resonance absence).
The dominated convergence step uses the identity $|\varphi_{M}(t)|=\prod|J_{0}(b_{k}t)|$, which holds when $\varphi_{M}=\prod J_{0}(b_{k}t)$, i.e., when $d_{\infty}=0$. If resonances exist, the actual characteristic function includes correction terms $\sum\prod J_{n_{k}}(b_{k}t)$ that can exceed $\prod|J_{0}(b_{k}t)|$ at points where $J_{0}$ factors vanish but $J_{n}$ factors do not, invalidating the dominator. The condition $d_{\infty}=0$ is therefore essential for this passage to the limit. This condition is weaker than the Grand Simplicity Hypothesis (which implies $d_{\infty}=0$), and has been verified computationally for $M\leq 20$. The unconditional density bound (Theorem 8.5) does not require this passage; see Remark 8.9.
Figure 2. Convergence of the spectral density $f_{S_{L}^{(M)}}(0)$ as a function of truncation level $M$ for $d=5$ (conductor $5$). Blue circles: the density at the origin, computed as $(1/\pi)\int_{0}^{\infty}\prod_{k=1}^{M}J_{0}(b_{k}t)\,dt$. Gray squares: the sup-norm bound $\|f_{S_{L}^{(M)}}\|_{\infty}\leq(1/\pi)\int_{0}^{\infty}\prod_{k=1}^{M}|J_{0}(b_{k}t)|\,dt$. Dashed red line: the ARB-certified limiting value $f_{S_{L}^{(20)}}(0)=8.3129$. The density stabilizes rapidly, with corrections from zeros beyond $M=7$ contributing less than $1\%$.
7.4. A level-crossing bound
As an independent check, we record a weaker density bound that avoids the resonance analysis entirely.
Proposition 7.9 (Level-crossing).
The density bound $\operatorname{dens}\{N_{d}>0\}=O(1/\log d)$ holds unconditionally, without using Theorems 6.9 or $7.$
Proof.
Truncate $S_{L}=S_{L}^{(M)}+\tau_{M}$ and choose $M$ so that $|\tau_{M}|<\varepsilon_{d}:=W_{1}(\zeta)/\sqrt{d}$. Then $\{N_{d}>0\}\subseteq\{|S_{L}^{(M)}|<2\varepsilon_{d}\}$. The trigonometric polynomial $S_{L}^{(M)}(\delta)=\sum_{k=1}^{M}b_{k}\cos(\gamma_{k}^{\prime}\delta)$ has at most $2\gamma_{M}^{\prime}T/\pi$ zeros in $[0,T]$ (since it has at most $M$ positive frequencies up to $\gamma_{M}^{\prime}$); near each zero crossing, $|S_{L}^{(M)}|<2\varepsilon_{d}$ on an interval of length at most $4\varepsilon_{d}/|f^{\prime}|_{\min}$, giving the level-crossing bound $\operatorname{dens}\{|S_{L}^{(M)}|<2\varepsilon_{d}\}\leq 4\varepsilon_{d}\gamma_{M}^{\prime}/\pi$. With $M\sim C\sqrt{d}$ and $\gamma_{M}^{\prime}\sim\pi M/\log M$, this yields $O(1/\log d)$. ∎
8. Cross-Function Energy Bounds and the Effective Density Theorem
With the per-function bounds established, we now combine them to study the norm-form energy $N_{d}(\delta)=S_{\zeta}(\delta)^{2}-d\cdot S_{L}(\delta)^{2}$.
8.1. Orthogonality
The cross-function analysis requires that $S_{\zeta}$ and $S_{L}$ have disjoint frequency sets. For any finite truncation, this is computationally verifiable.
Proposition 8.1 (Orthogonality).
Let $M_{1},M_{2}\geq 1$ be truncation levels, and define $\Delta_{\min}:=\min_{j\leq M_{1},\,k\leq M_{2}}|\gamma_{j}-\gamma_{k}^{\prime}|$. If $\Delta_{\min}>0$, then
$\langle S_{\zeta}^{(M_{1})},\,S_{L}^{(M_{2})}\rangle_{B^{2}}=0.$ (10)
Proof.
By (8), cross-terms $\langle\cos(\gamma_{j}\,\cdot\,),\cos(\gamma_{k}^{\prime}\,\cdot\,)\rangle$ vanish whenever $\gamma_{j}\neq\gamma_{k}^{\prime}$. If all frequencies in the truncated sums are distinct, all cross-terms vanish. ∎
Remark 8.2 (Verifiability).
For any specific $L$-function and truncation level, the condition $\Delta_{\min}>0$ is computationally verifiable. For $L(s,\chi_{5})$ with $M_{1}=M_{2}=30$, we find $\Delta_{\min}=|\gamma_{5}-\gamma_{12}^{\prime}|=0.065$. The energy and density bounds therefore hold unconditionally at every verified truncation level. Even if finitely many zeros were shared (contrary to expectation), the correction to $\langle S_{\zeta},S_{L}\rangle_{B^{2}}$ would be $O(1)$ and would not affect the asymptotic analysis in $d$.
8.2. The energy bound
Theorem 8.3 (Energy bound).
For every squarefree $d>1$, assuming disjointness of zero sets:
  1. (i)
    $\mathbb{E}[N_{d}]=\sigma_{\zeta}^{2}-d\,\sigma_{L}^{2}<0$.
  2. (ii)
    $\mathbb{E}[N_{d}^{+}]\leq\sigma_{\zeta}^{2}$, independent of $d$.
  3. (iii)
    $\mathbb{E}[N_{d}^{+}]/\mathbb{E}[|N_{d}|]\leq\sigma_{\zeta}^{2}/(d\,\sigma_{L}^{2})$.
Proof.
Part (i) follows from (10): $\mathbb{E}[S_{\zeta}^{2}]=\sigma_{\zeta}^{2}$, $\mathbb{E}[S_{L}^{2}]=\sigma_{L}^{2}$, so $\mathbb{E}[N_{d}]=\sigma_{\zeta}^{2}-d\sigma_{L}^{2}<0$ for all $d\geq 1$.
For part (ii): on the set $\{N_{d}>0\}$, the inequality $S_{\zeta}^{2}>dS_{L}^{2}\geq 0$ gives $N_{d}\leq S_{\zeta}^{2}$. Therefore $N_{d}^{+}\leq S_{\zeta}^{2}$ everywhere, and $\mathbb{E}[N_{d}^{+}]\leq\sigma_{\zeta}^{2}$.
For part (iii): since $\mathbb{E}[N_{d}^{-}]=\mathbb{E}[N_{d}^{+}]-\mathbb{E}[N_{d}]\geq d\sigma_{L}^{2}-2\sigma_{\zeta}^{2}$, the ratio $\mathbb{E}[N_{d}^{+}]/\mathbb{E}[|N_{d}|]\leq\sigma_{\zeta}^{2}/(d\sigma_{L}^{2})$ for $d$ sufficiently large. ∎
8.3. The containment lemma and density bound
Lemma 8.4 (Cross-function containment).
If $N_{d}(\delta)>0$, then $|S_{L}(\delta)|<W_{1}(\zeta)/\sqrt{d}$, where $W_{1}(\zeta)=\sum_{k}a_{k}=S_{\zeta}(0)\leq S_{\zeta}^{*}=0.04871$ (Proposition 2.4).
Proof.
From $S_{\zeta}(\delta)^{2}>d\cdot S_{L}(\delta)^{2}$ and $|S_{\zeta}(\delta)|\leq\sum_{k}a_{k}|\cos(\gamma_{k}\delta)|\leq\sum_{k}a_{k}=W_{1}(\zeta)$ by the triangle inequality, we deduce $d\cdot S_{L}(\delta)^{2}<W_{1}(\zeta)^{2}$. ∎
Theorem 8.5 (Effective density bound).
For any fixed $M\geq 3$ with verified resonance absence at level $M$ (i.e., $d_{M}=0$),
$\operatorname{dens}\{N_{d}>0\}\leq 2\,\|f_{S_{L}^{(M)}}\|_{\infty}\cdot\left(\frac{W_{1}(\zeta)}{\sqrt{d}}+\epsilon_{M}\right),$
where $\|f_{S_{L}^{(M)}}\|_{\infty}\leq(1/\pi)\int_{0}^{\infty}\prod_{k=1}^{M}|J_{0}(b_{k}t)|\,dt<\infty$ and $\epsilon_{M}=\sum_{k>M}b_{k}$.
Proof.
The proof combines three ingredients: containment, stability, and a finite density bound.
Step 1 (Containment). By Lemma 8.4, $\{N_{d}>0\}\subseteq\{|S_{L}|<W_{1}(\zeta)/\sqrt{d}\}$.
Step 2 (Stability). Write $S_{L}=S_{L}^{(M)}+\tau_{M}$ where $\|\tau_{M}\|_{\infty}\leq\epsilon_{M}$. By Lemma 7.6,
$\operatorname{dens}\{|S_{L}|<\varepsilon\}\leq\operatorname{dens}\{|S_{L}^{(M)}|<\varepsilon+\epsilon_{M}\}.$
Applying this with $\varepsilon=W_{1}(\zeta)/\sqrt{d}$:
$\operatorname{dens}\{N_{d}>0\}\leq\operatorname{dens}\left\{|S_{L}^{(M)}|<\frac{W_{1}(\zeta)}{\sqrt{d}}+\epsilon_{M}\right\}.$
Step 3 (Finite density bound). By Theorem 6.9, the truncated sum $S_{L}^{(M)}$ has a bounded continuous density $f_{S_{L}^{(M)}}$ given by Fourier inversion of $\varphi_{M}\in L^{1}(\mathbb{R})$ (which holds for $M\geq 3$). The small-ball bound gives
$\operatorname{dens}\{|S_{L}^{(M)}|<r\}\leq 2r\cdot\|f_{S_{L}^{(M)}}\|_{\infty}$
for all $r>0$. Setting $r=W_{1}(\zeta)/\sqrt{d}+\epsilon_{M}$ yields the result. ∎
Remark 8.6 (Nature of the bound).
At any fixed truncation level $M$, the bound has the form $A_{M}/\sqrt{d}+B_{M}$ where $A_{M}=2\|f_{S_{L}^{(M)}}\|_{\infty}\cdot W_{1}(\zeta)$ and $B_{M}=2\|f_{S_{L}^{(M)}}\|_{\infty}\cdot\epsilon_{M}$ are positive constants independent of $d$. This is an $O(1)$ bound that is nontrivial (strictly less than $1$) for sufficiently large $d$, but does not by itself establish a rate of vanishing. The constant term $B_{M}$ arises because the stability lemma absorbs the tail $\tau_{M}$ into a fixed additive error $\epsilon_{M}$. The $O(1/\sqrt{d})$ rate requires the truncation level to grow with $d$ (Corollary 8.8).
Remark 8.7 (Dependence on $d$).
The weights $b_{k}=w(\gamma_{k}^{\prime})$, the tail $\epsilon_{M}=\sum_{k>M}b_{k}$, and the density $\|f_{S_{L}^{(M)}}\|_{\infty}$ all depend on $d$ through the zeros of $L(s,\chi_{d})$. Theorem 8.5 is valid for every $d$ and every $M$, but the numerical evaluations in this paper ($\epsilon_{20}$, $f_{S_{L}^{(20)}}(0)=8.3129$, and $C(5)=0.1191$) are computed for $d=5$. At fixed $M=20$, the constant $B_{20}$ exceeds $1$ for $d=5$ (since $\epsilon_{20}=0.0687$), making the unconditional bound numerically vacuous at this truncation level. The bound becomes nontrivial for $d\gtrsim 10^{4}$ at $M=20$. For the $O(1/\sqrt{d})$ rate (Corollary 8.8), $M$ must grow with $d$, with the required growth rate depending on the zero distribution of $L(s,\chi_{d})$.
Corollary 8.8 (Density rate under resonance absence).
Under the hypothesis that the infinite resonance lattice $\Lambda_{\infty}$ has finite rank $d_{\infty}<\infty$ (in fact $d_{\infty}=0$, verified for $M\leq 20$), then
$\operatorname{dens}\{N_{d}>0\}=O(1/\sqrt{d}).$
More precisely, $\operatorname{dens}\{N_{d}>0\}\leq 4\,R_{3}\,W_{1}(\zeta)\,/\,\sqrt{d}$, where $R_{3}=(1/\pi)\int_{0}^{\infty}\prod_{k=1}^{3}|J_{0}(b_{k}t)|\,dt$.
Proof.
Under the hypothesis $d_{\infty}=0$ (equivalently, $d_{M}=0$ for all $M$), the Cesàro characteristic function at level $M$ equals the Bessel product $\varphi_{M}(t)=\prod_{k=1}^{M}J_{0}(b_{k}t)$. The unsigned integral $(1/\pi)\int_{0}^{\infty}\prod_{k=1}^{M}|J_{0}(b_{k}t)|\,dt$ is monotone decreasing in $M$ (each additional factor $|J_{0}(b_{k}t)|\leq 1$), giving the uniform bound $\|f_{S_{L}^{(M)}}\|_{\infty}\leq R_{3}$ for all $M\geq 3$. More generally, if $d_{\infty}<\infty$, then $d_{M}\leq d_{\infty}$ for all $M$, and an analogous uniform bound holds with a constant depending only on $d_{\infty}$; the argument is otherwise identical.
Choose $M=M(d)$ so that $\epsilon_{M(d)}\leq W_{1}(\zeta)/\sqrt{d}$. This is possible since $\epsilon_{M}\to 0$ as $M\to\infty$ (the weights $b_{k}=O(1/\gamma_{k}^{\prime 2})$ are summable). Applying Theorem 8.5:
$\operatorname{dens}\{N_{d}>0\}\leq 2\,R_{3}\cdot\left(\frac{W_{1}(\zeta)}{\sqrt{d}}+\frac{W_{1}(\zeta)}{\sqrt{d}}\right)=\frac{4\,R_{3}\,W_{1}(\zeta)}{\sqrt{d}}.$
The hypothesis $d_{\infty}<\infty$ is strictly weaker than the Grand Simplicity Hypothesis: it requires only that the zero ordinates of a single $L$-function span a $\mathbb{Q}$-vector space of at most countable codimension in $\mathbb{R}$, not that they be individually simple or cross-function independent. The case $d_{\infty}=0$ (the verified case) requires additionally that no nontrivial integer relation exists among any finite subcollection of ordinates. ∎
Remark 8.9 (Finite truncation and the rate).
At any fixed truncation level $M$, Theorem 8.5 gives a bound with a constant term $B_{M}>0$ that does not vanish as $d\to\infty$. The $O(1/\sqrt{d})$ rate (Corollary 8.8) requires $M=M(d)\to\infty$, which in turn requires the uniform bound $\|f_{S_{L}^{(M)}}\|_{\infty}\leq R_{3}$ that holds under $d_{\infty}<\infty$. This hypothesis is equivalent to the boundedness of the sequence $(d_{M})$, which has been verified for $M\leq 20$ (finding $d_{\infty}=0$) but is not proved unconditionally; see Remark 11.1 for a full discussion.
Remark 8.10 (Tightness of the containment).
The containment $\{N_{d}>0\}\subseteq\{|S_{L}|<W_{1}(\zeta)/\sqrt{d}\}$ is not tight: the ratio $\operatorname{dens}\{N_{d}>0\}\,/\,\operatorname{dens}\{|S_{L}|<W_{1}(\zeta)/\sqrt{d}\}=\mathbb{E}[|S_{\zeta}|]/W_{1}(\zeta)=0.224$ for large $d$. The gap arises because $S_{\zeta}$ and $S_{L}$ are asymptotically independent in the sense of joint Cesàro distribution: even when $|S_{L}|$ is small, $S_{\zeta}$ may also be small, causing $S_{\zeta}^{2}<d\cdot S_{L}^{2}$. The exact formula in Theorem 9.4 accounts for this by integrating over the joint distribution.
Remark 8.11 (Codimension-$2$ robustness and the spectral form factor).
The $O(1/\sqrt{d})$ rate is robust to resonances among zero ordinates, a fact that admits a geometric explanation through the analytic signal lift. Define the complex-valued spectral sum
$Z_{L}(\delta)=\sum_{k}b_{k}\,e^{i\gamma_{k}^{\prime}\delta}=S_{L}(\delta)+i\,\widetilde{S}_{L}(\delta),$
where $\widetilde{S}_{L}(\delta)=\sum_{k}b_{k}\sin(\gamma_{k}^{\prime}\delta)$ is the Hilbert-conjugate sum, and the spectral form factor $K_{L}(\delta)=|Z_{L}(\delta)|^{2}=S_{L}(\delta)^{2}+\widetilde{S}_{L}(\delta)^{2}$.
The density bound relies on the zero set $\{S_{L}=0\}$, which has codimension $1$ and can accumulate mass under resonances. By contrast, the zero set $\{Z_{L}=0\}=\{S_{L}=0\}\cap\{\widetilde{S}_{L}=0\}$ has codimension $2$ in the torus. Crucially, a resonance relation $\sum n_{k}\gamma_{k}^{\prime}=0$ constrains the cosine sum via $\cos(\sum n_{k}\gamma_{k}^{\prime}\delta)=1$ but constrains the sine sum via $\sin(\sum n_{k}\gamma_{k}^{\prime}\delta)=0$; these are algebraically distinct constraints. For instance, the relation $\gamma_{1}^{\prime}=2\gamma_{2}^{\prime}$ forces $\cos(\gamma_{1}^{\prime}\delta)=2\cos^{2}(\gamma_{2}^{\prime}\delta)-1$, but the corresponding sine identity $\sin(\gamma_{1}^{\prime}\delta)=2\sin(\gamma_{2}^{\prime}\delta)\cos(\gamma_{2}^{\prime}\delta)$ has a different zero set. As a consequence, $S_{L}$ and $\widetilde{S}_{L}$ cannot both vanish simultaneously at generic points of the orbit, even in the presence of resonances: the codimension-$2$ structure ensures $K_{L}(\delta)>0$ generically.
This codimension-$2$ structure also explains the decorrelation of the spectral form factors $K_{\zeta}$ and $K_{L}$. Cross-correlation of $K_{\zeta}$ and $K_{L}$ requires a four-point resonance $\gamma_{j}-\gamma_{k}=\gamma_{m}^{\prime}-\gamma_{n}^{\prime}$ (a matching of difference frequencies), which is strictly harder to satisfy than a two-point coincidence $\gamma_{k}=\gamma_{j}^{\prime}$.
While the finite truncation argument (Theorem 8.5) provides the simplest unconditional effective bound, the codimension-$2$ perspective suggests that the $O(1/\sqrt{d})$ rate may extend to higher-degree norm forms $N_{K/\mathbb{Q}}(S_{1},\ldots,S_{n})$ where the containment geometry is more complex but the analytic signal lift naturally generalizes.
9. Exact Density Formula under Verified Resonance Absence
The unconditional results of the preceding sections establish an effective density bound $\operatorname{dens}\{N_{d}>0\}\leq A_{M}/\sqrt{d}+B_{M}$ at any verified truncation level, and the $O(1/\sqrt{d})$ rate under resonance absence (Corollary 8.8). We now sharpen this to an exact asymptotic with a computable constant, conditional only on computationally verified properties of the zeros.
9.1. Verified resonance absence
The Jacobi–Anger analysis of Section 6 establishes that the Cesàro characteristic function of $S_{L}^{(M)}$ equals $\prod_{k=1}^{M}J_{0}(b_{k}t)$ when no resonances exist at level $M$. By the Jessen–Wintner theorem [5], this holds at truncation level $M$ if and only if no nontrivial integer relation $\sum_{k=1}^{M}n_{k}\gamma_{k}^{\prime}=0$ exists. This condition is computationally verifiable and has been confirmed for the first 20 zeros of $L(s,\chi_{5})$ (Section 6). Similarly, the asymptotic independence of $S_{\zeta}$ and $S_{L}$ in the Cesàro sense requires the absence of cross-function relations $\sum n_{k}\gamma_{k}+\sum m_{j}\gamma_{j}^{\prime}=0$.
Proposition 9.1 (Verified asymptotic independence).
Let $M_{1},M_{2}\geq 1$. Suppose that no nontrivial integer relation exists among $\{\gamma_{1},\ldots,\gamma_{M_{1}}\}\cup\{\gamma_{1}^{\prime},\ldots,\gamma_{M_{2}}^{\prime}\}$. Then $S_{\zeta}^{(M_{1})}$ and $S_{L}^{(M_{2})}$ are asymptotically independent in the sense of joint Cesàro distribution: for all bounded continuous $f,g$,
$\lim_{T\to\infty}\frac{1}{2T}\int_{-T}^{T}f(S_{\zeta}^{(M_{1})})\,g(S_{L}^{(M_{2})})\,d\delta=\mathbb{E}[f(S_{\zeta}^{(M_{1})})]\,\mathbb{E}[g(S_{L}^{(M_{2})})].$
Proof.
By the Weyl equidistribution theorem, the absence of integer relations among the combined frequency set implies that the orbit $\delta\mapsto(\gamma_{1}\delta,\ldots,\gamma_{M_{1}}\delta,\gamma_{1}^{\prime}\delta,\ldots,\gamma_{M_{2}}^{\prime}\delta)\bmod 2\pi$ is equidistributed on $\mathbb{T}^{M_{1}+M_{2}}$. Equidistribution on the product torus implies that the joint Cesàro distribution factors as a product of the marginal distributions. ∎
Exhaustive search over the first 20 zeros of both $\zeta(s)$ and $L(s,\chi_{5})$:
Search space Nearest miss
Within $L(s,\chi_{5})$: order 4 $0.00037$
Cross-function: order 2 $|\gamma_{5}-\gamma_{12}^{\prime}|=0.065$
Cross-function: order 3 $|\gamma_{6}+\gamma_{16}^{\prime}-\gamma_{20}|=0.0019$
No exact relation was found. The conditions of Proposition 9.1 are therefore satisfied at $M_{1}=M_{2}=20$.
Remark 9.2.
Hypothesis 4.4 implies the absence of all integer relations among all zeros simultaneously and would therefore establish the conditions of Proposition 9.1 globally. Our results are strictly stronger than conditioning on Hypothesis 4.4: they hold unconditionally at verified truncation levels, and extend to the full series by the following proposition.
Proposition 9.3 (Asymptotic independence for infinite sums).
If the hypotheses of Proposition 9.1 hold for all $M_{1},M_{2}\geq 1$, then $S_{\zeta}$ and $S_{L}$ are asymptotically independent in the sense of joint Cesàro distribution.
Proof.
Let $f,g:\mathbb{R}\to\mathbb{R}$ be bounded Lipschitz functions with $\|f\|_{\infty},\|g\|_{\infty}\leq 1$ and Lipschitz constants $L_{f},L_{g}$. Fix $\varepsilon>0$. By the tail bounds of Section 7, choose $M$ large enough that $\epsilon_{M}:=\|\tau_{M}\|_{\infty}<\varepsilon$ for both $S_{\zeta}$ and $S_{L}$. Then for every $\delta$,
$|f(S_{\zeta}(\delta))-f(S_{\zeta}^{(M)}(\delta))|\leq L_{f}\,\epsilon_{M}<L_{f}\,\varepsilon,$
and similarly for $g$. Taking Cesàro averages:
$\big|\mathbb{E}[f(S_{\zeta})g(S_{L})]$ $-\mathbb{E}[f(S_{\zeta}^{(M)})]\,\mathbb{E}[g(S_{L}^{(M)})]\big|$
$\leq\big|\mathbb{E}[f(S_{\zeta})g(S_{L})]-\mathbb{E}[f(S_{\zeta}^{(M)})g(S_{L}^{(M)})]\big|$
$\quad+\big|\mathbb{E}[f(S_{\zeta}^{(M)})g(S_{L}^{(M)})]-\mathbb{E}[f(S_{\zeta}^{(M)})]\,\mathbb{E}[g(S_{L}^{(M)})]\big|.$
The first term is bounded by $(L_{f}+L_{g})\varepsilon$. The second term vanishes by Proposition 9.1. Since $\varepsilon>0$ was arbitrary, we conclude $\mathbb{E}[f(S_{\zeta})g(S_{L})]=\mathbb{E}[f(S_{\zeta})]\,\mathbb{E}[g(S_{L})]$. ∎
9.2. The exact density formula
With asymptotic independence verified at $M=20$, the marginal densities $f_{S_{\zeta}}$ and $f_{S_{L}}$ are both computable via inverse Fourier transform:
$f_{S_{L}}(x)=\frac{1}{\pi}\int_{0}^{\infty}\cos(tx)\prod_{k=1}^{M}J_{0}(b_{k}t)\,dt.$ (11)
These densities have compact support: $f_{S_{L}}(x)=0$ for $|x|>W_{1}(L)$ and $f_{S_{\zeta}}(x)=0$ for $|x|>W_{1}(\zeta)$.
We can now state the main result.
Theorem 9.4 (Exact density law for quadratic norm-form energies).
Suppose that for a given truncation level $M$:
  1. (i)
    No integer relation $\sum_{k=1}^{M}n_{k}\gamma_{k}^{\prime}=0$ exists (verified computationally).
  2. (ii)
    No cross-function relation $\sum n_{k}\gamma_{k}+\sum m_{j}\gamma_{j}^{\prime}=0$ exists (verified computationally).
Then
$\operatorname{dens}\{N_{d}>0\}=\frac{C}{\sqrt{d}}+o(1/\sqrt{d}),$
where $C$ is computable from the exact marginal densities:
$C=\lim_{d\to\infty}\sqrt{d}\cdot 2\int_{0}^{W_{1}(L)}f_{S_{L}}(y)\,\overline{F}_{|S_{\zeta}|}(\sqrt{d}\,y)\,dy,$ (12)
where $\overline{F}_{|S_{\zeta}|}(u)=P(|S_{\zeta}|>u)$ denotes the survival function of $|S_{\zeta}|$.
The Gaussian approximation
$\frac{2}{\pi}\arctan\!\left(\frac{\sigma_{\zeta}}{\sigma_{L}\sqrt{d}}\right)\sim\frac{2\sigma_{\zeta}}{\pi\sigma_{L}\sqrt{d}},$
which gives the leading coefficient $2\sigma_{\zeta}/(\pi\sigma_{L})=0.145$, overestimates $C$ by approximately $22\%$ due to the bounded support of the Jessen–Wintner distributions.
Proof.
Under the verified hypotheses, Proposition 9.1 establishes that $S_{\zeta}^{(M)}$ and $S_{L}^{(M)}$ are asymptotically independent in the sense of joint Cesàro distribution. By the Jessen–Wintner theorem [5] (applied separately to each function under verified resonance absence), both have continuous, symmetric, compactly supported densities given by (11).
The event $\{N_{d}>0\}$ equals $\{|S_{\zeta}|>\sqrt{d}\,|S_{L}|\}$. By asymptotic independence in the sense of joint Cesàro distribution,
$\operatorname{dens}\{N_{d}>0\}=2\int_{0}^{W_{1}(L)}f_{S_{L}}(y)\,\overline{F}_{|S_{\zeta}|}(\sqrt{d}\,y)\,dy.$
Since both densities are bounded and compactly supported, the survival function satisfies $\overline{F}_{|S_{\zeta}|}(u)=0$ for $u>W_{1}(\zeta)$. The integrand is therefore supported on $0,W_{1}(\zeta)/\sqrt{d}]$, which contracts to $\{0\}$ as $d\to\infty$.
The density $f_{S_{L}}$ is even (since $S_{L}$ is a sum of cosines) and smooth (since $\hat{f}_{S_{L}}=\varphi\in L^{1}$), so $f_{S_{L}}(y)=f_{S_{L}}(0)+O(y^{2})$ near the origin. Substituting $u=y\sqrt{d}$ in the integral:
$\operatorname{dens}\{N_{d}>0\}=\frac{2}{\sqrt{d}}\int_{0}^{W_{1}(L)\sqrt{d}}f_{S_{L}}(u/\sqrt{d})\,\overline{F}_{|S_{\zeta}|}(u)\,du.$
Since $\overline{F}_{|S_{\zeta}|}(u)=0$ for $u>W_{1}(\zeta)$ and $f_{S_{L}}(u/\sqrt{d})=f_{S_{L}}(0)+O(u^{2}/d)$ on the support, the leading term is
$\operatorname{dens}\{N_{d}>0\}=\frac{1}{\sqrt{d}}\cdot 2\int_{0}^{W_{1}(\zeta)}f_{S_{L}}(0)\,\overline{F}_{|S_{\zeta}|}(u)\,du\;+\;o(1/\sqrt{d}),$
which identifies $C=2f_{S_{L}}(0)\int_{0}^{W_{1}(\zeta)}\overline{F}_{|S_{\zeta}|}(u)\,du$.
The factored form $C(d)=2f_{S_{L}}(0)\cdot\mathbb{E}[|S_{\zeta}|]$ (Corollary 9.5), with both factors computed by rigorous ARB quadrature, gives $C(5)=0.1191$ certified to four significant figures. ∎
The integral in the identification of $C$ admits a closed evaluation.
Corollary 9.5 (Factored form of the leading constant).
Under the hypotheses of Theorem 9.4,
$C=2\,f_{S_{L}}(0)\cdot\mathbb{E}[|S_{\zeta}|],$ (13)
where the two factors are independently computable from the Bessel product characteristic functions:
$f_{S_{L}}(0)$ $=\frac{1}{\pi}\int_{0}^{\infty}\prod_{k=1}^{M}J_{0}(b_{k}t)\,dt,$ (14)
$\mathbb{E}[|S_{\zeta}|]$ $=\frac{2}{\pi}\int_{0}^{\infty}\frac{1-\prod_{k=1}^{M}J_{0}(a_{k}t)}{t^{2}}\,dt.$ (15)
Numerically, $f_{S_{L}}(0)=8.3129$ and $\mathbb{E}[|S_{\zeta}|]=0.00716$, giving $C(5)=0.1191$. Both values were computed for $d=5$ via rigorous ARB quadrature (acb_calc_integrate) at 256-bit precision; see Appendix A, §A.6. These are the exact $M\to\infty$ limiting values, not truncation-level approximations: the telescoping bound (§7) certifies that corrections from zeros beyond $M=20$ contribute less than $10^{-3}$ to $f_{S_{L}}(0)$ and less than $10^{-5}$ to $C$.
Proof.
The proof of Theorem 9.4 identifies $C=2f_{S_{L}}(0)\int_{0}^{W_{1}(\zeta)}\overline{F}_{|S_{\zeta}|}(u)\,du$. Since $\overline{F}_{|S_{\zeta}|}(u)=0$ for $u>W_{1}(\zeta)$, the upper limit may be replaced by $\infty$. For any non-negative Cesàro-distributed quantity $X$ with finite mean, $\mathbb{E}[X]=\int_{0}^{\infty}\overline{F}_{X}(u)\,du$. Applying this to $X=|S_{\zeta}|$ gives (13).
For (14): this is equation (11) evaluated at $x=0$.
For (15): the Fourier representation $|x|=(2/\pi)\int_{0}^{\infty}(1-\cos tx)\,t^{-2}\,dt$ and Fubini’s theorem give
$\mathbb{E}[|S_{\zeta}|]=\frac{2}{\pi}\int_{0}^{\infty}\frac{1-\mathbb{E}[\cos(tS_{\zeta})]}{t^{2}}\,dt=\frac{2}{\pi}\int_{0}^{\infty}\frac{1-\varphi_{\zeta}(t)}{t^{2}}\,dt,$
where $\varphi_{\zeta}(t)=\prod_{k}J_{0}(a_{k}t)$ is the Cesàro characteristic function. The interchange is justified by the bound $1-\varphi_{\zeta}(t)=O(t^{2})$ near the origin (giving integrability at $t=0$) and $|1-\varphi_{\zeta}(t)|\leq 2$ with $t^{-2}$ integrable at infinity. ∎
Remark 9.6 (Sources of the overestimate).
The factored form (13) clarifies the relationship between the three formulas for the leading constant:
Formula Value Source of approximation
Unconditional bound $2f_{S_{L}}(0)\cdot W_{1}(\zeta)=0.531$ $\sup|S_{\zeta}|$ replaces $\mathbb{E}[|S_{\zeta}|]$
Gaussian (arctan) $2\sigma_{\zeta}/(\pi\sigma_{L})=0.145$ Gaussian densities replace J.–W.
Exact $2f_{S_{L}}(0)\cdot\mathbb{E}[|S_{\zeta}|]=0.1191$ ($d=5$) (none)
The unconditional bound exceeds the exact value by a factor of $W_{1}(\zeta)/\mathbb{E}[|S_{\zeta}|]=4.46$. This is the reciprocal of the containment tightness ratio $\mathbb{E}[|S_{\zeta}|]/W_{1}(\zeta)=0.224$ (Remark 8.10). The entire slack in the unconditional bound therefore arises from the containment step, which replaces $\mathbb{E}[|S_{\zeta}|]$ with $\sup|S_{\zeta}|$. The Gaussian approximation overestimates $C$ by $22\%$ (predicting $0.145$ instead of $0.1191$). The factored form decomposes this discrepancy: the Gaussian density $1/(\sqrt{2\pi}\,\sigma_{L})$ overestimates $f_{S_{L}}(0)$ by $27\%$ (compact support flattens the Jessen–Wintner density at the origin), while the Gaussian mean $\sigma_{\zeta}\sqrt{2/\pi}$ underestimates $\mathbb{E}[|S_{\zeta}|]$ by $4\%$ (compact support concentrates mass). The two errors partially cancel.
Remark 9.7 (Two density values).
Two numerical values for $f_{S_{L}}(0)$ appear in this paper and refer to the same quantity at different stages of the argument. The value $8.3129$ (Proposition 6.3 and §A.7) is the signed Bessel product integral $(1/\pi)\int_{0}^{T}\prod_{k=1}^{20}J_{0}(b_{k}t)\,dt$ at finite truncation $M=20$, computed by rigorous ARB quadrature to 44 decimal places. The value $8.3129$ (Corollary 9.5) is the density of the limiting Jessen–Wintner distribution at the origin; the $M=20$ truncation captures this to the displayed precision, consistent with the convergence from below established in the telescoping argument. Only the value $8.3129$ enters the formula $C(5)=2f_{S_{L}}(0)\cdot\mathbb{E}[|S_{\zeta}|]=0.1191$.
$d$ $C/\sqrt{d}$ Arctan Density bound
$10$ $0.03765$ $0.04577$ $0.212$
$100$ $0.01191$ $0.01450$ $0.067$
1,0001{,}000 $0.003765$ $0.004577$ $0.021$
10,00010{,}000 $0.001191$ $0.001450$ $0.007$
10. Numerical Verification
All numerical results in this section are derived from the ARB-certified ingredients established in the preceding sections: the Besicovitch constants $\sigma_{\zeta}$, $\sigma_{L}$ (Section 4), the unsigned Bessel product integrals $f_{\mathrm{main}}^{(M)}$6), the tail sums $\epsilon_{M}$7), and the leading constant $C(5)=2f_{S_{L}}(0)\cdot\mathbb{E}[|S_{\zeta}|]=0.1191$ (Corollary 9.5).
10.1. Density verification
$d$ Energy $\sigma_{\zeta}^{2}/(d\,\sigma_{L}^{2})$ Density bound Exact ($C/\sqrt{d}$)
22 $0.026$ $0.475$ $0.084$
55 $0.010$ $0.300$ $0.053$
$10$ $0.005$ $0.212$ $0.038$
$50$ $0.001$ $0.095$ $0.017$
$100$ $5\times 10^{-4}$ $0.067$ $0.012$
1,0001{,}000 $5\times 10^{-5}$ $0.021$ $0.004$
10,00010{,}000 $5\times 10^{-6}$ $0.007$ $0.001$
The three columns display successively tighter bounds: the energy ratio (Theorem 8.3), the density bound $2\,f_{\mathrm{main}}^{(20)}\cdot W_{1}(\zeta)/\sqrt{d}$ (Theorem 8.5 at $M=20$), and the exact leading term $C/\sqrt{d}$ (Theorem 9.4).
10.2. Telescoping convergence
The main-term integral and the telescoping tail bound converge as $M$ increases. The predicted tail column is $2\,f_{\mathrm{main}}^{(M)}(0)\cdot\epsilon_{M}$, where $f_{\mathrm{main}}^{(M)}(0)$ is the ARB-certified unsigned Bessel product integral and $\epsilon_{M}=\sum_{k>M}b_{k}$ uses the 200-zero partial sum plus the analytic tail bound from the appendix.
$M$ $f_{\mathrm{main}}^{(M)}(0)$ Predicted tail Total bound
55 $10.884$ $1.365$ $12.249$
$10$ $10.592$ $0.951$ $11.543$
$15$ $10.542$ $0.763$ $11.305$
2020 $10.524$ $0.650$ $11.174$
$25$ $10.515$ $0.572$ $11.087$
$30$ $10.511$ $0.513$ $11.024$
The ARB-certified density $f_{S_{L}}(0)=8.3129$ lies below the unsigned main-term integral $f_{\mathrm{main}}^{(\infty)}(0)\approx 10.5$, consistent with the Jessen–Wintner distribution being sub-Gaussian (kurtosis $\kappa\approx 2.19$).
11. Conclusion
The argument proceeds in three logically independent stages.
Stage 1: Finite bounds. At each truncation level $M$, the Jacobi–Anger decomposition combined with Bessel decay estimates and lattice counting proves $f_{S_{L}^{(M)}}(0)<\infty$ unconditionally (Theorem 6.9). The inactive $J_{0}$ factors provide exponential suppression at rate $\eta=e/4\approx 0.68$.
Stage 2: Finite truncation and stability. At any fixed verified truncation level $M$, the effective density bound (Theorem 8.5) gives $\operatorname{dens}\{N_{d}>0\}\leq A_{M}/\sqrt{d}+B_{M}$, which is nontrivial for sufficiently large $d$. The $O(1/\sqrt{d})$ rate (Corollary 8.8) follows under the hypothesis that $\Lambda_{\infty}$ has finite rank $d_{\infty}<\infty$ (verified $d_{\infty}=0$ for $M\leq 20$), which provides a uniform bound on the density and allows the truncation level to grow with $d$.
Stage 3: Verified refinements. Under computationally verified resonance absence at $M=20$, the telescoping argument (Theorem 7.7) establishes $f_{S_{L}}(0)<\infty$ for the full infinite series, and the exact formula $\operatorname{dens}\{N_{d}>0\}=C(d)/\sqrt{d}+o(1/\sqrt{d})$ follows from the computable marginal densities (Theorem 9.4), with $C(5)=0.1191$. No result in this paper assumes GRH or the Grand Simplicity Hypothesis.
Remark 11.1 (The $M\to\infty$ barrier).
The one step we cannot make unconditionally is the passage from verified resonance absence at a fixed $M$ to $d_{\infty}<\infty$. Proving $d_{\infty}=0$ would require either GSH or an unconditional proof that zero ordinates of $L$-functions admit no integer linear relations. The numerical vacuity of the fixed-$M$ bound (Remark 8.7) and the role of $d_{\infty}$ in the rate (Remark 8.9) are discussed in Section 8.
A structural asymmetry governs how GSH appears. The joint approach (studying $S_{\zeta}$ and $S_{L}$ on a combined torus) requires GSH for cross-independence, and its failure mode is a shared zero ordinate $\gamma_{k}=\gamma_{j}^{\prime}$. The separated approach used here handles each function independently and requires GSH only for taking $M\to\infty$ within a single function; its failure mode is an intra-function integer relation $\sum n_{k}\gamma_{k}^{\prime}=0$. Shared zeros are harmless in our framework, while intra-function relations affect the characteristic function only beyond the verified truncation level. The paper’s architecture selects the formulation in which the GSH obstruction lies behind the verified horizon.
Remark 11.2 (Coupled scaling).
When $d$ determines both the scaling factor in $N_{d}$ and the character $\chi_{d}$, the density decays faster than $O(1/\sqrt{d})$. The first zero $\gamma_{1}^{\prime}(d)$ of $L(s,\chi_{d})$ satisfies $\gamma_{1}^{\prime}(d)\sim A/\log d$ by the zero-counting formula, which implies $\sigma_{L}(d)\sim\log d$ and hence $C(d)\sim 1/\log d$. The true rate is therefore
$\operatorname{dens}\{N_{d}>0\}=O\!\left(\frac{1}{\sqrt{d}\,\log d}\right).$
This refinement applies when varying $d$ across different $L$-functions; for a fixed $L$-function with $d$ as a scaling parameter only, the rate $C/\sqrt{d}$ is sharp.
Remark 11.3.
We acknowledge that the depth of the proof is inversely proportional to the surprise of the result.
12. The Null Crossing and Spectral Phase Transition
12.1. The norm form on the real axis
For real $s>1$, define $N(s)=\zeta(s)^{2}-d\cdot L(s,\chi_{d})^{2}$.
Lemma 12.1 (Ratio monotonicity).
The ratio $G(s)=\zeta(s)/L(s,\chi_{d})$ is strictly decreasing on $(1,\infty)$ with $G(1^{+})=+\infty$ and $G(\infty)=1$.
Proof.
Taking logarithmic derivatives:
$\frac{d}{ds}\log G(s)=\frac{\zeta^{\prime}(s)}{\zeta(s)}-\frac{L^{\prime}(s,\chi_{d})}{L(s,\chi_{d})}=-\sum_{p}f(p,s),$
where $f(p,s)=\frac{\log p}{p^{s}-1}-\frac{\chi_{d}(p)\log p}{p^{s}-\chi_{d}(p)}$. Split primes ($\chi_{d}(p)=+1$) contribute $f=0$. Inert primes ($\chi_{d}(p)=-1$) contribute $f=\frac{2p^{s}\log p}{p^{2s}-1}>0$. Ramified primes contribute $f=\frac{\log p}{p^{s}-1}>0$. Since infinitely many primes are inert (Chebotarev), $(d/ds)\log G(s)<0$. ∎
12.2. The unique null crossing
Theorem 12.2 (Unique null crossing).
For every squarefree $d>1$, there exists a unique $s_{*}(d)\in(1,\infty)$ such that $N(s_{*}(d))=0$. For $1<s<s_{*}(d)$, the norm form is positive; for $s>s_{*}(d)$, it is negative.
Proof.
By Lemma 12.1, $G(s)$ is strictly decreasing from $\infty$ to $1$. Since $\sqrt{d}>1$, the equation $G(s)=\sqrt{d}$ has a unique solution $s_{*}(d)$. The norm form $N(s)=L(s,\chi_{d})^{2}(G(s)^{2}-d)$ vanishes precisely there. ∎
$d$ $L(1,\chi_{d})$ $s_{*}(d)$
2 0.6232 2.5635
3 0.7603 2.0000
5 0.4304 2.0492
7 1.0465 1.4445
13 0.6627 1.4608
Table 1. Null crossing $s_{*}(d)$ for small discriminants.
Values of $L(1,\chi_{d})$ were computed from the class number formula $L(1,\chi_{d})=2hR/\sqrt{|\Delta_{K}|}$, where $h$ is the class number and $R=\log\varepsilon$ is the regulator with fundamental unit $\varepsilon$ of $\mathcal{O}_{K}$; all five fields have $h=1$. The null crossings $s_{*}(d)$ were found by bisecting $G(s)=\zeta(s)/L(s,\chi_{d})$ against the target $\sqrt{d}$, with $L(s,\chi_{d})$ evaluated via the Hurwitz decomposition $L(s,\chi_{d})=q^{-s}\sum_{a=1}^{q}\chi_{d}(a)\,\zeta(s,a/q)$ using ARB interval arithmetic at 512-bit precision ($200$ bisection steps, residual $|G(s_{*})-\sqrt{d}|<10^{-60}$).
12.3. Integer crossing obstruction
Theorem 12.3 (Integer crossing).
Among all real quadratic fields, $\mathbb{Q}(\sqrt{3})$ is the unique field with integer null crossing: $s_{*}(3)=2$. No other real quadratic field has $s_{*}(d)\in\mathbb{Z}$.
Proof.
We show $G(m)\neq\sqrt{d}$ for all integers $m\geq 2$ and all squarefree $d\geq 2$, except $m=2$, $d=3$. The key tool is the Euler product
$G(s)=\frac{\zeta(s)}{L(s,\chi_{d})}=\prod_{p}\frac{1-\chi_{d}(p)\,p^{-s}}{1-p^{-s}}.$
Each local factor satisfies $F_{p}=1$ (split), $F_{p}=1/(1-p^{-s})$ (ramified), or $F_{p}=(1+p^{-s})/(1-p^{-s})$ (inert). Since $1\leq 1/(1-x)\leq(1+x)/(1-x)$ for $0<x<1$, every factor is bounded by the inert value, giving the all-inert bound
$G(s)\leq\prod_{p}\frac{1+p^{-s}}{1-p^{-s}}=\frac{\zeta(s)^{2}}{\zeta(2s)},$ (16)
valid for all $s>1$ and all squarefree $d\geq 2$.
Case 1: $s=2$ and $d=3$. Direct computation: $\zeta(2)=\pi^{2}/6$ and $L(2,\chi_{3})=\pi^{2}/(6\sqrt{3})$. Therefore
$\zeta(2)^{2}-3\cdot L(2,\chi_{3})^{2}=\frac{\pi^{4}}{36}-3\cdot\frac{\pi^{4}}{108}=\frac{\pi^{4}}{36}-\frac{\pi^{4}}{36}=0.$
Case 2: $s=2$ and $d\neq 3$. The Klingen–Siegel theorem gives $L(2,\chi_{d})=c_{d}\cdot\pi^{2}/\sqrt{d}$ for an explicit rational $c_{d}$ depending on $d$. The null crossing condition $\zeta(2)^{2}=d\cdot L(2,\chi_{d})^{2}$ becomes $\pi^{4}/36=c_{d}^{2}\cdot\pi^{4}$, requiring $c_{d}=1/6$. This holds only for $d=3$.
Case 3: $s\geq 4$ (all integers). Each factor $(1+p^{-s})/(1-p^{-s})$ is decreasing in $s$ (since $p^{-s}$ decreases), so the all-inert bound (16) is decreasing. At $s=4$, the Bernoulli number evaluations $\zeta(4)=\pi^{4}/90$ and $\zeta(8)=\pi^{8}/9450$ give
$\frac{\zeta(4)^{2}}{\zeta(8)}=\frac{(\pi^{4}/90)^{2}}{\pi^{8}/9450}=\frac{9450}{8100}=\frac{7}{6}.$
For all $s\geq 4$: $G(s)\leq 7/6$. Since $(7/6)^{2}=49/36<2\leq d$, the inequality $G(s)<\sqrt{d}$ holds for every squarefree $d\geq 2$ and every integer $s\geq 4$.
Case 4: $s=3$, $d\geq 3$. The all-inert bound gives $G(3)\leq\zeta(3)^{2}/\zeta(6)=\prod_{p}(p^{3}+1)/(p^{3}-1)$. The partial product through $p=5$ equals $(9\cdot 28\cdot 126)/(7\cdot 26\cdot 124)=567/403$. Each remaining factor satisfies $(p^{3}+1)/(p^{3}-1)<1+3p^{-3}$ (since $2/(p^{3}-1)<3/p^{3}$ for $p\geq 2$), so $\prod_{p>5}(p^{3}+1)/(p^{3}-1)<\exp(3\sum_{p>5}p^{-3})<\exp(3/50)$, using $\sum_{p>5}p^{-3}<\int_{5}^{\infty}x^{-3}\,dx=1/50$. Since $\exp(y)<1+y+y^{2}$ for $0<y<1$, the tail is bounded by $\exp(3/50)<2659/2500$, giving $G(3)\leq(567/403)(2659/2500)=1507653/1007500$. Then $1507653^{2}=2{,}273{,}017{,}568{,}409<3{,}045{,}168{,}750{,}000=3\cdot 1007500^{2}$, so $G(3)<\sqrt{3}\leq\sqrt{d}$.
Case 5: $s=3$, $d=2$. In $\mathbb{Q}(\sqrt{2})$ (discriminant $\Delta_{K}=8$), the Kronecker symbol $(8/p)$ determines splitting: $p=2$ is ramified, primes $p\equiv\pm 1\pmod{8}$ split, and primes $p\equiv\pm 3\pmod{8}$ are inert. The exact local factors at $p\leq 7$ are
$F_{2}=\frac{8}{7}\;\text{(ramified)},\quad F_{3}=\frac{14}{13}\;\text{(inert)},\quad F_{5}=\frac{63}{62}\;\text{(inert)},\quad F_{7}=1\;\text{(split)},$
giving a partial product of $504/403$. For the tail, each remaining factor satisfies $F_{p}\leq(1+p^{-3})/(1-p^{-3})<1+3p^{-3}$ (since $p^{-3}<1/3$), giving
$\prod_{p>7}F_{p}<\exp\!\Bigl(3\sum_{p>7}p^{-3}\Bigr)<\exp\!\Bigl(\tfrac{3}{98}\Bigr)<1+\tfrac{3}{98}+\tfrac{9}{9604}=\frac{9907}{9604},$
using $\sum_{p>7}p^{-3}<\int_{7}^{\infty}x^{-3}\,dx=1/98$ and $\exp(y)<1+y+y^{2}$ for $0<y<1$. Therefore $G(3)\leq(504/403)(9907/9604)=178326/138229$, and
$G(3)^{2}\leq\frac{178326^{2}}{138229^{2}}=\frac{31{,}800{,}162{,}276}{19{,}107{,}256{,}441}<2,$
since $31{,}800{,}162{,}276<2\times 19{,}107{,}256{,}441=38{,}214{,}512{,}882$. ∎
Remark 12.4 (Spectral phase transition).
The null crossing $s_{*}(d)$ separates a “zeta-dominated” regime ($1<s<s_{*}(d)$) from an “$L$-dominated” regime ($s>s_{*}(d)$). The spacelike theorem (Corollary 3.5) shows that at $\delta=0$ in the spectral shift variable, the zero data lies in the $L$-dominated (spacelike) region.
13. Connections and Open Problems
13.1. Relation to Katz–Sarnak statistics
The Katz–Sarnak density conjecture [6] predicts that zeros of $L(s,\chi_{d})$ near the central point follow symplectic statistics. The low-lying zero dominance theorem is consistent with this: as $d\to\infty$, the first zero height $\gamma_{1}^{\prime}(d)$ approaches zero, and its Lorentzian weight dominates.
Open Problem 13.1.
Derive quantitative predictions for $\gamma_{1}^{\prime}(d)$ from the spacelike constraint and compare with Katz–Sarnak.
13.2. Extensions to higher degree
The framework extends to cyclic extensions $K/\mathbb{Q}$ of prime degree $p$, with the norm form replaced by the determinant of a circulant matrix. The negativity persists, but the integer crossing obstruction strengthens: no cyclic extension of degree $p\geq 3$ has an integer null crossing.
Open Problem 13.2.
Does the spacelike constraint for cyclic $p$-extensions imply bounds on low-lying zeros of degree-$p$ Artin $L$-functions?
13.3. Further questions
The norm-form energy $N(s)=\zeta(s)^{2}-d\cdot L(s,\chi_{d})^{2}$ inherits symmetries from both factors.
Open Problem 13.3.
Is there a functional equation relating $N(s)$ to N(1s)N(1-s)?
The spacelike constraint provides global information about zero distribution. A natural question is whether it implies local constraints.
Open Problem 13.4.
Does the spacelike constraint imply bounds on zero spacing or pair correlation?
Under verified resonance absence, the density of positive excursions vanishes as $O(1/\sqrt{d})$. An unconditional proof of the rate (without the hypothesis $d_{\infty}<\infty$) would require a uniform bound on $\|f_{S_{L}^{(M)}}\|_{\infty}$ that does not rely on $d_{\infty}=0$.
Open Problem 13.5.
Characterize the fields $K$ for which $\operatorname{dens}(\{N>0\})=0$ unconditionally.
The $O(1/\sqrt{d})$ density rate emerges in this paper from a specific chain: low-lying zero dominance forces $S_{\zeta}<\sqrt{d}\cdot S_{L}$, the Lorentzian containment lemma converts this into a level-crossing problem for $S_{L}$, and the Jessen–Wintner theory identifies the crossing density as an integral over the marginal distribution of $S_{L}$, which is $O(1/\sqrt{d})$ because $S_{L}$ scales as $1/\sqrt{d}$ in the support width of $f_{S_{\zeta}}$. What is absent is an explanation of why this rate must be $C/\sqrt{d}$ rather than some other power of $d$, and why the constant $C$ is universal across all quadratic fields. The $\sqrt{d}$ arises from the coefficient of the norm form $N=S_{\zeta}^{2}-d\cdot S_{L}^{2}$; the density law is asking how often the Lorentzian energy crosses zero, and the crossing rate is governed by the oscillation amplitude of $S_{L}$ relative to $S_{\zeta}/\sqrt{d}$.
Open Problem 13.6 (Universal Density Law).
Is there a dynamical or geometric proof that $\operatorname{dens}\{N_{d}>0\}\sim C/\sqrt{d}$ as $d\to\infty$, valid unconditionally, in which the $\sqrt{d}$ exponent is derived directly from the spectral geometry of the norm form $N=S_{\zeta}^{2}-d\cdot S_{L}^{2}$? Concretely: does the indefinite quadratic structure of $N$ force the $O(1/\sqrt{d})$ rate by a level-crossing argument in the space of almost periodic functions, without passing through the Jessen–Wintner characteristic function machinery? A natural approach would identify the density $\operatorname{dens}\{N>0\}$ with the volume of a thin slab $\{|S_{L}|<S_{\zeta}/\sqrt{d}\}$ in an ergodic flow on a compact abelian group, and derive the slab width from the geometry of the quadratic form. The existing proof is analytic; a geometric proof would explain why the law is universal.
The construction exhibits a natural stability under truncation and limit.
Remark 13.7 (Functorial perspective).
The passage from spectral data to scalar energy is functorial with respect to inclusion: if $\Sigma_{M}\subset\Sigma_{M+1}$ denotes the inclusion of the first $M$ zeros into the first $M+1$ zeros, then the induced maps on weighted sums, characteristic functions, and density bounds commute with the truncation and limit operations. Concretely:
  1. (i)
    The spectral sum $S_{L}^{(M)}\to S_{L}^{(M+1)}$ extends by adding a single term $b_{M+1}\cos(\gamma_{M+1}^{\prime}\delta)$.
  2. (ii)
    The characteristic function $\varphi_{M}(t)\to\varphi_{M+1}(t)$ multiplies by the factor $J_{0}(b_{M+1}t)$.
  3. (iii)
    The density bound $f^{(M)}(0)\to f^{(M+1)}(0)$ changes by at most $O(b_{M+1}^{2})$ (Theorem 7.4).
This functorial structure is why the telescoping argument works: each level controls the next, and the limit inherits finiteness from the tower.
13.4. Weight robustness
The Lorentzian weight $w(\gamma)=2/(\tfrac{1}{4}+\gamma^{2})$ arises naturally from the Weil explicit formula with test function $g(x)=e^{-|x|/2}$. However, the spacelike property is not an artifact of this particular choice.
Definition 13.8 (Admissible weight).
A function $w:\mathbb{R}^{+}\to\mathbb{R}^{+}$ is an admissible spectral weight if it satisfies:
  1. (i)
    $w$ is monotonically decreasing,
  2. (ii)
    $w(\gamma)=O(1/\gamma^{2})$ as $\gamma\to\infty$,
  3. (iii)
    $w(0)<\infty$.
Condition (ii) ensures absolute convergence of the spectral sums $S_{f}=\sum_{\rho}w(\rho)$. Condition (i) is the essential structural requirement: it ensures that zeros at lower height receive greater weight.
Proposition 13.9 (Weight robustness).
Let $w$ be any admissible spectral weight. Define the weighted spectral sums
$S_{\zeta}^{(w)}:=\sum_{\rho:\,\zeta(\rho)=0}w(\gamma),\qquad S_{L}^{(w)}:=\sum_{\rho^{\prime}:\,L(\rho^{\prime},\chi_{d})=0}w(\gamma^{\prime}),$
and the norm-form energy $N_{w}:=(S_{\zeta}^{(w)})^{2}-d\cdot(S_{L}^{(w)})^{2}$. Then $N_{w}<0$ for all fundamental discriminants $d>1$.
Proof.
We show $S_{L}^{(w)}>S_{\zeta}^{(w)}$ for any admissible weight $w$ by comparing the zero-counting functions.
Step 1: Zero-counting domination. Let $N_{\zeta}(T)=\#\{\gamma_{k}\leq T\}$ and $N_{L}(T)=\#\{\gamma_{j}^{\prime}\leq T\}$ count zeros with positive imaginary part up to height $T$. For $T\in[0,\gamma_{1}(\zeta))$, $N_{\zeta}(T)=0$ while $N_{L}(T)\geq 1$ (Proposition 3.9). For $T\geq\gamma_{1}(\zeta)$, the Riemann–von Mangoldt formula gives $N_{L}(T)-N_{\zeta}(T)=(T/\pi)\log q+O(\log T)$, which exceeds $2$ for all $T\geq 14.13$ and $q\geq 5$.
Step 2: Abel summation. For any decreasing weight $w$ with $w(\gamma)=O(1/\gamma^{2})$, the spectral sum admits the Abel representation
$S_{f}^{(w)}=\sum_{\rho}w(\gamma)=\int_{0}^{\infty}w(t)\,dN_{f}(t)=-\int_{0}^{\infty}w^{\prime}(t)\,N_{f}(t)\,dt,$
where the boundary terms vanish since $w(t)N_{f}(t)\to 0$ as $t\to\infty$ (by the decay condition on $w$ and the polynomial growth of $N_{f}$). Since $w$ is decreasing, $w^{\prime}(t)\geq 0$, and the integral is a positive linear functional of $N_{f}$.
Step 3: Comparison. Since $N_{L}(T)\geq N_{\zeta}(T)$ for all $T\geq 0$ (the $L$-function has its first zero below $14.13$ while $\zeta$ has none, and maintains a surplus at every height by the zero-counting formula), and $w^{\prime}(t)\geq 0$:
$S_{L}^{(w)}=-\int_{0}^{\infty}w^{\prime}(t)\,N_{L}(t)\,dt\geq-\int_{0}^{\infty}w^{\prime}(t)\,N_{\zeta}(t)\,dt=S_{\zeta}^{(w)}.$
Equality is excluded because $N_{L}(T)>N_{\zeta}(T)$ on the interval $0,\gamma_{1}(\zeta))$ (where $N_{\zeta}=0$ but $N_{L}\geq 1$), and $w^{\prime}$ is strictly positive on this interval. Therefore $S_{L}^{(w)}>S_{\zeta}^{(w)}$ strictly.
Since $d>1$, we have $S_{\zeta}^{(w)}<S_{L}^{(w)}<\sqrt{d}\cdot S_{L}^{(w)}$, so $N_{w}=(S_{\zeta}^{(w)})^{2}-d\cdot(S_{L}^{(w)})^{2}<0$. ∎
Remark 13.10 (Alternative weights).
Several weights arise naturally from mathematical structure:
  1. (a)
    The sech weight $w(\gamma)=\operatorname{sech}(\gamma)$ is self-dual under the Fourier transform: $\mathcal{F}[\operatorname{sech}(\pi x)](t)=\operatorname{sech}(t)$. It has no free parameters and decays exponentially.
  2. (b)
    The heat kernel $w(\gamma,t)=e^{-t\gamma^{2}}$ arises from spectral zeta function theory. As $t\to 0^{+}$, it approaches zero counting; as $t\to\infty$, it isolates the lowest zero.
  3. (c)
    The Li coefficients [8] $\lambda_{n}=\sum_{\rho}[1-(1-1/\rho)^{n}]$ form a sequence directly related to the Riemann Hypothesis: RH holds if and only if $\lambda_{n}>0$ for all $n\geq 1$.
Table 2 records $S_{\zeta}(w)$, $S_{L}(w)$, and $N_{w}=S_{\zeta}(w)^{2}-d\cdot S_{L}(w)^{2}$ for four admissible weights across four fundamental discriminants, computed using the first 60 zeros of $\zeta(s)$ (Appendix D), the first 60 certified zeros of $\chi_{5}$, and the first 20 certified zeros of $\chi_{2}$, $\chi_{3}$, $\chi_{13}$ (from Appendices $C$). All values of $N_{w}$ are negative, confirming the spacelike property unconditionally at these truncation levels. The sech and heat-kernel sums for $\zeta$ are near-zero because all Riemann zeros satisfy $\gamma_{k}>14$, so exponentially decaying weights are negligible on the $\zeta$ side; the spacelike sign is then driven entirely by $d\cdot S_{L}(w)^{2}<0$.
Table 2. Spacelike verification across admissible weights. All computations use ARB interval arithmetic at 512-bit precision, so every entry is a rigorous certified value. The $\zeta$ sum uses 60 zeros (Appendix D); the $L$-function sums use 60 certified zeros for $\chi_{5}$ and 20 certified zeros for $\chi_{2}$, $\chi_{3}$, $\chi_{13}$ (from Appendices $C$). The norm form is $N_{w}=S_{\zeta}(w)^{2}-d\cdot S_{L}(w)^{2}$ with squarefree $d$. Every $N_{w}$ entry is rigorously negative: the ARB ball lies entirely below zero. Entries rounded to four significant figures.
Weight Quantity d=5(q=5)d=5\ (q=5) d=2(q=8)d=2\ (q=8) d=3(q=12)d=3\ (q=12) d=13(q=13)d=13\ (q=13)
Lorentzian $\frac{2}{1/4+\gamma^{2}}$ $S_{\zeta}(w)$ $0.03793$ $0.03793$ $0.03793$ $0.03793$
$S_{L}(w)$ $0.1406$ $0.1979$ $0.2863$ $0.3514$
$N_{w}$ 0.09742-0.09742 0.07693-0.07693 0.2444-0.2444 1.604-1.604
sech $\operatorname{sech}(\gamma)$ $S_{\zeta}(w)$ $1.5\times 10^{-6}$ $1.5\times 10^{-6}$ $1.5\times 10^{-6}$ $1.5\times 10^{-6}$
$S_{L}(w)$ $2.713\times 10^{-3}$ $1.592\times 10^{-2}$ $4.730\times 10^{-2}$ $9.008\times 10^{-2}$
$N_{w}$ $3.68\times 10^{-5}$ $5.07\times 10^{-4}$ $6.71\times 10^{-3}$ 0.1055-0.1055
Heat kernel $e^{-0.01\gamma^{2}}$ $S_{\zeta}(w)$ $0.1497$ $0.1497$ $0.1497$ $0.1497$
$S_{L}(w)$ $1.417$ $2.080$ $2.652$ $2.765$
$N_{w}$ 10.02-10.02 8.630-8.630 21.08-21.08 99.37-99.37
Heat kernel $e^{-0.1\gamma^{2}}$ $S_{\zeta}(w)$ $3.5\times 10^{-9}$ $3.5\times 10^{-9}$ $3.5\times 10^{-9}$ $3.5\times 10^{-9}$
$S_{L}(w)$ $1.210\times 10^{-2}$ $9.361\times 10^{-2}$ $2.469\times 10^{-1}$ $3.839\times 10^{-1}$
$N_{w}$ $7.32\times 10^{-4}$ $1.75\times 10^{-2}$ 0.1828-0.1828 1.916-1.916
Remark 13.11 (Lorentzian truncation).
[…]
Remark 13.12 (Non-decaying weights).
The decay condition (ii) is essential. The de Branges reproducing kernel $K_{\Delta}(\gamma,\gamma)\to\Delta$ as $\gamma\to\infty$ does not decay, and the corresponding weighted sums count zeros rather than emphasizing low-lying zeros. For this kernel, $N>0$ for most discriminants.
Remark 13.13 (Lorentzian as minimal decay).
Among polynomial-decay weights $w(\gamma)=1/(\tfrac{1}{4}+\gamma^{2})^{\alpha}$, the Lorentzian corresponds to $\alpha=1$, the minimal exponent ensuring absolute convergence of the spectral sums. In this sense, the Lorentzian is the “least biased” polynomial weight toward low-lying zeros while still being admissible. The spacelike result holds for all $\alpha\geq 1$.
Remark 13.14 (Scope of this series).
The open problems collected in this section range from the immediately tractable (the functional equation for $N(s)$, zero spacing implications) to the deeply conditional (unconditional density rate, universal density law). The Persistent Heuristics Series is committed to pursuing only unconditional results because doing so forces one to find the exact limitations of currently available tools. Therefore, while various open problems are presented for others to pursue, this series neither confirms nor denies that any follow-on paper will actually pursue those precise problems.
Acknowledgments
The author thanks the LMFDB collaboration [9] for access to verified zero data. Claude (Anthropic) was used for manuscript preparation, including drafting, editing, and computational verification of numerical results. The author takes full responsibility for all mathematical content.
Appendix A Computational Methods and Numerical Verification
A.1. Software and rigorous arithmetic
Computations in this paper fall into two categories. Rigorous computations use interval arithmetic via ARB [1], accessed through python-flint $\geq 0.8.0$ (Python 3.12); ARB represents numbers as balls $m\pm r]$ with midpoint $m$ and radius $r$, and arithmetic operations propagate error bounds rigorously, so the true value is guaranteed to lie within the output ball as a mathematical theorem. This includes all zero certification, spectral sum bounds, Besicovitch constants, and Bessel product integrals (via acb_calc_integrate). The PSLQ integer relation search uses mpmath $\geq 1.3$; the Ferguson–Bailey–Arno bounds [3] ensure that None results rigorously certify the absence of integer relations with bounded coefficients (see §A.7).
Working precisions: 512-bit ($\approx 154$ decimal digits) for $S_{\zeta}^{*}$ certification, spacelike verification (Table 2), and null crossing computation (ARB); 500-bit for auxiliary $L$-function zeros (ARB); 1500-bit ($\approx 451$ decimal digits) for all $\chi_{5}$ zero refinement and certification (ARB); 256-bit for all Bessel product integrals including density, resonance, and decay estimates (ARB); and 80-digit decimal precision for PSLQ searches (mpmath).
A.2. $L$-function evaluation and zero-finding
All Dirichlet $L$-functions in this paper are evaluated via the Hurwitz zeta decomposition
$L(s,\chi_{d})=q^{-s}\sum_{a=1}^{q}\chi_{d}(a)\,\zeta(s,a/q),$
where $q=|\Delta_{K}|$ is the conductor. Each Hurwitz zeta value $\zeta(s,a/q)$ is computed by ARB’s built-in acb.zeta, which returns a certified complex ball. Character values $\chi_{d}(a)$ are computed from the Kronecker symbol $(\Delta_{K}/a)$.
Character values. The Hurwitz sum runs over all $a$ with $1\leq a\leq q$; terms with $\gcd(a,q)>1$ vanish since $\chi_{d}(a)=0$ for such $a$. For reproducibility, the nonzero character values for all eight characters used in this paper are recorded below. All characters are even ($\chi_{d}(-1)=+1$) and primitive.
Character $q$ Nonzero values $\chi_{d}(a)$ for $1\leq a<q$, $\gcd(a,q)=1$
$\chi_{5}$ 55 $\chi(1)=\chi(4)=+1$;  $\chi(2)=\chi(3)=-1$
$\chi_{2}$ 88 $\chi(1)=\chi(7)=+1$;  $\chi(3)=\chi(5)=-1$
$\chi_{3}$ $12$ $\chi(1)=\chi(11)=+1$;  $\chi(5)=\chi(7)=-1$
$\chi_{13}$ $13$ $\chi(1,3,4,9,10,12)=+1$;  $\chi(2,5,6,7,8,11)=-1$
$\chi_{6}$ $24$ $\chi(1,5,19,23)=+1$;  $\chi(7,11,13,17)=-1$
$\chi_{7}$ $28$ $\chi(1,3,9,19,25,27)=+1$;  $\chi(5,11,13,15,17,23)=-1$
$\chi_{10}$ $40$ $\chi(1,3,9,13,27,31,37,39)=+1$;  $\chi(7,11,17,19,21,23,29,33)=-1$
$\chi_{11}$ $44$ $\chi(1,5,7,9,19,25,35,37,39,43)=+1$;  $\chi(3,13,15,17,21,23,27,29,31,41)=-1$
Hardy $Z$-function. For a real primitive Dirichlet character $\chi_{d}$ with conductor $q$, define the Hardy phase
$\theta_{\chi}(t)=\begin{cases}\operatorname{Im}\!\bigl(\log\Gamma(s/2)\bigr)-\tfrac{t}{2}\log(q/\pi)&\text{if }\chi_{d}(-1)=+1\text{ (even)},\\[4.0pt] \operatorname{Im}\!\bigl(\log\Gamma((s{+}1)/2)\bigr)-\tfrac{t}{2}\log(q/\pi)&\text{if }\chi_{d}(-1)=-1\text{ (odd)},\end{cases}$
where $s=\tfrac{1}{2}+it$. The Hardy $Z$-function $Z_{\chi}(t)=e^{i\theta_{\chi}(t)}L(\tfrac{1}{2}+it,\chi_{d})$ is then real-valued for $t\in\mathbb{R}$, and its real zeros correspond to zeros of $L(s,\chi_{d})$ on the critical line. The phase is computed via ARB’s built-in acb.lgamma.
Stage 1: Sign-change scan. Evaluate $Z_{\chi}(t)$ on a uniform grid with step $0.25$ over $0,T_{\max}]$, where $T_{\max}$ is chosen to exceed the expected height of the last zero sought. The zero-counting formula $N(T,\chi_{d})\sim\frac{T}{\pi}\log\frac{qT}{2\pi e}$ provides a conservative estimate; for example, $200$ zeros of $L(s,\chi_{5})$ require $T_{\max}\approx 290$ (the asymptotic formula gives $N(290,\chi_{5})\approx 205$). Low precision (128-bit ARB) suffices since only the sign of $Z_{\chi}$ is required. The scan produces $\lceil T_{\max}/0.25\rceil$ evaluations; for $T_{\max}=290$ this is $1160$ evaluations, each taking a fraction of a second at 128-bit precision.
Stage 2: Bisection. For each consecutive pair $t_{a},t_{b}$ with $Z_{\chi}(t_{a})\,Z_{\chi}(t_{b})<0$, bisect on $Z_{\chi}$ to isolate the root. At 512-bit ARB precision, $100$ bisection steps produce $\sim$30 decimal digits from a bracket of width $0.25$, yielding seeds accurate to at least 20 decimal places. Certification at this stage: accept $\gamma^{\prime}$ when $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{d})|<10^{-30}$. (Note: Stages 1 and 2 need be performed only once per character; the resulting 20-digit seeds are tabulated in Appendix C and may be reused indefinitely as input to Stage 3.)
Stage 3: Newton refinement. Refine each seed by Newton iteration applied to $Z_{\chi}(t)$. The derivative is approximated by the finite difference $(Z_{\chi}(t+h)-Z_{\chi}(t))/h$ with $h=10^{-50}$; the midpoint of the resulting ARB ball is extracted after each Newton step (see Remark A.1 below). At 1500-bit precision, $8$ Newton iterations from a 20-digit seed converge to full working precision ($\approx 451$ digits), with certified bounds on $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{d})|$ typically in the range $10^{-358}$$10^{-400}$. The convergence is quadratic: each iteration approximately doubles the number of correct digits, so $20\to 40\to 80\to 160\to 320\to 451$ digits in five to six effective steps.
Certification. A zero ordinate $\gamma^{\prime}$ is accepted when $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{d})|<\text{threshold}$, with the threshold depending on application (typically $10^{-30}$ for 20-digit values, $10^{-68}$ for 70-digit values). The certified bound is computed by evaluating $L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{d})$ in ARB ball arithmetic and extracting the upper endpoint of the absolute value ball; this is a rigorous upper bound as a mathematical theorem. Readers who already possess 20-digit seeds (for example, from the tables in Appendix C) may skip directly to Stage 3.
Precomputation. The constants $\log q$, $\log(q/\pi)$, and the Hurwitz parameters $a/q$ for each nonzero character value should be computed once and stored as arb objects at the working precision. This avoids redundant string-to-ball conversions in the inner loop and reduces the per-evaluation cost by approximately $30\%$.
Remark A.1 (Implementation pitfalls).
Three points are essential for correct implementation.
(1) Newton iteration must target $Z_{\chi}$, not $L$. On the critical line, $L(\tfrac{1}{2}+it,\chi_{d})$ is complex-valued for non-trivial characters; its real part does not vanish at zeros of $L$, so Newton iteration on $\operatorname{Re}L$ converges to a spurious fixed point rather than to a zero. The Hardy phase rotation reduces the problem to root-finding for a real function of one real variable, where Newton’s method achieves quadratic convergence (doubling the number of correct digits per iteration).
(2) Midpoint extraction is required in ball arithmetic. The Newton update $t\mapsto t-Z(t)/Z^{\prime}(t)$ introduces radius from the division. Without extracting the ball midpoint after each step, the radius grows geometrically and the computation produces NaN within a few iterations. This is specific to rigorous interval arithmetic and does not arise in ordinary floating-point computation.
(3) The finite-difference step $h$ must be appropriate for the current precision. The step $h=10^{-50}$ is suitable for 1500-bit ($\approx 451$-digit) working precision: it is large enough to avoid catastrophic cancellation (the difference Z(t+h)Z(t)Z(t+h)-Z(t) retains $\approx 400$ significant digits) and small enough to make the $O(h)$ truncation error negligible relative to the quadratic convergence. A step that is too small (for example, $h=10^{-200}$ at 512-bit precision) produces a derivative estimate dominated by rounding error, causing Newton iteration to stall.
A.3. Certification of $S_{\zeta}^{*}$
The upper bound $S_{\zeta}\leq 0.04625<S_{\zeta}^{*}=0.04871$ is established by rigorous interval arithmetic with no numerical quadrature at any stage. All arithmetic uses ARB ball arithmetic at $512$-bit working precision ($\approx 154$ decimal digits).
Input data. The first $K=6000$ positive zero ordinates $\gamma_{1},\ldots,\gamma_{6000}$ of $\zeta(s)$, each accurate to $31$ decimal places, sourced from the LMFDB [9] and Odlyzko’s tables [10]. Each ordinate is loaded into an arb ball whose radius encloses the rounding error of the $31$-digit string representation.
Step 1: Direct sum. The partial sum
$S_{\zeta}^{(6000)}=\sum_{k=1}^{6000}\frac{2}{\tfrac{1}{4}+\gamma_{k}^{2}}$
is accumulated in ARB ball arithmetic. The certified result is
$S_{\zeta}^{(6000)}\in[0.045795374824191\ldots\;\pm\;7.4\times 10^{-152}].$
Step 2: Tail bound. Set $T_{0}=\gamma_{6000}+0.01\approx 6365.86$. Abel summation gives
$\sum_{\gamma>T_{0}}w(\gamma)\leq\int_{T_{0}}^{\infty}\frac{4t\,[N(t)-K]}{(\tfrac{1}{4}+t^{2})^{2}}\,dt,$
where $N(t)$ is the zero-counting function and the boundary term $2N(t)/(\tfrac{1}{4}+t^{2})]_{T_{0}}^{\infty}$ is non-positive (discarded as a further upper bound). The integral is evaluated without quadrature using the exact antiderivative identity
$\int_{a}^{b}\frac{4t}{(\tfrac{1}{4}+t^{2})^{2}}\,dt=\frac{2}{\tfrac{1}{4}+a^{2}}-\frac{2}{\tfrac{1}{4}+b^{2}}.$
The interval $T_{0},\,10^{6}]$ is partitioned into $235$ subintervals (width $100$ for $T_{0}\leq t\leq 20{,}000$; width 10,00010{,}000 for $20{,}000<t\leq 10^{6}$). On each subinterval $[a,b]$, the zero-counting function is bounded above by Theorem 2.5 evaluated at $t=b$ in ARB:
$N(t)-K\leq N_{\mathrm{upper}}(b)-K\quad\text{for all }t\in[a,b],$
giving a certified contribution of $(N_{\mathrm{upper}}(b)-K)\cdot(2/(\tfrac{1}{4}+a^{2})-2/(\tfrac{1}{4}+b^{2}))$ per subinterval. The sum over all $235$ subintervals is $\leq 4.31\times 10^{-4}$.
For $t\geq 10^{6}$, the bound $N(t)\leq t\log t$ (valid for all $t\geq 10$) and $4t/(\tfrac{1}{4}+t^{2})^{2}\leq 4/t^{3}$ yield the analytic remainder
$\int_{10^{6}}^{\infty}\frac{4\,N(t)}{(\tfrac{1}{4}+t^{2})^{2}}\,dt\leq\frac{2(\log U+1)}{\pi U}+\frac{4}{U}\bigg|_{U=10^{6}}\leq 1.35\times 10^{-5},$
evaluated in ARB. The total tail bound is $\leq 4.45\times 10^{-4}$.
Step 3: Certified result. Combining in ARB:
$S_{\zeta}\leq S_{\zeta}^{(6000)}+\text{tail}\in[0.046240\;\pm\;2.8\times 10^{-152}].$
The right endpoint of this interval is $0.046240\ldots<0.04625$, which lies $5.1\%$ below the declared bound $S_{\zeta}^{*}=0.04871$.
Cross-checks. The partial sum $S_{\zeta}^{(50)}=0.03708$ agrees with independent double-precision evaluation. The direct sum from zeros $51$ through $6000$ is $0.00871$; combined with the analytic tail $0.00045$, the total tail from $k>50$ is $0.00916$, which tightens the Abel summation estimate $0.00986$.
A.4. Verification of spacelike constraint
The low-lying zero dominance $S_{L}>S_{\zeta}^{*}$ is verified by the three-case argument of Theorem 3.1: the BMOR bound covers $q\geq 12$ (the smallest conductor of a fundamental discriminant exceeding those of $d\in\{2,5\}$), and the exceptions $d\in\{2,5\}$ are verified by direct computation of $L$-function zeros via the Hardy $Z$-function method.
$d$ $q$ Zeros $\leq 11$ $S_{L}$ (lower bound) Method
22 88 33 $0.13376$ $Z_{\chi}$ computation
55 55 22 $0.06563$ $Z_{\chi}$ computation
$q\geq 12$ $\geq 12$ $\geq 4$ $0.06598$ BMOR bound
In all cases, $S_{L}>S_{\zeta}^{*}=0.04871$.
A.5. Computation of $\epsilon_{M}$
The tail bounds $\epsilon_{M}=\sum_{k>M}b_{k}$ reported in the stability table of Section 7 were computed using the methodology of §A.1A.2. The first 200 zeros of $L(s,\chi_{5})$ were located and certified as described therein; the Lorentzian weights $b_{k}=2/({\tfrac{1}{4}+{\gamma_{k}^{\prime}}^{2}})$ were computed in ARB and summed. The analytic tail beyond zero 200 (at height $T=283.93$) was bounded by the integral $\frac{1}{\pi}\bigl(\frac{\log(5T/2\pi)}{T}+\frac{1}{T}\bigr)=0.0072$, derived from Abel summation against the zero-counting formula $N(T,\chi_{5})\sim\frac{T}{\pi}\log\frac{5T}{2\pi e}$.
A.6. Bessel product integrals
The density at the origin $f_{S_{L}}(0)$ and the mean absolute value $\mathbb{E}[|S_{\zeta}|]$ appearing in Corollary 9.5 are computed via rigorous numerical quadrature using ARB’s acb_calc_integrate function, which implements the Petras algorithm with adaptive Gauss-Legendre quadrature and certified error bounds [1]. All results are rigorous ARB balls.
For the density integral (14): $f_{S_{L}}(0)=\frac{1}{\pi}\int_{0}^{T}\prod_{k=1}^{20}J_{0}(b_{k}t)\,dt$ with $T=2000$. The Lorentzian weights $b_{k}=2/(\tfrac{1}{4}+{\gamma_{k}^{\prime}}^{2})$ are computed in ARB from the 70-digit certified zeros of Appendix B. The integrand is entire (products of $J_{0}$), so no branch cut handling is required. ARB quadrature at 256-bit precision yields the certified result:
$f_{S_{L}^{(20)}}(0)=8.312883337454846\ldots\;\pm\;5.84\times 10^{-44}.$
For the mean absolute value (15): $\mathbb{E}[|S_{\zeta}|]=\frac{2}{\pi}\int_{\epsilon}^{T}\frac{1-\prod_{k=1}^{20}J_{0}(a_{k}t)}{t^{2}}\,dt$ with $\epsilon=10^{-3}$ and T=50, [^0] T=50{,}000, where $a_{k}=2/(\tfrac{1}{4}+\gamma_{k}^{2})$ for $\zeta$ zeros. The integrand is $O(1)$ near $t=0$ (since $1-\prod J_{0}(a_{k}t)=O(t^{2})$); the contribution from $0,\epsilon]$ is bounded by $(\sum_{k}a_{k}^{2}/4)\cdot\epsilon<4\times 10^{-8}$ and added to the main integral. ARB quadrature yields:
$\mathbb{E}[|S_{\zeta}|]=0.00716132817045\ldots\;\pm\;2.20\times 10^{-24}.$
The product $C(5)=2\cdot f_{S_{L}}(0)\cdot\mathbb{E}[|S_{\zeta}|]=0.1191$ is certified to four significant figures.
A.7. PSLQ integer relation search
This subsection and the next complete the proof of Proposition 6.3. The 70-digit zero ordinates of Appendix B, certified by ARB (Appendix E), serve as input to the PSLQ algorithm.
Integer relation detection uses the PSLQ algorithm (Ferguson–Bailey–Arno [3]) via mpmath.pslq. The 70-digit ARB-certified zero ordinates from Appendix B serve as input at 80-digit working precision.
Search class Bound on $\max|n_{k}|$ Result
Full 20-dimensional $100$, $500$, $1000$ None
All $190$ pairs $(j,k)$ 10,00010{,}000 None (all pairs)
All $120$ triples from first $10$ zeros $500$ None (all triples)
The pairwise None result rules out all rational ratios $\gamma_{j}^{\prime}/\gamma_{k}^{\prime}=p/q$ with $\max(|p|,|q|)\leq 10{,}000$.
Remark A.2 (Rigorous certification via Ferguson–Bailey–Arno bounds).
[…]
Remark A.3 (Cautionary experiment).
As a precision control, PSLQ was also applied to the original $20$-digit values at $40$-digit working precision with $\max|n_{k}|\leq 1000$. This returned the vector
[…]
with $\max|n_{k}|=19$ and $\ell^{1}$ norm $184$, whose weighted sum vanishes to all $20$ displayed digits. Evaluating at $70$-digit precision gives $\sum n_{k}\gamma_{k}^{\prime}=2.72\times 10^{-18}$, unambiguously nonzero. The spurious detection is explained by the effective precision of a linear combination of $20$ terms with $\ell^{1}$ norm $184$: $20-\log_{10}(184)\approx 17.7$ digits, below the residual threshold. The $70$-digit certification provides ample margin above this threshold.
A.8. Bessel tail bound
For any integer vector $\mathbf{n}$ with $\max|n_{k}|\geq N=1001$, the standard bound $|J_{N}(x)|\leq(x/2)^{N}/N!$ gives, at the most pessimistic integration range $T=2000$ and with $b_{\max}=b_{1}=4.499\times 10^{-2}$:
$\left(\frac{b_{1}\cdot 2000}{2}\right)^{1001}\frac{1}{1001!}<10^{-916}.$
The count of integer vectors with $\max|n_{k}|=N$ in dimension $20$ is at most $(2N+1)^{20}\leq(2003)^{20}<10^{66}$. Summing over all $N\geq 1001$ gives total resonance contribution bounded by $10^{66}\cdot 10^{-916}=10^{-850}$.
The resonance correction is bounded by $10^{-850}$ and does not affect any reported digit of $f_{S_{L}^{(20)}}(0)=8.3129$. The PSLQ search (§A.7) and Bessel tail bound together certify $d_{20}=0$. This completes the proof of Proposition 6.3. ∎
A.9. Reproducibility
All algorithms and parameters required for independent verification are documented in §A.1A.8. Rigorous computations (zero certification, $S_{\zeta}^{*}$ bounds, spacelike verification, null crossings, Cesàro threshold verification, Bessel product integrals) use ARB via python-flint; PSLQ integer relation search uses mpmath with rigorous certification via Ferguson–Bailey–Arno bounds [3]. The character values required for the Hurwitz decomposition are tabulated in §A.2; all zero ordinates are given in Appendices B, C, and D. The certification of $S_{\zeta}^{*}$A.3) uses 6000 Riemann zeta zeros from the LMFDB [9]; this is the only external data dependency. A reader requires only python-flint $\geq 0.8.0$, mpmath $\geq 1.3$ (both freely available), and the LMFDB zeta zero tables (for the $S_{\zeta}^{*}$ tail bound only) to reproduce every result in this paper.
Appendix B Zeros of $L(s,\chi_{5})$ to 70 Decimal Places (Conductor 5, $d=5$)
The first 200 zeros of $L(s,\chi_{5})$ are given here to 70 decimal places. Certified upper bounds on $|L(\tfrac{1}{2}+i\gamma_{k}^{\prime},\chi_{5})|$ appear in the rightmost column and are summarized in Appendix E. For computational methodology, see §A.2.
ARB Bound
$n$ $\gamma_{n}^{\prime}$ (70 decimal places) on $|L(\tfrac{1}{2}+i\gamma_{n}^{\prime})|$
11 6.64845334472771471612327845997931784729858542324449837237471294671536936.6484533447277147161232784599793178472985854232444983723747129467153693 $3.65\times 10^{-400}$
22 9.83144443288666961634832134745844382188813289401550743042668652269440359.8314444328866696163483213474584438218881328940155074304266865226944035 $5.71\times 10^{-379}$
33 11.958845626083514530265658688262841817293112765723287717381364444659796711.9588456260835145302656586882628418172931127657232877173813644446597967 $2.45\times 10^{-377}$
44 16.033821128384235674593253782248083508749303711393763454343960473652264016.0338211283842356745932537822480835087493037113937634543439604736522640 $1.67\times 10^{-374}$
55 17.566994292325555202701595268144486403485240895508468763100055336820239617.5669942923255552027015952681444864034852408955084687631000553368202396 $6.87\times 10^{-379}$
66 19.540732622784750250378690022998553485526493445007156424107667950060473619.5407326227847502503786900229985534855264934450071564241076679500604736 $4.30\times 10^{-381}$
77 22.227405454459410911877624963081473365280642640560187108353392333259926322.2274054544594109118776249630814733652806426405601871083533923332599263 $1.80\times 10^{-378}$
88 24.588466217408195207656269976081979066247857188503256974763128082376573724.5884662174081952076562699760819790662478571885032569747631280823765737 $3.05\times 10^{-376}$
99 26.776095948004140116523574965270948586289370096100720747141444091727181626.7760959480041401165235749652709485862893700961007207471414440917271816 $4.82\times 10^{-383}$
$10$ 28.461035100177522475186978272325256271200734674528650404495427630566281928.4610351001775224751869782723252562712007346745286504044954276305662819 $3.67\times 10^{-379}$
$11$ 29.707909350480965569230986518657313605970196432418479892145402797880542429.7079093504809655692309865186573136059701964324184798921454027978805424 $5.77\times 10^{-383}$
$12$ 33.000456006870514367949759177211937317376500804631238343447920775517672933.0004560068705143679497591772119373173765008046312383434479207755176729 $2.60\times 10^{-381}$
$13$ 34.728812978904808674143729833981379136564007992078974829862022029197196934.7288129789048086741437298339813791365640079920789748298620220291971969 $7.03\times 10^{-380}$
$14$ 35.868638371812274594595048638876426333384181725562636611431705494857515935.8686383718122745945950486388764263333841817255626366114317054948575159 $8.26\times 10^{-388}$
$15$ 38.129184721436531850151418270372918926216187170437541987827769860251830338.1291847214365318501514182703729189262161871704375419878277698602518303 $4.26\times 10^{-378}$
$16$ 39.560572946403181705055095059954798425566825204646381762421875276690628839.5605729464031817050550950599547984255668252046463817624218752766906288 $1.04\times 10^{-377}$
$17$ 41.842438545791694308509306885312933146785226936948497507540546923147676541.8424385457916943085093068853129331467852269369484975075405469231476765 $1.47\times 10^{-382}$
$18$ 44.031290061441695044700908058424922514756441978971928142545909588119921544.0312900614416950447009080584249225147564419789719281425459095881199215 $3.53\times 10^{-371}$
1919 45.427300082782289388841561909609720649965874840453313174253112295195133545.4273000827822893888415619096097206499658748404533131742531122951951335 $1.04\times 10^{-383}$
2020 46.492727159491405345339199351675833673263826528873731752681947727399791446.4927271594914053453391993516758336732638265288737317526819477273997914 $1.44\times 10^{-379}$
2121 48.345661821067846176548201304493486929945215885626654705310956748852704448.3456618210678461765482013044934869299452158856266547053109567488527044 $7.47\times 10^{-384}$
$22$ 51.087751926746491355258257204131245513931822764598874889643028952539913351.0877519267464913552582572041312455139318227645988748896430289525399133 $4.99\times 10^{-375}$
$23$ 52.125902231316974188604130695930300524969293830813142026749809937021618752.1259022313169741886041306959303005249692938308131420267498099370216187 $1.86\times 10^{-381}$
$24$ 53.830445195442163351054269022338568937339170713109887047091903708283035253.8304451954421633510542690223385689373391707131098870470919037082830352 $6.33\times 10^{-386}$
$25$ 55.589280335404810150964588998492970666633756779445931871191432982474752955.5892803354048101509645889984929706666337567794459318711914329824747529 $6.76\times 10^{-379}$
$26$ 56.838865942890988462381701087979625873787502823421640673322572691726700956.8388659428909884623817010879796258737875028234216406733225726917267009 $3.20\times 10^{-377}$
$27$ 58.386117485358380826793925742317062296785252761744412031426483942467746058.3861174853583808267939257423170622967852527617444120314264839424677460 $3.82\times 10^{-376}$
$28$ 61.138864751921745109095463915572425456187565195524483622740465340768912261.1388647519217451090954639155724254561875651955244836227404653407689122 $2.98\times 10^{-378}$
$29$ 62.132904720233877934071529440383543136671977562274208379683708348891512562.1329047202338779340715294403835431366719775622742083796837083488915125 $3.12\times 10^{-386}$
$30$ 63.709543060981146771798494836335680246320729913038089464768398907875187863.7095430609811467717984948363356802463207299130380894647683989078751878 $6.60\times 10^{-377}$
$31$ 64.637359665374574101088589957514469169614418340871541926806467731667825864.6373596653745741010885899575144691696144183408715419268064677316678258 $3.08\times 10^{-374}$
$32$ 66.747641845937807830433496332522277799056278904231710999500592219175209366.7476418459378078304334963325222777990562789042317109995005922191752093 $1.44\times 10^{-384}$
$33$ 68.590727924825006540833515068030970994309372068387999238433409618374964168.5907279248250065408335150680309709943093720683879992384334096183749641 $7.49\times 10^{-385}$
$34$ 70.115837159193504702254162003288501122534880816675087043045923668594310970.1158371591935047022541620032885011225348808166750870430459236685943109 $1.94\times 10^{-376}$
$35$ 71.605959479243017199477227381292487650674837775870501191061771430349257871.6059594792430171994772273812924876506748377758705011910617714303492578 $7.09\times 10^{-376}$
$36$ 73.350105431468259324986481309687463592380080206128610550028406269746719673.3501054314682593249864813096874635923800802061286105500284062697467196 $2.28\times 10^{-381}$
$37$ 74.293688438016568417563919573323746752656162649388588394478627746168005574.2936884380165684175639195733237467526561626493885883944786277461680055 $9.97\times 10^{-379}$
$38$ 75.534904117228113405345989547687724838297800656054992950178626192713539375.5349041172281134053459895476877248382978006560549929501786261927135393 $1.36\times 10^{-378}$
$39$ 78.072170864382337194051431823672802856073274377354409700849470519410328178.0721708643823371940514318236728028560732743773544097008494705194103281 $2.97\times 10^{-372}$
$40$ 79.786062490097939106371571243171617631261869341105718896147463944735173779.7860624900979391063715712431716176312618693411057188961474639447351737 $3.45\times 10^{-378}$
$41$ 80.567300856410698261373757677661867641972294500859119722736169220177608280.5673008564106982613737576776618676419722945008591197227361692201776082 $3.03\times 10^{-379}$
$42$ 81.923380893154115311967989553905914195519790752160387751303590547918086981.9233808931541153119679895539059141955197907521603877513035905479180869 $5.41\times 10^{-383}$
$43$ 83.699157001874113162883288392240055676480712410299839420963639633825894883.6991570018741131628832883922400556764807124102998394209636396338258948 $3.38\times 10^{-377}$
$44$ 84.952507103754973661846904660485113866224814086703961086051329413504388084.9525071037549736618469046604851138662248140867039610860513294135043880 $8.50\times 10^{-372}$
$45$ 86.809996679338770496962453824123865539107306122702912569153552151872514786.8099966793387704969624538241238655391073061227029125691535521518725147 $9.72\times 10^{-386}$
$46$ 88.281776599709835593621908069365568911371943199556447973711012167487531688.2817765997098355936219080693655689113719431995564479737110121674875316 $7.96\times 10^{-391}$
$47$ 90.376573182128156802963266298552786636345751535570441691904969596990152990.3765731821281568029632662985527866363457515355704416919049695969901529 $2.68\times 10^{-371}$
$48$ 90.715525933146974977269052021613460765711009835158005422179562303740599890.7155259331469749772690520216134607657110098351580054221795623037405998 $4.10\times 10^{-379}$
$49$ 92.450243253774403222656651609940166532625651637342691786476584587897979992.4502432537744032226566516099401665326256516373426917864765845878979799 $9.68\times 10^{-380}$
$50$ 93.394423937828017400982646092448145456052891559124280035714176687016475693.3944239378280174009826460924481454560528915591242800357141766870164756 $5.09\times 10^{-377}$
$51$ 95.970001646037706939832706231680914850344362198662140910938755275316429995.9700016460377069398327062316809148503443621986621409109387552753164299 $3.46\times 10^{-382}$
$52$ 97.331588252898071590582881239627993663344456879633576511321148041911903697.3315882528980715905828812396279936633444568796335765113211480419119036 $1.34\times 10^{-379}$
$53$ 98.227457869901077077061547196562023327793547734165391303478765895425843798.2274578699010770770615471965620233277935477341653913034787658954258437 $7.80\times 10^{-377}$
$54$ 99.838050807536549892458178740443727323094718661873207723162307044585847899.8380508075365498924581787404437273230947186618732077231623070445858478 $2.09\times 10^{-387}$
$55$ 101.1877170969186106879893673648615779856303956787129404948013652775331402101.1877170969186106879893673648615779856303956787129404948013652775331402 $8.26\times 10^{-389}$
$56$ 102.6052879410500105919016338097244199966334774046924761659561257554903097102.6052879410500105919016338097244199966334774046924761659561257554903097 $7.88\times 10^{-387}$
$57$ 103.6661310797677629747891494548931064951051111049995242765910102899167622103.6661310797677629747891494548931064951051111049995242765910102899167622 $1.11\times 10^{-399}$
$58$ 105.9731937977587861297183750394317655839089623673242427608883134019122134105.9731937977587861297183750394317655839089623673242427608883134019122134 $9.41\times 10^{-387}$
$59$ 107.2971036375259575240109457236330672603545936270064737698887145263170400107.2971036375259575240109457236330672603545936270064737698887145263170400 $4.62\times 10^{-386}$
$60$ 108.4981066242615983601377772869913646076568520675142124199793643754286051108.4981066242615983601377772869913646076568520675142124199793643754286051 $1.63\times 10^{-382}$
$61$ 109.5502586071841272439535415119528096196195061162254833594241784691252750109.5502586071841272439535415119528096196195061162254833594241784691252750 $9.23\times 10^{-388}$
$62$ 110.6361976726278518372473829197474157598718943493419745206490444961344614110.6361976726278518372473829197474157598718943493419745206490444961344614 $1.65\times 10^{-384}$
$63$ 112.6157699450494525603420477631256573284904001732412552640056168054705385112.6157699450494525603420477631256573284904001732412552640056168054705385 $1.47\times 10^{-378}$
$64$ 114.2306398723819524205236259115864715634752773232857784617817436415063754114.2306398723819524205236259115864715634752773232857784617817436415063754 $8.49\times 10^{-385}$
$65$ 115.5784139823770731492864674907752976644587354074111098082626495892785777115.5784139823770731492864674907752976644587354074111098082626495892785777 $2.08\times 10^{-382}$
$66$ 116.6356717597837992860716616244380290420850400876742741600872550143161217116.6356717597837992860716616244380290420850400876742741600872550143161217 $5.60\times 10^{-393}$
$67$ 118.3711623063552621519304519550474639914856549056003421824775094839317423118.3711623063552621519304519550474639914856549056003421824775094839317423 $1.09\times 10^{-367}$
$68$ 119.5614916258800030014640284654174824520291754354058843504968760508448926119.5614916258800030014640284654174824520291754354058843504968760508448926 $5.21\times 10^{-386}$
$69$ 120.2817341918527579929633987893433395906991345403172727276829320830538688120.2817341918527579929633987893433395906991345403172727276829320830538688 $6.70\times 10^{-376}$
$70$ 122.0616322838941436366460069970312459167329049609093140518854982869671603122.0616322838941436366460069970312459167329049609093140518854982869671603 $2.75\times 10^{-383}$
$71$ 124.2428371489503574456280462431539381366392074620427334384975693951880954124.2428371489503574456280462431539381366392074620427334384975693951880954 $2.98\times 10^{-375}$
$72$ 125.2815361690544786849959629329922579004633315004421413848554281307878485125.2815361690544786849959629329922579004633315004421413848554281307878485 $4.25\times 10^{-376}$
$73$ 126.3986993784245731137991303757675004132192700408244704238571151980971242126.3986993784245731137991303757675004132192700408244704238571151980971242 $3.04\times 10^{-387}$
$74$ 127.3206348281152351166040431083884153111288443677405189033098037679391082127.3206348281152351166040431083884153111288443677405189033098037679391082 $2.80\times 10^{-381}$
$75$ 129.0483000943208951329053997018161493807340037666172892353902573266257761129.0483000943208951329053997018161493807340037666172892353902573266257761 $1.49\times 10^{-385}$
$76$ 130.1970382668577478738346244894136617196218666973159138428989678080490833130.1970382668577478738346244894136617196218666973159138428989678080490833 $4.04\times 10^{-397}$
$77$ 132.0860053031175217593818811988220993124226919197273881005207706354373250132.0860053031175217593818811988220993124226919197273881005207706354373250 $1.69\times 10^{-374}$
$78$ 133.0302283883319444007316023118379973226202504376344453082369748548055878133.0302283883319444007316023118379973226202504376344453082369748548055878 $5.31\times 10^{-377}$
$79$ 134.9249638941115606306389354461440563212716929502899197778761945615660047134.9249638941115606306389354461440563212716929502899197778761945615660047 $1.13\times 10^{-378}$
$80$ 135.8744976900983831529154942879362038033124413255633587171958644086561842135.8744976900983831529154942879362038033124413255633587171958644086561842 $8.03\times 10^{-394}$
$81$ 137.2010834668406301716830996653811071822066893852372048164365956206970149137.2010834668406301716830996653811071822066893852372048164365956206970149 $3.01\times 10^{-379}$
$82$ 138.0356990313049020042749619065482729585857816235533229199603352798733897138.0356990313049020042749619065482729585857816235533229199603352798733897 $2.52\times 10^{-379}$
$83$ 139.4105995951457328667296699990735809900833487908551282462172352268054339139.4105995951457328667296699990735809900833487908551282462172352268054339 $1.40\times 10^{-375}$
$84$ 141.9462405516049128078346136764138632079930804595332431920642151809404665141.9462405516049128078346136764138632079930804595332431920642151809404665 $1.77\times 10^{-375}$
$85$ 142.5245124329877516499138087023775533114539886176690581797696212087823187142.5245124329877516499138087023775533114539886176690581797696212087823187 $1.05\times 10^{-380}$
$86$ 143.9615911232873249679700260946989275500715387277893452624660630780480223143.9615911232873249679700260946989275500715387277893452624660630780480223 $1.42\times 10^{-378}$
$87$ 144.9483276132617096716684583857500230575404813669059877812679180958528794144.9483276132617096716684583857500230575404813669059877812679180958528794 $1.37\times 10^{-375}$
$88$ 146.3848096091094233623717818286364786730892073299368170400054147265509705146.3848096091094233623717818286364786730892073299368170400054147265509705 $2.53\times 10^{-378}$
$89$ 147.8797291031824647001555355926284591276893974747116792549589248487923432147.8797291031824647001555355926284591276893974747116792549589248487923432 $4.67\times 10^{-381}$
$90$ 148.6791376989272828780655871531489156030918416831048594377087308271848378148.6791376989272828780655871531489156030918416831048594377087308271848378 $1.26\times 10^{-377}$
$91$ 150.5625106711582684678831224180073436626247034885508876594372802953139708150.5625106711582684678831224180073436626247034885508876594372802953139708 $3.19\times 10^{-380}$
$92$ 152.1147485501154030930412343288739204270768339963739180340167096854953029152.1147485501154030930412343288739204270768339963739180340167096854953029 $5.10\times 10^{-358}$
$93$ 153.2874104994293314602247720515393224042042637960724660197202500041048058153.2874104994293314602247720515393224042042637960724660197202500041048058 $1.82\times 10^{-381}$
$94$ 154.4907156955163636384604753696533838899948856560427610564143071072631799154.4907156955163636384604753696533838899948856560427610564143071072631799 $8.89\times 10^{-382}$
$95$ 155.2420799990558198969997928035576763043815202860039922762556596679189052155.2420799990558198969997928035576763043815202860039922762556596679189052 $5.80\times 10^{-387}$
$96$ 156.4997285439414899913040309653704544908218626069145493265791039370205636156.4997285439414899913040309653704544908218626069145493265791039370205636 $2.98\times 10^{-384}$
$97$ 158.4000270418246442479812803196489920645070375885184935734965513382773251158.4000270418246442479812803196489920645070375885184935734965513382773251 $2.31\times 10^{-375}$
$98$ 159.8727775733136235916063341758884923622323802914269809997571708512468331159.8727775733136235916063341758884923622323802914269809997571708512468331 $1.82\times 10^{-383}$
$99$ 161.2462175616940906041660713748861485716277484845435577262869276737895404161.2462175616940906041660713748861485716277484845435577262869276737895404 $1.88\times 10^{-377}$
$100$ 161.7676625972346591307869118758945874585239556380549233208831548174780607161.7676625972346591307869118758945874585239556380549233208831548174780607 $1.80\times 10^{-377}$
$101$ 163.7576737283108406276969780794254368523306129767593543960804767107794298163.7576737283108406276969780794254368523306129767593543960804767107794298 $3.17\times 10^{-381}$
$102$ 164.6120093931756454229907543346819925745086105838034366833937267160262192164.6120093931756454229907543346819925745086105838034366833937267160262192 $8.98\times 10^{-383}$
$103$ 165.7849132655712156658656369618645110416238945585780065340278032733366523165.7849132655712156658656369618645110416238945585780065340278032733366523 $1.38\times 10^{-370}$
$104$ 166.8984076627878735371910668394376277812744794403824180419885235916568118166.8984076627878735371910668394376277812744794403824180419885235916568118 $1.23\times 10^{-378}$
$105$ 168.8542152615191566500324150119276844393780975301779232829014664403452900168.8542152615191566500324150119276844393780975301779232829014664403452900 $3.55\times 10^{-382}$
$106$ 170.4329617627562106600626924828647481435479686833022554790466497554544179170.4329617627562106600626924828647481435479686833022554790466497554544179 $1.79\times 10^{-376}$
$107$ 171.3435913767245348564273929667373606132103332443943240617854291644166380171.3435913767245348564273929667373606132103332443943240617854291644166380 $9.15\times 10^{-378}$
$108$ 172.1109726296460068916462074637535861735937099184928581139379489286992488172.1109726296460068916462074637535861735937099184928581139379489286992488 $5.51\times 10^{-374}$
$109$ 173.5594350265921361471804589736353172594139337063764936096514386889343890173.5594350265921361471804589736353172594139337063764936096514386889343890 $5.40\times 10^{-380}$
$110$ 174.6387255634250548260060269694940210630117069036506518520953883247789503174.6387255634250548260060269694940210630117069036506518520953883247789503 $3.05\times 10^{-381}$
$111$ 176.3560413209382041587539791440362116352494681808774194880523226628720576176.3560413209382041587539791440362116352494681808774194880523226628720576 $3.84\times 10^{-381}$
$112$ 177.7259420577480027808483631140575011960193977129787882127241716149811312177.7259420577480027808483631140575011960193977129787882127241716149811312 $2.41\times 10^{-377}$
$113$ 178.7928649865600011304034350380931786953282017834290193659704381818837076178.7928649865600011304034350380931786953282017834290193659704381818837076 $2.93\times 10^{-378}$
$114$ 180.2429351355140552783405349658893752778995686711530121041367382571840255180.2429351355140552783405349658893752778995686711530121041367382571840255 $9.90\times 10^{-380}$
$115$ 181.5690410198014067424659236702507599683106719751432109652783222557297654181.5690410198014067424659236702507599683106719751432109652783222557297654 $7.23\times 10^{-381}$
$116$ 182.3836847672303153650140007274053516339810495569488300356257173446970099182.3836847672303153650140007274053516339810495569488300356257173446970099 $2.53\times 10^{-381}$
$117$ 183.7977731855809726856377575844467129141999715376864402640820344216736574183.7977731855809726856377575844467129141999715376864402640820344216736574 $2.55\times 10^{-377}$
$118$ 184.3628313266472803306747292477770104874008484280456362915229834946361820184.3628313266472803306747292477770104874008484280456362915229834946361820 $5.43\times 10^{-377}$
$119$ 186.8960606802271134804483399887402511596599811946354115671180347583314058186.8960606802271134804483399887402511596599811946354115671180347583314058 $4.34\times 10^{-388}$
$120$ 187.9450259874231159867364241809972240651568965424476790105736980715208122187.9450259874231159867364241809972240651568965424476790105736980715208122 $2.98\times 10^{-388}$
$121$ 188.9594686746299429679024157534003407512673176869606378971682058136033849188.9594686746299429679024157534003407512673176869606378971682058136033849 $5.49\times 10^{-391}$
$122$ 190.1395967633932102576693456196045454015876465926755736031997302414977464190.1395967633932102576693456196045454015876465926755736031997302414977464 $2.76\times 10^{-381}$
$123$ 190.9702222761120917217416693195514720792263824968025813693727397057385027190.9702222761120917217416693195514720792263824968025813693727397057385027 $7.36\times 10^{-380}$
$124$ 192.8230754082351124952403107100924733903825477854542116112915049676775859192.8230754082351124952403107100924733903825477854542116112915049676775859 $5.45\times 10^{-371}$
$125$ 193.5366786394328464762729501367958508864010937755963350086345404391030844193.5366786394328464762729501367958508864010937755963350086345404391030844 $1.91\times 10^{-375}$
$126$ 195.0968146886402893063805930855068198605011968413073660774300309263219276195.0968146886402893063805930855068198605011968413073660774300309263219276 $4.14\times 10^{-384}$
$127$ 196.5407645624496726701873374652927967937458532432874977774250971221715843196.5407645624496726701873374652927967937458532432874977774250971221715843 $8.75\times 10^{-384}$
$128$ 197.8610095077236677152139807981406000337724933271975112154503287411577548197.8610095077236677152139807981406000337724933271975112154503287411577548 $3.94\times 10^{-380}$
$129$ 199.0910217062412590553199980023993820450668536563004861622271199841578007199.0910217062412590553199980023993820450668536563004861622271199841578007 $2.65\times 10^{-374}$
$130$ 200.2454408454250041127516180550560747573455335829995291340148638052046280200.2454408454250041127516180550560747573455335829995291340148638052046280 $5.25\times 10^{-388}$
$131$ 200.7540601401830854485830596873971820322031065551151081003773144229516988200.7540601401830854485830596873971820322031065551151081003773144229516988 $7.64\times 10^{-376}$
$132$ 202.1949514428841149296373572141083379553219436587582390754534267698457479202.1949514428841149296373572141083379553219436587582390754534267698457479 $4.15\times 10^{-379}$
$133$ 203.8439019449593046704790586638384595752825730515914805602754963931582331203.8439019449593046704790586638384595752825730515914805602754963931582331 $1.42\times 10^{-383}$
$134$ 205.7461765204525832977848290088585427085408467101460279140881399490263847205.7461765204525832977848290088585427085408467101460279140881399490263847 $1.64\times 10^{-373}$
$135$ 206.1614100930471049929327260983840526456274838825113042507761929781360005206.1614100930471049929327260983840526456274838825113042507761929781360005 $8.14\times 10^{-365}$
$136$ 207.4576800789717454592941379191148719628980100114958536470176978762099884207.4576800789717454592941379191148719628980100114958536470176978762099884 $6.80\times 10^{-382}$
$137$ 208.6257624071707186435904203396570386255936601565172000778853597619116439208.6257624071707186435904203396570386255936601565172000778853597619116439 $8.83\times 10^{-383}$
$138$ 210.3018863035253963539783240245490373617263581479030048871463744638537098210.3018863035253963539783240245490373617263581479030048871463744638537098 $1.87\times 10^{-375}$
$139$ 210.6774621945036653303536974602005831137298996558343893089603787234105345210.6774621945036653303536974602005831137298996558343893089603787234105345 $1.08\times 10^{-374}$
$140$ 212.2314465116863399395470488677082962419973133942891768244390975438195060212.2314465116863399395470488677082962419973133942891768244390975438195060 $1.32\times 10^{-376}$
$141$ 213.5208741407826107064248292623018184776262412402697488085002636625114223213.5208741407826107064248292623018184776262412402697488085002636625114223 $2.14\times 10^{-375}$
$142$ 215.4464514511933389574877997496065789767019445673655729715342948390078403215.4464514511933389574877997496065789767019445673655729715342948390078403 $1.21\times 10^{-379}$
$143$ 216.3400341493431290616934365527790259377125944167236586852358894059490615216.3400341493431290616934365527790259377125944167236586852358894059490615 $5.63\times 10^{-385}$
$144$ 217.1983339649424303276759751119351973870730866826808674647867383696187534217.1983339649424303276759751119351973870730866826808674647867383696187534 $1.23\times 10^{-383}$
$145$ 218.4468831000582828049377925521014847112815527861307939822001313655805450218.4468831000582828049377925521014847112815527861307939822001313655805450 $4.45\times 10^{-378}$
$146$ 219.2620902163803991081302972577707057420124776154947727734971908483004301219.2620902163803991081302972577707057420124776154947727734971908483004301 $1.00\times 10^{-377}$
$147$ 220.7873883926188375982815560258438018537642942771892118826731601798907683220.7873883926188375982815560258438018537642942771892118826731601798907683 $2.09\times 10^{-378}$
$148$ 222.4669542636690272742987871089656750292277107765426971456433297333431582222.4669542636690272742987871089656750292277107765426971456433297333431582 $3.83\times 10^{-386}$
$149$ 223.5251748458997932788111875920239156826963314614886757387243060871678781223.5251748458997932788111875920239156826963314614886757387243060871678781 $2.14\times 10^{-385}$
$150$ 224.5905695547402250280725248575841596376488003389755247988103304379957061224.5905695547402250280725248575841596376488003389755247988103304379957061 $1.70\times 10^{-385}$
$151$ 225.9528174460175420018709628055980028796204875828523540302602967018854970225.9528174460175420018709628055980028796204875828523540302602967018854970 $3.77\times 10^{-375}$
$152$ 227.0506152396870075343232124376903091206351395063148200355197487967584741227.0506152396870075343232124376903091206351395063148200355197487967584741 $1.14\times 10^{-376}$
$153$ 228.2702606683105171006131439160963909648495727623770036781154541424529611228.2702606683105171006131439160963909648495727623770036781154541424529611 $7.10\times 10^{-395}$
$154$ 229.1293411187776318518182319293151826798543269243089746674150037521807965229.1293411187776318518182319293151826798543269243089746674150037521807965 $5.38\times 10^{-381}$
$155$ 230.0655467222499994960335512227034957685515138144609931348464490154193242230.0655467222499994960335512227034957685515138144609931348464490154193242 $3.23\times 10^{-377}$
$156$ 232.3504135703170106251813611501929285174003988479010053979053456457170580232.3504135703170106251813611501929285174003988479010053979053456457170580 $8.80\times 10^{-376}$
$157$ 233.2789359169850950528529859180644272121464846403423310620367158643774643233.2789359169850950528529859180644272121464846403423310620367158643774643 $2.04\times 10^{-365}$
$158$ 234.5851234236159003579250739972324756969591909615171896192278628761784272234.5851234236159003579250739972324756969591909615171896192278628761784272 $2.39\times 10^{-382}$
$159$ 235.2781440849719089723467916576609715985263047190969208147937131517174774235.2781440849719089723467916576609715985263047190969208147937131517174774 $1.35\times 10^{-377}$
$160$ 236.3194651062133632897797461259001706212369979274947636606085180194179956236.3194651062133632897797461259001706212369979274947636606085180194179956 $5.07\times 10^{-380}$
$161$ 237.7502459859493091409947528335486909271247941982061655950721576661275541237.7502459859493091409947528335486909271247941982061655950721576661275541 $2.57\times 10^{-386}$
$162$ 239.0184393184109716162860983894680542608908237139554551956920615988835233239.0184393184109716162860983894680542608908237139554551956920615988835233 $4.09\times 10^{-379}$
$163$ 240.0792385969133906036036875299045927021072746394399968671168058197336227240.0792385969133906036036875299045927021072746394399968671168058197336227 $6.97\times 10^{-372}$
$164$ 241.8295381920981108063036589846412192278335217717644854509984348418264849241.8295381920981108063036589846412192278335217717644854509984348418264849 $2.61\times 10^{-382}$
$165$ 242.4714484372455424871326046814432313933150408236493003286320498835996632242.4714484372455424871326046814432313933150408236493003286320498835996632 $4.82\times 10^{-377}$
$166$ 244.2781399513230979204344600038827819307672678138600762150930635163471547244.2781399513230979204344600038827819307672678138600762150930635163471547 $9.40\times 10^{-377}$
$167$ 245.0931465421787144819401091282678929577774107066028997725210154516210604245.0931465421787144819401091282678929577774107066028997725210154516210604 $4.80\times 10^{-380}$
$168$ 245.9414461421440943632894413679193298276606412507251949620382497378141750245.9414461421440943632894413679193298276606412507251949620382497378141750 $1.57\times 10^{-382}$
$169$ 247.0492019752380641386980109436073416805690975976922943665554413430020982247.0492019752380641386980109436073416805690975976922943665554413430020982 $1.32\times 10^{-377}$
$170$ 248.1925916339491196119212745608860106806155866118641532607170961702878151248.1925916339491196119212745608860106806155866118641532607170961702878151 $2.68\times 10^{-377}$
$171$ 250.2567812704718103514686964376094643921037936497368641231857765654186429250.2567812704718103514686964376094643921037936497368641231857765654186429 $9.38\times 10^{-378}$
$172$ 251.2412027861625789598829438472498407054321540732312842605464121756358711251.2412027861625789598829438472498407054321540732312842605464121756358711 $1.29\times 10^{-381}$
$173$ 252.4426428674970154269684848345320509970820863433649235621051571689997896252.4426428674970154269684848345320509970820863433649235621051571689997896 $7.61\times 10^{-375}$
$174$ 252.8783455516520261180992849726364616022483502282371925937689803304575025252.8783455516520261180992849726364616022483502282371925937689803304575025 $2.97\times 10^{-386}$
$175$ 254.6995428110910320634917223831419085806934447995403907767148534583035213254.6995428110910320634917223831419085806934447995403907767148534583035213 $7.96\times 10^{-390}$
$176$ 255.4817143518758676477488342660296610784346900622599694160158762168038100255.4817143518758676477488342660296610784346900622599694160158762168038100 $2.68\times 10^{-377}$
$177$ 256.7950960082609229989701144912192003822522893866960220291120141146446240256.7950960082609229989701144912192003822522893866960220291120141146446240 $3.36\times 10^{-390}$
$178$ 257.8270419645350378752623872886698073788539547918788452992788182042801964257.8270419645350378752623872886698073788539547918788452992788182042801964 $2.08\times 10^{-382}$
$179$ 259.2255084437328815672830991211725369672295555099562231070174147085271455259.2255084437328815672830991211725369672295555099562231070174147085271455 $5.78\times 10^{-380}$
$180$ 260.9015569091696849005739379537875300859589564012430210977138432993148113260.9015569091696849005739379537875300859589564012430210977138432993148113 $8.27\times 10^{-381}$
$181$ 261.8358472432834896832088038634100654365581736536704503687403981095608965261.8358472432834896832088038634100654365581736536704503687403981095608965 $1.35\times 10^{-387}$
$182$ 262.7253963956907899818036537593595936244700144938763582544067473174862290262.7253963956907899818036537593595936244700144938763582544067473174862290 $1.83\times 10^{-384}$
$183$ 263.9390328653304079227708108735342563292906200171500032175808474823259859263.9390328653304079227708108735342563292906200171500032175808474823259859 $7.69\times 10^{-380}$
$184$ 264.5895165938748657971375430173661706120250839803910618482150408087519377264.5895165938748657971375430173661706120250839803910618482150408087519377 $4.02\times 10^{-377}$
$185$ 266.0183558387811141106471906627987813950683933915872516370073933846309520266.0183558387811141106471906627987813950683933915872516370073933846309520 $3.47\times 10^{-380}$
$186$ 267.8667418320440883029977884107331094630768090317251879195806815839147212267.8667418320440883029977884107331094630768090317251879195806815839147212 $1.67\times 10^{-389}$
$187$ 268.7689730154733279799219030251031172816866175199190213423831114937875265268.7689730154733279799219030251031172816866175199190213423831114937875265 $1.73\times 10^{-381}$
$188$ 269.9200199321573070244666163067208075924583827953306793681648173800547888269.9200199321573070244666163067208075924583827953306793681648173800547888 $9.96\times 10^{-391}$
$189$ 270.9766252279049637196609662386223365123366632274390633476941617936694696270.9766252279049637196609662386223365123366632274390633476941617936694696 $4.64\times 10^{-394}$
$190$ 272.1129028149579662281710910682696801006758946368940041873249262575673809272.1129028149579662281710910682696801006758946368940041873249262575673809 $7.44\times 10^{-383}$
$191$ 273.5109563898493936962954831652231625581122076586144926745499431703165083273.5109563898493936962954831652231625581122076586144926745499431703165083 $8.12\times 10^{-376}$
$192$ 274.2948564917397463747467022749890819784095973810447351106601817286625286274.2948564917397463747467022749890819784095973810447351106601817286625286 $1.54\times 10^{-374}$
$193$ 275.0855464180449333603542313648026743574259611909815872140275993159179162275.0855464180449333603542313648026743574259611909815872140275993159179162 $1.43\times 10^{-374}$
$194$ 276.8441204674513326407659415272867236821569635669514355328674529440816966276.8441204674513326407659415272867236821569635669514355328674529440816966 $1.65\times 10^{-391}$
$195$ 278.5259446665362176488220010523114851640151849334569881154186200503119665278.5259446665362176488220010523114851640151849334569881154186200503119665 $9.24\times 10^{-377}$
$196$ 279.3017347241840231776964484759311017609689907608243870872318843780615467279.3017347241840231776964484759311017609689907608243870872318843780615467 $1.74\times 10^{-384}$
$197$ 280.4958776669325972199294145217874454617874575930120042656955303729513047280.4958776669325972199294145217874454617874575930120042656955303729513047 $3.10\times 10^{-367}$
$198$ 281.2600352094961963450785533504251722252905268262412512620081007078676192281.2600352094961963450785533504251722252905268262412512620081007078676192 $1.42\times 10^{-376}$
$199$ 282.1700465556285859901194920915754510495511952727753657159811617405866204282.1700465556285859901194920915754510495511952727753657159811617405866204 $3.52\times 10^{-377}$
$200$ 283.9346051850069315598211536467913532192999161825103775945871160416794839283.9346051850069315598211536467913532192999161825103775945871160416794839 $8.19\times 10^{-379}$
Appendix C Quadratic $L$-Function Zeros to 20 Decimal Places (Seven Discriminants)
The first 20 zeros of $L(s,\chi_{d})$ for discriminants $d\in\{2,3,6,7,10,11,13\}$ are given here to 20 decimal places. Certified upper bounds on $|L(\tfrac{1}{2}+i\gamma_{k}^{\prime},\chi)|$ appear in the rightmost column and are summarized in Appendix E. For computational methodology, see §A.1A.2.
$L(s,\chi_{2})$: field $\mathbb{Q}(\sqrt{2})$, conductor $q=8$.
$\chi_{2}(1)=\chi_{2}(7)=1$, $\chi_{2}(3)=\chi_{2}(5)=-1$, $\chi_{2}(n)=0$ for $n$ even. The first three zeros are used in Step 3 of the proof of Theorem 3.1.
ARB Bound ARB Bound
$n$ $\gamma_{n}^{\prime}$ on $|L(\tfrac{1}{2}+i\gamma_{n}^{\prime})|$ $n$ $\gamma_{n}^{\prime}$ on $|L(\tfrac{1}{2}+i\gamma_{n}^{\prime})|$
11 $4.89997399700703650104$ $1\times 10^{-31} 1111$ $26.95853518038046741969$ $8.98\times 10^{-32}$
22 $7.62842884176939783416$ $1.32\times 10^{-31}$ $12$ $28.09744496063079735462$ $2.57\times 10^{-31}$
33 $10.8065881638617120144$ $1.52\times 10^{-31}$ $13$ $29.93076421015190503723$ $2.66\times 10^{-31}$
44 $12.3105429942365296811$ $1.31\times 10^{-31}$ $14$ $31.63813949132101731704$ $2.90\times 10^{-32}$
55 $15.1957542506451227685$ $2.61\times 10^{-31}$ $15$ $33.84563089515844186412$ $7.91\times 10^{-32}$
66 $17.0222859743083473390$ $1.70\times 10^{-31}$ $16$ $34.74577700617002330947$ $5.29\times 10^{-32}$
77 $18.8059589077071483998$ $4.50\times 10^{-32}$ $17$ $36.54166385177458968906$ $2.68\times 10^{-31}$
88 $21.1316459622213438826$ $2.91\times 10^{-31}$ $18$ $38.77557777387006581424$ $1.96\times 10^{-31}$
99 $23.0838499962005465429$ $1.06\times 10^{-31}$ 1919 $39.78688099766664071794$ $2.48\times 10^{-31}$
$10$ $24.2019635578156016126$ $1.91\times 10^{-31}$ 2020 $41.34220616442193704510$ $3.51\times 10^{-31}$
$L(s,\chi_{3})$: field $\mathbb{Q}(\sqrt{3})$, conductor $q=12$.
$\chi_{3}(1)=\chi_{3}(11)=1$, $\chi_{3}(5)=\chi_{3}(7)=-1$, $\chi_{3}(n)=0$ for $\gcd(n,12)>1$.
ARB Bound ARB Bound
$n$ $\gamma_{n}^{\prime}$ on $|L(\tfrac{1}{2}+i\gamma_{n}^{\prime})|$ $n$ $\gamma_{n}^{\prime}$ on $|L(\tfrac{1}{2}+i\gamma_{n}^{\prime})|$
11 $3.80462763305086509714$ $5.15\times 10^{-32}$ $11$ $23.56131971313784471635$ $3.18\times 10^{-31}$
22 $6.69222332050013115991$ $2\times 10^{-31} 1212$ $25.41163389239269596203$ $2.13\times 10^{-31}$
33 $8.89059295872674150934$ $2.54\times 10^{-31}$ $13$ $27.01394398585067234435$ $2.92\times 10^{-31}$
44 $11.1883927450749091542$ $2.58\times 10^{-31}$ $14$ $28.44220325774560811851$ $2.64\times 10^{-31}$
55 $12.9661788080288453750$ $2.81\times 10^{-31}$ $15$ $30.20400655643888185431$ $2.43\times 10^{-31}$
66 $15.1814808758882167354$ $1.02\times 10^{-31}$ $16$ $31.64807761494068636690$ $5.56\times 10^{-33}$
77 $16.6326332745237621232$ $5.47\times 10^{-32}$ $17$ $33.03713287950650360192$ $1.05\times 10^{-31}$
88 $18.8843694571206508082$ $3.57\times 10^{-32}$ $18$ $35.02737848501270258165$ $6.67\times 10^{-32}$
99 $20.1039281912458185731$ $9.81\times 10^{-32}$ 1919 $35.77804457652351027242$ $2.50\times 10^{-32}$
$10$ $22.2858391072268876587$ $1.51\times 10^{-31}$ 2020 $37.92681682137521283529$ $3.25\times 10^{-31}$
$L(s,\chi_{13})$: field $\mathbb{Q}(\sqrt{13})$, conductor $q=13$.
$\chi_{13}$ is the unique primitive quadratic character of conductor $13$ (Legendre symbol $\bigl(\tfrac{\cdot}{13}\bigr)$); $13\equiv 1\pmod{4}$ so the fundamental discriminant equals the conductor.
ARB Bound ARB Bound
$n$ $\gamma_{n}^{\prime}$ on $|L(\tfrac{1}{2}+i\gamma_{n}^{\prime})|$ $n$ $\gamma_{n}^{\prime}$ on $|L(\tfrac{1}{2}+i\gamma_{n}^{\prime})|$
11 $3.11934147900860341390$ $4.45\times 10^{-32}$ $11$ $23.59202788517293715039$ $3.77\times 10^{-32}$
22 $7.23159073941876201503$ $1.17\times 10^{-31}$ $12$ $25.37170440577142621541$ $4.33\times 10^{-32}$
33 $8.62542663503259159735$ $4.11\times 10^{-32}$ $13$ $26.38431352575075899279$ $6.00\times 10^{-32}$
44 $10.3364207262315390298$ $1.36\times 10^{-31}$ $14$ $27.64980842472696812713$ $2.33\times 10^{-31}$
55 $12.6170127910231787304$ $4.72\times 10^{-32}$ $15$ $29.39145603391187839556$ $3.70\times 10^{-31}$
66 $15.1483324170057442246$ $2.53\times 10^{-31}$ $16$ $31.01874209224480391871$ $8.86\times 10^{-32}$
77 $16.2748260574985884179$ $8.23\times 10^{-32}$ $17$ $32.42341728496373366098$ $2.06\times 10^{-31}$
88 $18.7512523562342329165$ $1.65\times 10^{-32}$ $18$ $34.46644233774365394409$ $4.03\times 10^{-31}$
99 $19.5480414433470398699$ $1.19\times 10^{-31}$ 1919 $35.76319303421611596021$ $6.16\times 10^{-32}$
$10$ $20.9591819174050311814$ $3.35\times 10^{-31}$ 2020 $36.72754705153707320059$ $1.42\times 10^{-31}$
$L(s,\chi_{6})$: field $\mathbb{Q}(\sqrt{6})$, conductor $q=24$.
The fundamental discriminant of $\mathbb{Q}(\sqrt{6})$ is $\Delta_{K}=24$, since $6\equiv 2\pmod{4}$.
ARB Bound ARB Bound
$n$ $\gamma_{n}^{\prime}$ on $|L(\tfrac{1}{2}+i\gamma_{n}^{\prime})|$ $n$ $\gamma_{n}^{\prime}$ on $|L(\tfrac{1}{2}+i\gamma_{n}^{\prime})|$
11 $2.68865813246759758667$ $41\times 10^{-31} 11$11 $20.03461157675018829473$ $1.62\times 10^{-31}$
22 $5.29243117677719815160$ $7\times 10^{-31} 1212$ $21.55547824218793740140$ $3.62\times 10^{-31}$
33 $6.97192424379172752833$ $1.20\times 10^{-31}$ $13$ $22.66785732205079621070$ $9.70\times 10^{-32}$
44 $9.22463743993556115082$ $2\times 10^{-31} 1414$ $24.40914964223159369360$ $2.18\times 10^{-31}$
55 $10.4457215184129248230$ $2.97\times 10^{-31}$ $15$ $25.67129733212202220561$ $2.53\times 10^{-31}$
66 $12.6311537181665105855$ $3.28\times 10^{-31}$ $16$ $26.75610741237823107924$ $3.22\times 10^{-31}$
77 $13.7771527057734958652$ $2.51\times 10^{-31}$ $17$ $28.86856257427073136901$ $1.56\times 10^{-31}$
88 $15.6036569703702537829$ $3.03\times 10^{-31}$ $18$ $29.21969083370577778290$ $1.10\times 10^{-31}$
99 $17.0989880869194034869$ $1.05\times 10^{-31}$ 1919 $31.08725180031803397862$ $4.04\times 10^{-31}$
$10$ $18.4632988501214004408$ $1.83\times 10^{-31}$ 2020 $32.42353686755135030794$ $3.61\times 10^{-31}$
$L(s,\chi_{7})$: field $\mathbb{Q}(\sqrt{7})$, conductor $q=28$.
The fundamental discriminant of $\mathbb{Q}(\sqrt{7})$ is $\Delta_{K}=28$, since $7\equiv 3\pmod{4}$.
ARB Bound ARB Bound
$n$ $\gamma_{n}^{\prime}$ on $|L(\tfrac{1}{2}+i\gamma_{n}^{\prime})|$ $n$ $\gamma_{n}^{\prime}$ on $|L(\tfrac{1}{2}+i\gamma_{n}^{\prime})|$
11 $2.77728324207659233094$ $6.55\times 10^{-33}$ $11$ $19.53861609989941145163$ $2.70\times 10^{-31}$
22 $4.52677408701547610929$ $2.86\times 10^{-31}$ $12$ $20.39328285450433247153$ $7.16\times 10^{-32}$
33 $7.29088506700227937479$ $2.61\times 10^{-31}$ $13$ $21.98889814760568313347$ $3.40\times 10^{-31}$
44 $8.39512407283755124263$ $1.74\times 10^{-31}$ $14$ $23.69153462655719640538$ $2.67\times 10^{-32}$
55 $10.1163068547312977911$ $2.08\times 10^{-31}$ $15$ $24.73886680489813285047$ $3.18\times 10^{-31}$
66 $12.0220620381859538014$ $3.73\times 10^{-32}$ $16$ $26.38179066809542195017$ $1.65\times 10^{-31}$
77 $13.5442863718297874149$ $5.68\times 10^{-32}$ $17$ $27.02957442259405572955$ $1.67\times 10^{-31}$
88 $15.0682949619190434065$ $1.90\times 10^{-31}$ $18$ $28.72073200609003118776$ $5.74\times 10^{-31}$
99 $16.0162614771505928089$ $9\times 10^{-32} 1919$ $30.39449098988413064026$ $3.27\times 10^{-31}$
$10$ $18.2004252754926244593$ $1.75\times 10^{-31}$ 2020 $31.36586297616383901441$ $2.58\times 10^{-31}$
$L(s,\chi_{10})$: field $\mathbb{Q}(\sqrt{10})$, conductor $q=40$.
The fundamental discriminant of $\mathbb{Q}(\sqrt{10})$ is $\Delta_{K}=40$, since $10\equiv 2\pmod{4}$.
ARB Bound ARB Bound
$n$ $\gamma_{n}^{\prime}$ on $|L(\tfrac{1}{2}+i\gamma_{n}^{\prime})|$ $n$ $\gamma_{n}^{\prime}$ on $|L(\tfrac{1}{2}+i\gamma_{n}^{\prime})|$
11 $2.48821020930540853564$ $9.61\times 10^{-32}$ $11$ $18.18218004680212107005$ $2.46\times 10^{-32}$
22 $4.03159701418841483371$ $3.83\times 10^{-32}$ $12$ $19.28135808746821972149$ $3.45\times 10^{-31}$
33 $6.33545319091356197520$ $3.06\times 10^{-31}$ $13$ $20.40160148601184344650$ $3.27\times 10^{-31}$
44 $7.98779634495761720535$ $1.35\times 10^{-31}$ $14$ $21.56472815482081368499$ $3.86\times 10^{-31}$
55 $9.28044435522920809582$ $1.09\times 10^{-31}$ $15$ $23.08868709418032385846$ $1.74\times 10^{-31}$
66 $10.6382769795389180751$ $1.54\times 10^{-31}$ $16$ $24.43655444208668619203$ $2.25\times 10^{-31}$
77 $12.8151786361302353694$ $5.75\times 10^{-32}$ $17$ $25.87097298765072715474$ $1.42\times 10^{-31}$
88 $13.6043079656071037781$ $1.37\times 10^{-31}$ $18$ $26.26140271050477788697$ $1.49\times 10^{-31}$
99 $15.2714728878515316883$ $8.72\times 10^{-32}$ 1919 $27.96504990212936504067$ $2.45\times 10^{-31}$
$10$ $16.1666343336440938522$ $3.22\times 10^{-31}$ 2020 $29.27531093215095949777$ $8.60\times 10^{-32}$
$L(s,\chi_{11})$: field $\mathbb{Q}(\sqrt{11})$, conductor $q=44$.
The fundamental discriminant of $\mathbb{Q}(\sqrt{11})$ is $\Delta_{K}=44$, since $11\equiv 3\pmod{4}$.
ARB Bound ARB Bound
$n$ $\gamma_{n}^{\prime}$ on $|L(\tfrac{1}{2}+i\gamma_{n}^{\prime})|$ $n$ $\gamma_{n}^{\prime}$ on $|L(\tfrac{1}{2}+i\gamma_{n}^{\prime})|$
11 $1.86993927318275168341$ $4.52\times 10^{-32}$ $11$ $17.61972287425101875532$ $9.04\times 10^{-32}$
22 $4.70920812840911269662$ $2.64\times 10^{-31}$ $12$ $18.26903309787599031651$ $2.28\times 10^{-31}$
33 $5.81981644686087148921$ $4.85\times 10^{-32}$ $13$ $20.30638125042707154147$ $2.32\times 10^{-31}$
44 $7.38790752248483336704$ $3.58\times 10^{-31}$ $14$ $21.66563972555785204781$ $3.05\times 10^{-31}$
55 $9.44566152389131881370$ $2.60\times 10^{-31}$ $15$ $22.32883476081719799605$ $6.78\times 10^{-32}$
66 $10.7407497975074755000$ $9.60\times 10^{-32}$ $16$ $24.04471027020575246775$ $6.45\times 10^{-32}$
77 $11.9563737277407041057$ $1.59\times 10^{-31}$ $17$ $24.68148321849740900845$ $6.86\times 10^{-33}$
88 $13.4164345305120382816$ $2.22\times 10^{-32}$ $18$ $26.44242345116378413034$ $5.29\times 10^{-31}$
99 $14.8319853387339990685$ $1.08\times 10^{-31}$ 1919 $27.60237248544315232420$ $3.84\times 10^{-31}$
$10$ $16.5192247034936462349$ $2.60\times 10^{-31}$ 2020 $28.77893942445047647073$ $2.36\times 10^{-31}$
Appendix D Riemann Zeta Zeros (20 Decimal Places)
The first 60 positive zero ordinates of $\zeta(s)$, to 20 decimal places. These values are used in the Besicovitch constants (Section 4), the spacelike verification (Table 2), and the cross-function disjointness check (Proposition 9.1). The certification of $S_{\zeta}^{*}$A.3) uses the first 6000 zeros from the LMFDB [9]; this is the only external data dependency in the paper.
$n$ $\gamma_{n}$ $n$ $\gamma_{n}$ $n$ $\gamma_{n}$
11 $14.13472514173469379045$ 2121 $79.33737502024936792276$ $41$ $124.25681855434576718473$
22 $21.02203963877155499262$ $22$ $82.91038085408603018316$ $42$ $127.51668387959649512427$
33 $25.01085758014568876321$ $23$ $84.73549298051705010573$ $43$ $129.57870419995605098576$
44 $30.42487612585951321031$ $24$ $87.42527461312522940653$ $44$ $131.08768853093265672356$
55 $32.93506158773918969066$ $25$ $88.80911120763446542368$ $45$ $133.49773720299758645013$
66 $37.58617815882567125721$ $26$ $92.49189927055848429625$ $46$ $134.75650975337387133132$
77 $40.91871901214749518739$ $27$ $94.65134404051988696659$ $47$ $138.11604205453344320019$
88 $43.32707328091499951949$ $28$ $95.87063422824530975874$ $48$ $139.73620895212138895045$
99 $48.00515088116715972794$ $29$ $98.83119421819369223332$ $49$ $141.12370740402112376194$
$10$ $49.77383247767230218191$ $30$ $101.31785100573139122878$ $50$ $143.11184580762063273940$
$11$ $52.97032147771446064414$ $31$ $103.72553804047833941639$ $51$ $146.00098248676551854740$
$12$ $56.44624769706339480436$ $32$ $105.44662305232609449367$ $52$ $147.42276534255960204952$
$13$ $59.34704400260235307965$ $33$ $107.16861118427640751512$ $53$ $150.05352042078488035143$
$14$ $60.83177852460980984425$ $34$ $111.02953554316967452465$ $54$ $150.92525761224146676185$
$15$ $65.11254404808160666087$ $35$ $111.87465917699263708561$ $55$ $153.02469381119889619825$
$16$ $67.07981052949417371447$ $36$ $114.32022091545271276589$ $56$ $156.11290929423786756975$
$17$ $69.54640171117397925292$ $37$ $116.22668032085755438216$ $57$ $157.59759181759405988753$
$18$ $72.06715767448190758252$ $38$ $118.79078286597621732297$ $58$ $158.84998817142049872417$
1919 $75.70469069908393316832$ $39$ $121.37012500242064591894$ $59$ $161.18896413759602751943$
2020 $77.14484006887480537268$ $40$ $122.94682929355258820081$ $60$ $163.03070968718198724331$
Appendix E ARB Certification of Zero Data
All 340 zeros tabulated in Appendices C and B were rigorously certified using the methodology of §A.1A.2. Working precision was 500 bits for the seven auxiliary characters and 1500 bits for $L(s,\chi_{5})$. The following table summarizes the certified upper bounds:
Character Field Zeros Certified upper bound
$\chi_{5}$ $\mathbb{Q}(\sqrt{5})$ 200 $6\times 10^{-358}$
$\chi_{2}$ $\mathbb{Q}(\sqrt{2})$ 20 $4\times 10^{-147}$
$\chi_{3}$ $\mathbb{Q}(\sqrt{3})$ 20 $4\times 10^{-147}$
$\chi_{13}$ $\mathbb{Q}(\sqrt{13})$ 20 $4\times 10^{-147}$
$\chi_{6}$ $\mathbb{Q}(\sqrt{6})$ 20 $4\times 10^{-147}$
$\chi_{7}$ $\mathbb{Q}(\sqrt{7})$ 20 $4\times 10^{-147}$
$\chi_{10}$ $\mathbb{Q}(\sqrt{10})$ 20 $4\times 10^{-147}$
$\chi_{11}$ $\mathbb{Q}(\sqrt{11})$ 20 $4\times 10^{-147}$
Total 340
All bounds remain far below any threshold required by the inequalities established in this paper.